A Complete Bibliography of Publications in Scientometrics: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

st [Kuk99]. _{c} [Chu92b]. N [BH95]. R [Tur94].

-rays [BH95].

1 [Hai92a, Kre99]. 1980s [Abt94, Mar94, ODOM98]. 1985 [Kla90]. 1989 [Sha90, Tod90]. 1993 [Ano93a]. 1995 [Ano95, Bon95, Le 95]. 1997 [Ano97a, Pav97, Sma97b]. 1999 [Ano99, Le 99]. 19th [WD97]. 19th-century [WD97]. 1st [Bar91b].

2 [Hai92b, KK99]. 20th [BWD96, WD97]. 20th-century [WD97].

80/20 [CCT93].

aboutness [Wil98]. abroad [Bon93, JCFA96, Ste98]. abstracts [BB99, RM97, GP99b, Spa97]. Academic [For92, Yam91, CM99, Esc99, GBL95, Kea91, Lin91b, RWG^{+}97, Ura90]. academician [Vin98d]. Academy [Vin98b]. accelerators [SB96b]. Access [Ada92, Whi92]. accomplishment [Ric91]. according [Lem92a]. account [RT95, RR98]. accounting [Hog90]. Accumulation [dHLR94, Fuj98]. accumulative [Luu90]. achievement [Ada92]. achievements [Kea91, LMHZYYS96]. achieving [Hic99]. acid [AL91a]. acidification [LW92]. acknowledgement [Lew94]. act [Bro96]. action [CCC94, Lew94]. activities [Hel95, Kla90, Kut98, LL93, PV91, PA91, RA98, San92, Sig96, Yam94]. activity [BZ97, CA92, DS91, GSM90, Mas91, NT96, SBG93, Ste93, Syl93, TW91]. adding [Ves95]. additional [Ley91]. address [Bar94, Gab95, Per97]. addresses [DM93]. adoption [RDC^{+}96]. Advanced [NV98, vR96, vR99, Dic96, PA91]. Advances [LW99, Fau90]. advantage [DLV99, WD95, dHLR94]. advantages [Luu90]. adverse [RV94, RKV95]. advertising [Lin96a]. advisors [Ano97b, FVWD98]. Advisory [Bor99, Lis99, Man99, BK90, Mey99, vH99]. afar [Shr97]. affect [Col94]. affecting [LOK99, Mea91]. affinity [So90]. Africa [Dah99]. African [Pou91]. Age [Bre96a, SL93, Coh91, GPK97, Gup97, LMHZYYS96, RT97]. age-structured [RT97]. ageing [GS94b, LM98, SMPK93]. agencies [Pap99, Pou93, Sei99]. aggregated [Ley94b]. aggregation [JJ92]. aging [Vin98d]. agreement [HH90]. Agricultural [SCQ93, DAD99, GS96, MJ93, MWHG93, NHHA94, NMMvR93, RP92]. Agronomique [DCL^{+}96]. agronomy [RMP^{+}98]. aided [DB95]. aids [MCRCNB98, MCSAMSL99, SK91, Sma94]. Aims [Sch97c]. air [Rou98]. algorithmic [HHM^{+}96]. alienation [MWHG93]. alive [BG96]. alternative [Ric91]. Ambiguities [POV91]. Ambiguity [BW97]. America [Can95, DHS99, KVKWD95, MCRCNB98, MCSAMSL99, NBFM92, NB95, Ves95, Zum97]. American [MVR96, FAM^{+}92, GOR97, GSMF99, GSS98, KETP97, LFJK93, MC92, MC94, Sen93, Yam91]. among [And99, HH90, Ley90b, Ley90c, MC92, Moe96a, PF95, Ura90]. amongst [Egg91]. analyses [Wil98, Wil99b]. analysing [FR98, HQD92]. Analysis [GFS99, Gup97, Kla95, MG93, UPS97, Ark99, Aru98, Bar91a, Bar92, BMSV91, BWD96, BSS97, BB98, BGJB92, BB92, Buc95, BDT97, CCL91, CLCL93, CS98, CI96, Cou94, De 96, DL97, DV92, EK95, FR98, Gio93, GSC99a, GSC99b, GSSC99, Gup99, HH90, HF99, HN93, JJ92, Kar90, KM97, Kla90, KW90, Kos97, Kos98, Lau95, Lin96a, LS95, LS91, Luu97, MM96a, MC95, MCRCNB98, MCSAMSL99, Mak98, MS94, McC98, McG96, MVR96, NS94, Nag99, NV97, NMV99, ODOM98, Opp97, OW97, Per94b, PV91, Phe99, PA91, QZ90, Raj96b, Rao98, RC94, RDQ^{+}98, RM97, RR97, Rou98, SK91, SB96b, Spa97, Syl93, TM99, TW98, TV98, TW90, Tod92, Tro90, Tsa95, TR91, Uzu90, UMÖ93, Uzu96, VGMP^{+}94, Van97b, Zit91, ZB94, dAVA99, dL94, eKOM^{+}98]. analytical [Ley92a, Ste93]. Analyzing [HH06]. Anatomy [QZ90]. András [Ano93a]. Angewandte [MVR96]. announce [Ano97b]. Announcement [Ano98]. annual [MM90]. Anthony [Ano95]. anything [Ley90b, Ley90c]. Apples [JJ92]. Applicability [GKR98]. applicant [LOK99]. application [BES90, DL97, GC96, JW99, Kre90, MS92, MP98, Pla98, Ste90, WD99, dL94]. Applications [Cun98, ER92, Kos95, MDV95]. applied [De 96, FR98]. Applying [Kor97]. appraisal [Pes92]. Approach [SBG93, Bar91a, Bar92, BG90a, CG97, DG97, Glä96a, GC96, HHM^{+}96, JD94, KV97, Kre99, Ley92a, NM93, RA93, SPC92, WD99]. approaches [Kos95, MK99, dMR99]. appropriate [MVR99]. aquaculture [Syl93]. aquatic [PF95]. Arab [EDM92]. area [BB98, Gup99]. areas [Bha97, BGG95a, BGG95b, Glä96a, HG96b, Mil94, MOO^{+}95, Pou91, VT93a]. Argentina [QAG95]. aromatic [GSM90]. art [OPMU92, Van97c]. article [GKR98, Oku93, Seg96]. articles [AL91b, Bal96, CH93, GOR97, Ove90, SH91, Yit94]. Artificial [PFK98]. arts [HG96a]. Asia [ODOM98]. Asian [ASR94]. asking [vR98]. aspects [Egg97, GH94, Kyr95, RG98, Vin95, Vin96c, ZT96]. assesment [Pla98]. assess [Kra92]. Assessing [Van97a, VT93a, Gil91, MGB93, RM97]. Assessment [KVA95, KV97, NN92, SRMP99, BGS91, Bun99b, Coz95, Eto94, HHM^{+}96, Kos95, KETP97, Lin99, Mar91, Mar96, MDV95, NP93, Nag95a, Nag95b, Nyb99, Pom99, SCQ93, Tho95, dA92]. assessments [SB93]. assignment [Bal95, NVBG95, POV91, Sni95]. associated [DRC^{+}99, MWHG93, UH97]. association [HH06]. astronomers [DS91, Tri91, UÖ96]. astronomical [Jas91, Jas92, LS95]. astronomy [Fer98, Van94]. attitudes [And99]. attraction [Laz94]. Australia [DWH99]. Australian [BB96b, DR96, TW90]. Author [Ano93b, GK96, GK99, PV91, POV91, Row99, eKOM^{+}98]. authored [Bri91, NSW91]. authors [Chu94, GK97, GKR98, JCFA96, Lew96, PB95, Yit94]. Authorship [Ano96a, Bir97, Bre96b, CD97, De 97, Nor93, O’N98, Sam95, BGMS92, DG97, GD97, LFJK93, LI99, Vin93b, Whi93]. authorships [MP96]. autism [CG97]. autoimmune [Hin99]. automated [Bal95, NVBG95]. Average [ER96]. award [Ano93a, Ano95, Ano97a, Ano99, Bon95, Le 95, Le 99, Pav97, Sha90, Sma97b, Tod90]. Awards [Bon94b]. awareness [MDNT91]. Axioms [Kos95].

B. [Kea91]. Back [Rus94]. background [Kos96]. Balance [Vin93a, BSZ97]. BAP [Lew94]. base [TV98]. based [CG97, Dha98, ER92, GS93, Glä97, GLMM92, GSM90, HF99, Kre90, MVR96, Moe96a, NS95, NB95, O’N98, RWG^{+}97, RM97, SBG93, SB93, Sch96a, Sch96b, Sch96c, Sch96d, Sch99c, TM99, Tei91, Vin98b, ZB98, vR96]. bases [PN92, SQRD95]. basic [CCL91, Mar96]. basis [Fau90, Ley94b, MS94]. be [Bal95, Bal96, BG96, Kun90, NVBG95]. been [Luu97]. before [Yam91]. behavior [De 98, LLD90, MGCG96, RMP^{+}98]. behavioral [Bon91, NV97, NV99]. behaviour [CMS91, IW99, Kre97, Luu92, dS92]. behaviours [DV92]. Belgium [Rey92, vB92]. Belver [Ano97a, Sma97b]. Ben [Ano97a, Ley91, Pav97]. Bertram [Sha90]. Besides [GBL95]. best [Bon90b]. better [Avk97]. Between [LW99, Sig96, Agu96, CCL91, Chu92a, Chu92b, CCC94, CFKS92, DEP97, Egg99a, Fas99a, FQD97, FAM^{+}92, FVWD98, God98, HH06, HS98, HF99, KM90, Lin91b, Mak98, NO98, Nor90, PMR96, Rou94b, Rus98, Sch97a, Ste90, Tup99, Uzu90, Vin96b]. between-year [Agu96]. Beyond [Wou99, Egg94b, GS94a, McG94, Mea94, Miq94, Per94a, Rao94, Sch94b, TV95, vR94]. bias [OD99]. biased [Vin98c, Luw99]. bibliograph [Col94]. Bibliographic [SC98, DV92, GHL92, GC96, HW99, Phi92, Whi92, Whi93]. Bibliographical [Hin94, JSBCF91]. bibliographies [Col93, Sch95, SC98]. bibliography [Sch90a, Sch90b, Sch91, Sch96a, Sch96b, Sch96c, Sch96d, Sch99c]. Bibliometric [BGJB92, BB92, Bra99, DV92, GSS98, Kat96, LC91, MCRCNB98, MM96b, NH96, NHHA94, NvR94, QBR^{+}98, SPC92, SK91, SB96b, TV98, BGS90, BI97, Bha97, BB96b, BGS91, CDC95, Cun98, DHS99, Glä96a, Glä96b, GSC99a, GFS99, GH94, Gup99, Hai94, Her95, Hic99, KH97b, Laz96, Lew98, LS91, Luu97, MC95, Mar91, MDV95, Moe96a, Moe96b, Nag95a, NN92, NM95, PvR94, PA91, Rao96, RC94, SST98, SS92, Sni95, SQRD95, Syl93, Tho92b, TW98, Tij93, Uzu96, VMP98, VT93a, Zac91, dAVA99, lMD91, vR96, vR99]. Bibliometrical [QZ90]. Bibliometrics [Kre92, Sko99, Bar94, BMSV91, McG96, Nar94, Per90a, RG98]. bibliometry [Kea91]. bibliotheconomic [VGMP^{+}94]. Biennial [BB99]. big [Egg94b, GS94a, GPK97, Mar92, McG94, Mea94, Miq94, Per94a, Rao94, Sch94b, vR94, Mar92]. Bilateral [Vin93a]. binomial [LMRR92, PN92]. biochemistry [Men92, YBT91]. bioelectronics [Hin94]. Biofísica [PD98]. biogeography [Nor90]. Biological [YBT91, CLCL93, KVA95, RDC^{+}96]. biologists [WW93]. Biology [BB99, Her95, PF95, RA93, eKOM^{+}98]. biomass [Tho92b]. Biomedical [BGF^{+}96, Ano96a, BZ97, Bre96b, LD98, Lew99, SH96, Ste93]. biomedicine [DB95]. biotechnologies [dL94]. biotechnology [HG96b, Lew94, MS93, MS94, McC95, UH97]. birthday [Vin98c]. Black [Dah99]. Blind [LOK99]. blot [Tho92a]. Board [Ano97b]. bodies [Mey99, vH99]. body [BK90]. Book [Glä90, Ley94a, Rou90, Sch94a]. Bosković [Kla90]. Bovine [RWG^{+}97, RDQ^{+}98]. BPB [Vin93a]. Bradford [Boo94b, Bur91, CZ98, Chu94, Col93, Col94, GSS93, Nor90, Per90a, Rao98, WD97]. Braun [Mar91]. Brazil [GH95, LD96, LLD98, Nun99, Sik97]. Brazilian [De 98, Men92, Men96, PEZ98, PE99, PD98, Spa90]. breeds [MHZ98]. Bridging [Egg94a]. Britain [Kea91]. Brite [BCML^{+}95]. British [Abt94, CDC95, HN93, KH96, Ley91, Mar94]. Brookes [Sha90]. budget [Tod97]. Building [Sma94, BBG^{+}95]. built [ZB94]. business [Dou94].

C [Ano97a, Sha90, Sma97b]. Côte [CA92]. calculated [Vin98b]. calibrated [KH97b]. calibration [Lew99]. Call [Ano96b, Kre92]. Can [Bal95, DLV99, Kun90, NVBG95]. Canada [SA91]. Canadian [AD92, Smi95, Syl93]. cancer [SH91, UPS97]. candidates [For92]. capabilities [KVA95, KV97]. capacity [Fra91, MLH^{+}98]. capita [Tei94]. cardiovascular [RA93]. care [MC95]. careers [Tri91]. Caribbean [KVKWD95, MCRCNB98, MCSAMSL99]. Carlo [Cze90]. Carlos [PD98]. cartography [Hin94, NvR94, PFK98]. case [AL91a, ARR96, BI98b, BB96b, CCL91, CA92, CH93, DLV99, Fer98, GSS93, GI96, Her95, Hin94, HB92, Kra92, KVA95, Lem92a, Mau91, NB95, PD98, Per91, PvR94, RGK95, RRMS99, RMP^{+}98, RA93, Rou93, SH96, Sma94, Tho92a, Tho92b, TW98, WDB94, YBT91, BWD96]. categories [Sig96]. Category [Rou92]. Catolica [KVA95]. causes [Vin91b]. CD [Ley94b]. CED [BB96c]. CED-le [BB96c]. center [DS95, Laz94]. centers [MJ93]. Central [BS96, CR98, KRI99, PMR96, RA98]. centre [DCL^{+}96, Rin95]. century [BWD96, Kuk99, Nal91, WD97]. CERN [DV97, Mar92]. CFTRI [SR99b]. Chagas [PD98]. chairs [Bal95, NVBG95, Sni95]. challenge [Bon96]. challenger [DEP97]. Challenges [Kru99, Rip97, Smi95]. chance [Vin96b]. change [Bun99b, GD97, Lin99, Nyb99, Pom99]. Changes [RMP^{+}98, BSS97, MLH^{+}98, Sma93]. channels [Rab92]. chaos [Laz96]. characterisation [Hei95]. Characteristics [BBS97, Can95, Prp96a, GHL92, SOM91]. characterization [GS93]. characterizing [Vin93a]. Chemical [KETP97, MVR96, BMSV91, EK95, HS98, JGSK98, MG93, PvR94, Sch97b, Sch97c]. Chemie [MVR96]. Chemistry [Qur91, BGMS94b, BGG95a, CCL91, DQH91, Her93, KG99, MZ92, NP93, Sch97b, Sch97c, Vin99, YBT91]. chemists [Kla90]. Chile [KVA95, KV97, SMPK93]. Chilean [Kra92, Vog97, Zum97]. China [Dha98, Hai96, HH97, LS95, MADO98]. Chinese [Chu92a, Chu92b, Hai94, JW99, Lau95, RZ92]. circuit [Rab92]. circulation [LMRR92]. Citation [Agu96, Aru98, Cam96, CL91, EK95, FU92, JSBCF91, Kos97, MS96a, OD99, Rou98, Sch98, So98, UMÖ93, Win96, Wou97, ARR96, Bal96, BS91a, BY99, BB96a, BGMS94b, BGMS94a, CI96, Cro98, Cza97, CH93, Dah99, DB95, Egg98, ERY99, Fuj98, Gio93, Glä97, HF99, JW99, Kos98, Kre90, LLD90, LF99, Lau95, Ley94b, Ley98, LW99, Lin96a, Luu97, Luw99, MM96a, Mak98, MG93, Mil94, NSW91, Nie96, Opp97, OW97, Per90b, Per91, Per94b, Phe99, Pic95, PF95, Rou94a, Rou99a, SR99a, SB93, Sch96a, Sch96b, Sch96c, Sch96d, Sch99c, SL92, SA91, Sma93, So90, TV95, Tsa95, Wou99, Yit98, ZB96, ZB98, vR91, MZ92]. Citation-networks [Sch98]. citationist [KVKWD95]. citationology [Gar98]. Citations [Sma98a, ZMM91, Bal96, BH95, Bur95, Gup97, Hei95, Lew94, Luu97, MGB93, MVR96, RBBMJCC99b, TM99, Vin96b, WD94]. cited [BP98, GS92, LBR92, MG91, Plo90, Plo94, Ric91]. Citedness [SH91, Vin96b]. city [DQH91]. civil [Yam91]. CJNR [Nor93]. claims [RG98]. classical [Egg91]. Classics [Cam96, CL91]. Classification [ER92, BZ99, Chu94, GSC99b, GSSC99, GBFM96, PFK98, Spa97]. cleansing [LI99]. clearinghouse [vB92]. climate [SW99]. clinical [IW99, Per94b]. closed [MCGC95]. Closing [Bun99a, Ein99a, Rou99b]. Clustering [Bar91a, Bar92, DB95, vR91]. clusters [Mil94, RBBMJCC99b, Sma93]. CNRS [BMSV91]. Co [CCL91, Spa97, BB98, BGMS92, DG97, DL97, GD97, LW92, LFJK93, LI99, MP96, MG93, Mil94, MDNT91, NSW91, PV91, RBBMJCC99b, Sma93, TW90, Tod92, TR91, Van94, ZB96, vR91]. co-author [PV91]. co-authorship [BGMS92, DG97, GD97, LFJK93, LI99]. co-authorships [MP96]. co-citation [MG93, Mil94, Sma93, ZB96, vR91]. co-citations [RBBMJCC99b]. Co-classification [Spa97]. co-heading [TW90, Tod92]. co-operation [MDNT91, NSW91, Van94]. Co-word [CCL91, BB98, DL97, LW92, TR91]. coal [AD92]. Coauthorship [Kre94a, Kre97]. cocitation [McC98, Row99, ZB94]. cognition [CCC94]. Cognitive [UH97, And99, DM93, Prp94, Row99, Tij93]. cohesion [CdSM93, Hig91, Lew93]. Colima [Le 99]. Collaboration [BGF^{+}96, GKK97, GK98, GK99, Hin99, KG98, RGK95, ASR94, Avk97, BZRB99, BG96, DG97, EDM92, GW92, GD97, GI96, HH97, Kat94, KH97b, Kre99, KK99, LS91, LTPS93, MP96, Mel99, MP98, MO94, NBFM92, NB95, OMFD92, Oku93, PMDK97, Qin94, SB90, Van98, Vog97, KK99]. Collaborative [Gup93, MG92, Kun96, Men96, Row99]. colleagues [Rab92]. collections [Wil98, Wil99b]. collective [Bon90b, CM99, Fer98, Nun99]. collectivity [Bon90a, Bon90a]. colleges [Kre94a, Kre97]. Colombian [MCGC95]. combined [BGMS94b, BGG95a, Lin96a]. come [Luu94, WL94]. comment [Kea91]. Comments [Bon95, Cou92, Oku93, Pav97, Sha90, Sma97b, Tod90, Wes99, BGS91, Ley92b]. commission [Rin95]. committee [For92]. Common [BB96c]. commune [BB96c]. Communication [Chu92a, Chu92b, YPC^{+}93, Bon91, Ehi90, Lau95, LV97, MC92, MS93, RG99, SSS99, So98, TV98, Vin94a, ZB99]. communications [Bon96]. Communist [KRI99]. communities [AD92, Eto99a, Kre94a, Mak98, Mas91]. community [Jas91, Lew93, Lew94, Luu97, MDNT91, RF92, RT97, SMPK93]. companies [LCS^{+}91]. Comparative [Dha98, KETP97, MOO^{+}95, TM96, Vin98a, Bar91a, GD97, JSBCF91, Kea91, Pes92, TW98, Yit97]. comparision [GKA99]. Comparison [BZ97, Bha97, BS91a, Can95, DR96, DHS99, GHL92, GOR97, NO98, NN92, Qur93, SR99b, Tei94, Van94]. Comparisons [MO94, IM89]. comparitive [Bar92]. compendium [Phe99]. compete [GBL95]. competencies [CM99]. Competition [Fic90]. competitive [DLV99, MM96b]. competitors [GBL95]. Complete [QT90, Mil94]. complex [Kre90]. complexity [Lyo90]. composite [Dic96]. Computer [DB95, GPK97, Lyo90]. Computer-aided [DB95]. computing [CB95]. Concentration [BBS99, Rou93]. concept [Gru90a, Tsa95]. concepts [Kor97]. Conceptual [Egg94a, RA98, Kam95, Rao96]. conditions [Rip97]. conducted [BK90, Nun99]. Conference [Bar91b, BB99, Kre92, Van96, Le 99, MS94, Per97, Le 95]. Conferences [MS93]. Conforming [Bro96]. congestion [Egg99b]. Connection [KM90]. consequences [BK90, Hic99]. considerations [HH90]. consilience [Sma98a]. constraints [SBA90]. construct [Nag95b]. constructing [RR97]. construction [JW99, Moe96a]. consultancy [Cur99, Sch99a, Wil99a]. consumer [Kea91]. contact [KL94, TS94]. content [Lin96a, MCRC97, MCSAMSL99, QZ90, Tod92, Tro90]. contents [Ano93e, Seg96]. Context [FR98, McC98]. contexts [LW99]. continentalization [LG94]. continued [Abt94, Mar94]. continuous [Fuj98]. continuously [SW94]. contract [CR91]. contracts [Mau91]. contraints [SBNA90]. contributing [Kra94]. Contribution [Nag95a, AL91a, DWH99, KM90, Kre93, Lew93, Pic93, SA91, Tho95, Vin93b]. contributions [TM99]. contributors [SW94]. convictions [And99]. cooperate [Her95]. Cooperation [CFKS92, GSC99a, GFS99, LG94, NS94, Rus95, Ste98]. Cooperative [FAM^{+}92]. cooperativeness [Vin93b]. coordinates [Ada92]. Coping [GBFM96]. core [BBS99, Chu94, DCL^{+}96, vR96]. corpora [Hei95]. corporate [Ano93c, DM93]. Corpus [DL97]. corrected [SRMP99]. Correlation [Hal92, Hal93, EK95]. cosmic [SB96b]. cost [Raj96a]. Council [FAM^{+}92, GS96, MS92, RMP^{+}98]. count [Ley90b, Ley90c]. counting [HW99]. Countries [OW97, BBS97, BS96, Can95, Dic96, EDM92, Eto91, Gai92, GW92, God98, GSS98, HS98, Hel95, KW90, KRI99, Kra92, Kra94, LL93, MC92, MO94, MOO^{+}95, NS95, NB95, OMFD92, Pes92, PA91, PMR96, Rab92, San92, SGB90, SB92, Uzu96, Van94]. country [Egg99a, GI96, GH95, OW97, Pfe90, Tei94]. country-by-country [OW97]. counts [Gio93, Sam95]. coupling [GC96]. courses [MS93]. Courtial [Ley92b]. coverage [GSMF99, Hic99]. covered [FU92, GSSC99]. cow [BI97]. coword [CLCL93, Cou94]. Creating [Sma97a, Sma98b]. creation [Win96]. Crisis [LW94, Pou94, lP94]. Criteria [Ale90, Buc95]. critical [MVR96, Mar92]. critique [LW94]. Croatia [Kla95]. Croatian [Gol98]. Cross [Bha97, NP93, SB96a, Hin99, Kos97]. cross-disciplinarity [Hin99]. Cross-field [SB96a, Kos97]. Cross-national [Bha97, NP93]. Crowfoot [KKJ99]. crystals [SA91]. CSIR [GKA99, Lem92a, SRMP99]. Cuba [MF90]. Cuban [SBG93]. culture [YPC^{+}93]. cultures [GOR97]. cumulative [WD95]. currency [BNR90]. Current [AB97, GP99b, Ano96a, Bre96a, Bre96b, Sch90a, Sch90b, Sch91, Sch95]. cycle [MB94]. cycles [CL91, WD98, Wou97].

D [Kuh95, O’H95]. Dacca [Qur91]. Danish [And99]. Data [RR97, DLV99, Fau90, GHL92, God98, HQD92, LMRR92, MS94, PN92, QT90, RZ92, SQRD95, TLG94]. Database [KETP97, MDV95, ARR96, BB96b, GSSC99, Hai94, JSBCF91, JW99, Mau91, RM94, Spa97]. databases [Art96, BBG^{+}95, GSMF99, HW99, JSBCF91, Pla98, TV95, Whi92, Whi93, Wil99b]. datafiles [BMS92, BGS93, BS97, SGB90]. death [Ano97b]. Debrecen [BG91]. DEC [BB96c]. deciles [Cza97]. decision [Fas99a, HQD92, Tup99, WD94]. decisions [For92, HH06]. decline [Abt94, Ley91, Mar94]. deep [Ano97b]. defence [MG99]. defined [SAE95].

Defining [Wil98, RA93]. definition [Lew99]. definitional [Col93]. degrees [Tri91]. delays [LM98]. Delimitation [DM93, RDM93]. delineation [LC93]. democracy [RM94]. demographic [JD94]. demonstrated [Ste93]. department [CM99, Her93, NN92, DCL^{+}96, Kuh95]. departmental [ZMM91]. Departments [Cun95, Zac91, Ale90, CR98, CDC95, De 96, NMMvR93, Qur93, RH90, SBA90, Qur91]. departments/agencies [CR98]. Dependence [Qur93]. dependencies [WD97]. Derek [Ano93a, Ano95, Ano97a, Ano99, Bon95, Le 95, Le 99, Pav97, Sha90, Sma97b, Tod90, Bon94b]. derived [Art96]. describe [Sig96]. describing [CCL91, MS93]. Description [For92, MDV95]. descriptor [RBBMJCC99b]. design [Bon91, Per94b, Rou93]. Designing [Kam95]. Desktop [KH97a]. DESY [DV97]. detection [LOK99, ZB94]. Determinants [BS98, Prp96a, UH97]. determination [CIW97]. Determining [DCL^{+}96, Die92]. Deutsche [HN90]. developed [San92]. Developing [Lin91b, Eto91, Gai92, Kra92, NB95, Rab92, SB92, Shr97]. Development [NM93, NM95, ZB94, Ada92, Air99, Bla92, Bon96, CM90, Far99, GK95, GH95, Kam95, KW90, MF90, O’H95, RM96, RP92, Sen93, Smi95, Tho92a, Van94, Vin96b, WDB94, DG97]. developments [Coz95, GK95, NvR94]. devices [Bla97]. difference [Fuj98]. Differences [HF99, Moe96a, Rus98, MP98, Rou99a, Vin91b]. different [Bon90a, Bon90b, Cza97, HS98, HW99, Vin91b]. differentiation [RDC^{+}96]. difficulties [vCMvdB96]. difficulty [Hic99]. diffusion [Kor97]. digital [Cun98]. Dimensions [BS97, Sma98b]. direct [TW98]. disadvantage [dHLR94]. disciplinarity [Hin99]. Disciplinary [Col94, Boo94b, KRI99, MOO^{+}95, RDC^{+}96]. discipline [Hin94, Mat94, Pic93, Tho95]. discipline-contribution [Pic93, Tho95]. disciplines [HF99, MM90, OW97, RGK95, RRMS99, SS94, Ura90, VGMP^{+}94, WDB94]. discovery [Cam96, Chu92a, Chu92b, WD99]. discussion [Hor99, Luu97]. discussions [Egg94a]. disease [BI97]. diseases [Hin99]. Displaying [Tod92]. dispute [KM90, Kre93]. dissertation [RM97]. dissertations [BB96c]. Distribution [LMHZYYS96, PS91, POV91, Bur91, Chu94, Gup97, HW99, Kun90, LMRR92, Lem92a, LL93, Per90a, PN92, WD94, WD95]. distributions [Col93, GS93, GKR98, SR99a, WD97]. diversity [GBFM96]. d’Ivoire [CA92]. Do [Aru98, Bar94, BB99, Vla94, WSW90]. doctoral [Bre96a, GLMM92, Hog90, Tri91]. doctorate [BB96c, BB96c]. document [Laz94, Sma97a]. Does [Her95, LLD90, LF99, Avk97]. doing [BM91]. domain [DWH99]. Domestic [BGF^{+}96, HS98, Men92, RRMS99]. Dominancy [Vin93a]. Dorothy [KKJ99]. Double [Rou94a]. downloaded [HQD92]. Dr [Le 95]. dream [WL94]. dreams [Lyo90]. droop [QT90]. Drs. [Le 99]. drug [RV94, RKV95]. dual [Pic95]. duplicate [HW99]. duration [GK97, WD95]. during [LS95, MLH^{+}98]. Dutch [De 97, Rey92, VT93a, Van97b]. Dutch-speaking [Rey92]. dwarfs [Van97a]. dynamic [LC93, Mil94, SW99, Zit91]. dynamics [BWD96, Bla97, DQH91, LCS^{+}91, Mau91, RKV95, RBBMJCC99a, RBBMJCC99b, SG91].

each [Lew96]. Early [Fau90, MLH^{+}98]. earth [RRMS99]. east [GW92, BG96, CH93, GH94, LBR92, ODOM98]. Eastern [PMR96, KRI99, RA98, Uzu96]. EC [Bob91, DL91, GSK91, Lar95, Lew91, Tei91, Tsi91]. Ecclesiastes [HH93]. ecology [BDT97, Nor90, RM96]. econometrics [Egg92, Rou94b]. Economic [Lin96b, Pou92, BCML^{+}95, Bla92, CdSM93, DV92, Eto99b, Lew93, Van94, WD98]. economic/regional [Eto99b]. economics [BP98, Per90b, RH90, SBA90, SBNA90]. Economies [Pou93, RA98]. economists [Van97a]. Economy [DCL^{+}96]. editor [Bon94a, HH06, Per91]. Editorial [Bra92, Bra98, Ano97b, Ano97b]. editorship [LF99]. education [MC92, RWG^{+}97, Wes99]. educational [FCB99]. Effect [ZB96, LD98, Wil98, BBS97]. effective [FQD97]. effectiveness [LOK99, RR98]. effectivity [MHZ98]. Effects [So98, Tsi91, BCML^{+}95, BZRB99, DL91, HH93, Lar95, MLH^{+}98, Tho95]. efficiency [DAD99, GH94]. efforts [NV97]. Egghe [Bur01]. eighties [BGMS94b, BGMS94a, NS95, Sch95]. elite [Gol98, dAVA99]. EMBL [eKOM^{+}98]. emerging [Bar91b, BH91, Cun98, Hin94, McC98]. eminent [Prp96a, Prp96b, Prp98]. Empirical [Lin96a, Ley92a, Nie96, SSS99, SC98]. energy [Cze90, Mau91, VT93a]. engineering [BGG95a, HS98, MG93, PvR94, RS99, RM97, Zum97]. engineers [WW93, dS92]. Engines [Bri99, Fas99b, Obo99, Ska99, VB99, BI98b]. English [JCFA96, Lau95, RC94]. English-language [Lau95]. entities [Phi92]. entitled [Abt94]. entropy [Gru90a, Mat94]. envelopment [RR97]. environment [CS98, MCRC97, MM96b, Sig96, dS92]. Environmental [Kar90]. Epistemometria [Kra94]. EPO [NO98]. EPO/US [NO98]. eponymy [Tho92a]. Equalization [PM96]. era [Luu94]. Erdos [BI98a, BI98b]. Erratum [Gab96]. estimated [BI98b]. Estimates [MM90]. estimation [NT96]. ethics [Prp98]. ethnic [LI99]. études [SB92]. EU [Fay99, Led99]. Euram [BCML^{+}95]. Europe [KRI99, NSW91, RA98, WD98]. European [Bar91b, BCML^{+}95, Bar91b, Bar97, Bor99, BG96, BS96, BB96c, Cun97, Dav99, DLV99, GW92, GSC99a, Hig91, Jas92, Kne99, Kuk99, LBR92, Lew93, Lew94, Lis99, Man99, Mas91, Mau91, MDNT91, Pes92, PMR96, Rin95, Roj99, RR98, Str99, Whi93, eKOM^{+}98]. Europeanization [CR91]. Européen [BB96c]. evaluate [Cza97, vCMvdB96]. evaluated [Bal96]. Evaluating [Kam95, Smi95, TR91, Lin96b, Mau91, So98]. Evaluation [BCML^{+}95, CZ98, Dav99, Hel95, Hel99, Kne99, Pap96, SS92, Tei91, Tsi99, BG91, Bob91, BB96a, Bra99, CDC95, Cun95, Cun97, DF90b, Esc99, Fas99a, Fuj98, GS96, Gei96, Geo99, GK95, GLMM92, Hil95, Hor99, Kos98, Kru95, Kru99, Kuh95, Kyr95, LC91, Lew98, MS96b, Mas91, Mey99, MOO^{+}95, Nar95, NHHA94, NR95, O’H95, Pap99, PS91, Pra99, Rin95, RWG^{+}97, Sei99, Sir99, Tho92b, Tup99, Vin95, Wes99, Zac91, vH99, vR96, vR99]. evaluations [BK90, Geo99, Hor99, PMA90, Pra99]. Evidence [Row99, WD98]. evolon [PMA90]. Evolution [MADO98, BES90, GSS93, RT97, SR99a, ZB99]. evolutionary [Sch98]. exact [Egg91]. example [CM99, DCL^{+}96, Ste93, Tho95]. exceedances [Hub99]. exercises [vR96]. existence [WD98]. expansion [DG97]. expenditures [Tei94]. Experience [Rin95, Str99, Cun95, Hil95, Kat96, Kru95, Kuh95, ZT96, Zum97]. Experiences [ARR96, Roj99, Van97b]. experimental [SB96b]. Expert [Tij93, AL91a, RM97, So98]. expert-based [RM97]. experts [Cur99, Sch99a, Wil99a]. explain [Rao98]. explanation [Egg99a]. explanatory [Nie96]. exploration [Gup93, RM97]. exploratory [MS93]. exponential [Rou94a, Sch97b, SH98]. express [Kra94]. extent [GD97]. external [NM95]. eyes [BI98a, Hem93].

F [Ano95, Ano99, Le 95]. factor [Agu96, CIW97, MS96b, Nie96, RT95, Vog97]. Factors [Sen93, DR96, LOK99, Mea91, Mic95, MVR96, Prp96b, VGMP^{+}94, Vin99]. Facts [Rey92, Tod94, GS92]. faculties [Hog90]. faculty [Her93, KVA95, BG91]. Fair [JCFA96]. Fairness [Bec95]. fast [HB92]. fault [vR98]. favoured [Lew91, Lew93]. features [Vin94b]. federal [SR90, DF90a, Fic90, Kos95, RH90]. fellow [Tei91]. fellowship [Tei91]. female [GKA99]. field [AL91a, BH91, CM90, Cze90, EK95, FR98, GSS93, Kos97, LM98, MC92, MC94, MS94, McC98, MG92, Moe96b, NS95, NvR94, Pla98, RMP^{+}98, RM96, RA93, SPC92, San99, SW94, SB96a]. fields [Bar91a, Bar92, BZ97, BGMS94b, BGG95a, Cza97, DV97, DWH99, MS96b, Prp96b, RV94, Rub92, SR99a, YBT91]. fifteen [Van94]. figure [Gei96]. figures [GS92, Rey92]. file [KM97]. Filho [PD98]. finance [Avk97]. firms [CMS91]. first [DEP97, MDV95, Moe96a, Rou94a, Win96, Yam91]. first-citation [Rou94a]. fisheries [DS95]. five [HS98, Pes92]. fixation [dL94]. Flanders [DLV99]. Flemish [LMRR92, MLH^{+}98]. flow [Ste90, TLG94]. flows [God98, Ura90]. focusing [Fuj98]. food [De 98, HG96b, SR99b]. forecasting [CCS93, TS94]. foreign [CFKS92, UMÖ93]. foresight [vR96]. Forest [Le 95]. Foreword [Bra91, Bra94]. form [Wil99b]. formal [Chu92a]. Former [RA98, Dic96]. Forschungsgemeinschaft [HN90]. foundation [CCT93]. Four [RG98]. Fourth [Van96, Kre92]. Fractal [Egg97, vR91]. fraction [Egg99a]. fractional [Egg99a]. fractionation [NM93]. Framework [Fay99, Led99, Kam95, Sma98b]. frameworks [Prp94]. franca [RC94]. France [DCL^{+}96, Tur91]. French [De 96, Lec98, lMD91]. frequencies [Lew96]. frequency [Lem92a, RZ92, Vin98a, WD94, WD95]. FRG [Gie90]. friction [Rou98]. frontier [Bha97]. fronts [DJD98, ZB94]. fuels [Hal92]. full [For92, Hic99, VMP98]. fullerenes [Gup99]. function [Bro96, MG91]. functions [Bor99, ERR95, Gru90b, Lis99, Man99]. fundamental [SQRD95]. Fundamentals [RBBMJCC99a]. funded [Kea91, Per90b]. funding [JGSK98, LD98, Lin91b, MLH^{+}98, Pao91, Pao92, PEZ98, Ste93]. funds [Hel99, Tsi99]. future [Cun97, Rus94].

G [Oku93]. Gaining [LF99]. gaps [Egg94a]. GDP [HS98]. General [Vin98b, GSC99b, Mea91, NT96, Sma98b]. generalising [TV95]. Generalized [Egg92]. generally [Ley90b, Ley90c]. generation [Ley94b, RDQ^{+}98]. genesis [GH94]. genetics [GK96, Gup97, GKK97, GKR98, GK98, Gup98, KG98]. genome [BM91]. Geode [TLG94]. Geographical [Ano93c, Kat94]. geographically [RRMS99]. geographically-oriented [RRMS99]. geological [RC94]. geometry [vR91]. geophysics [Gup93, TW90]. geopolitical [BMS92]. German [CH93, DF90b, Kuh95, WD94]. Germany [BG90b, DF90a, Fic90, GW92, GH94, RH90]. getting [Laz94]. giants [ASR94, Van97a]. Gini [Bur91]. Glänzel [Ano99, Le 99, Mar91]. global [ER96, Kuk99, TV98, UMÖ93]. Goffman [WD99]. golden [Luu94]. gone [WD95]. good [Kea91]. Government [Cun95, Kea91, Kuh95, Sen93, Bro96, Kos96, Lin96b, PEZ98, SR90]. Government-funded [Kea91]. government-sponsored [Kos96, Lin96b]. graduate [GH95]. graduates [Bre96a]. Grail [Wou99]. grant [Gil91]. graph [Ley91]. gravity [Laz94]. Greece [Qur93, Tsi91, Zac91]. Griffith [Ano97a, Sma97b]. Groos [QT90]. gross [HS98]. group [Bur95, LMHZYYS96, Plo94]. group-Weibull [LMHZYYS96]. grouping [KW90]. groups [CS98, Coh91, Eng91, Mak98, Qur91, Qur93, dHLR94]. groups-relation [Mak98]. growing [SW94]. Growth [GSK97, Gup98, GSK99, SR99b, AL91b, CCC94, ER92, Egg93, GSK95, Hal92, Hal93, HS98, Men96, PM96, PMA90, RM96, Sch97b, SH98, WD98]. Guardians [Bec95]. guidelines [RMP^{+}98].

hard [BNR90, WL94]. hard-currency [BNR90]. Hargens [Lin91a]. harmonic [GS93]. heading [TW90, Tod92]. Health [lMD91, MC94, MC95, MCRC97, Nun99, PE99, dA92, dAVA99]. held [Le 95, Le 99]. Henk [Ano99, Le 99]. hero [GS91a]. Herting [Lin91a]. heuristic [Per94b]. Hierarchical [eKOM^{+}98]. high [Bal96, Chu92a, Chu92b, Cze90, NV99]. high- [Bal96]. high-energy [Cze90]. high-T [Chu92b]. high-Tc [Chu92a]. higher [Wes99]. Highly [MG91, SW99, GS92, Plo90, Plo94]. Historical [CLCL93]. histories [EK95]. histortical [RG98]. History [KRI99, Mar92]. Hodgkin [KKJ99]. Holy [Wou99]. homogeneity [Col93]. Human [LLD98, SN90, BM91, FVWdM97, GH95]. humanities [Fin90, HG96a, Yit97]. Hungarian [Vin98b]. Hungary [BG91]. hypertext [Egg97]. Hyphenation [BBG^{+}95]. hypothesis [Kre93, Eto91, KM90].

Iberian [GSS98]. IBSS [KM97]. ICT [TV98]. Identification [Pic93, CC91, Fau90, MB94]. Identifying [NV99, Pou91, CCS93, DRC^{+}99]. if [MZ92]. ignored [Luu97]. II [Bon90b, BB92, BGG95b, Chu92b, GD97, MO94, RBBMJCC99b, Sch97c]. ILL [vB92]. Illinois [Le 95]. illustrated [Rou93]. Impact [DV97, DR92, JGSK98, Mel99, Vog97, Agu96, Bal96, BK90, Bon91, BBS97, BGMS94b, BGMS94a, CIW97, Cza97, DR96, ER96, Fay99, GH95, Her93, IW99, Joh95, Kos95, KETP97, Led99, MS96b, Mic95, MVR96, MLH^{+}98, MVR99, Ove90, Per90b, Per91, SH96, SST98, SRMP99, Tri96, Tsa95, VGMP^{+}94, VMP98, Van98, Vin91a, Vin91b, Vin96a, Vin97, Vin99]. implementation [BGS90]. implications [Spa90, TW98]. importance [FVWdM97, GH95]. imposed [Tod94]. improving [ZB96]. inaccuracies [MVR96]. incentives [SBA90, SBNA90]. incidence [Cam96]. income [Gil91, Tei94]. incomplete [QT90]. increase [LF99, YG99]. increased [ERR95]. Increasing [Agu96, Rus95]. Index [BBS93, Win96, Ano93b, Ano93c, Ano93d, Bur91, Dah99, Luw99, NB95, SL92, So90, Vin97, Agu96, JSBCF91, Ley94b, BBS99]. indexes [Sch95, Vin98b]. India [Bas99, CR98, Dha98, JG92, Kar90, KG95, Lem92a, RG98, SR99b, Sik97, SA91]. Indian [GSK99, KG99, Kun96, KK99, Nag95a, Nag99, RGK95, Raj96b, RS99, SL92, SRMP99, TM99]. indicator [BS97, Hai92a, Hai92b, MO94, NT96, OMFD92, PS91, Plo90, Plo94, Wou99, Yit98]. indicator-latent [Hai92a, Hai92b]. Indicators [Ban98, Bar91b, BZ99, BS96, CC91, Hig91, RDC^{+}96, Sch97a, Art96, BGS90, Bar97, Bas99, Bha97, Bla92, Bob91, BMS92, BGS93, BBG^{+}95, Bra99, CJ92, CdSM93, CA92, Cze90, DF90a, DHS99, Fra91, GFS99, Gru90b, GSK91, GSM90, Hal92, Hal93, KW90, Kra92, Ley90a, Ley90b, Ley90c, LV97, Lin91b, MB94, Mar96, Mas91, MS92, MGB93, MM96b, MDV95, Moe96a, Moe96b, MVR99, Nar95, NMMvR93, NM95, RR97, SBG93, SR90, Sch93, SGB90, SG91, SB96a, Sir99, SOM91, Sni95, So90, Tho92b, TM96, Van97b, Vin91a, Vin93a, Vin94c, Vin96a, Vin96c, Vin97, Vin98b, Wes99, Zac91, ZB96, ZT96, dMR99, Bar91b, Van96]. Indices [PF95]. Individual [Mil94, CM99, Fer98, So98, Bon90b]. individuals [Vin95]. INDSCAL [RKV95]. Industrial [KG95, Air99, De 96, Far99, Nar95, Pap96, RM97, Smi95, Tka99, Tsa95]. Industry [KG95, BCML^{+}95, De 98, Hal92, Hal93, Smi95]. infancy [MZ92]. Influence [CH93, GSMF99, AL91a, Egg93, ERR95, LLD90, Van98]. influences [LBR92]. Info [Tur94]. infometrics [Tur94]. informal [Chu92b]. informatics [Ste93]. Information [Bri99, BH91, Fas99b, GLMM92, MGCG96, Obo99, Rou92, Ska99, Ura90, VB99, dS92, Ada92, CD97, DJD98, Eto94, Ley90a, PFK98, QBR^{+}98, Row99, SN90, Ste90, TV98, Tsa95, TLG94, Vin91b, Vin94b, WW93, vR91]. Information-based [GLMM92]. informativeness [TS94]. informativity [Yit97]. informed [LV97]. informetric [DWH99, Egg91, Egg97, Kor97, Mil94, Wil98, Wil99b]. Informetrics [Kre92, Egg92, Egg94a, Rou92, Rou94b, Rus94]. infrastructure [Can95]. inherent [Tho92b]. inherited [KRI99]. initial [Lew96]. Initiatives [Spi90]. Innovation [Bri99, Fas99b, Obo99, Sir99, Ska99, VB99, Wes99, Ban98, CMS91, Gru90a, JSBCF91, NR95, Smi95, dS92]. innovations [RDC^{+}96, SAE95]. innovative [CMS91, De 98]. Inorganic [Qur91]. input [Hei95, TR91]. input/output [Hei95, TR91]. INSERM [lMD91]. INSPEC [Dha98]. Institut [DCL^{+}96]. Institute [Kla90, lMD91, BZ99, De 96, UPS97]. institutes [NV99, Vin98b]. Institutional [TW98, GP99a, GC96]. institutionalization [RG98]. institutionalized [Kre94a]. Institutions [SCQ93, FAM^{+}92, NB95]. instituto [PD98]. Integrated [Gei96]. Integrating [NMV99]. integration [PM96]. intelligence [Bra99, KETP97]. Intensity [MHZ98]. inter [NS95]. inter-field [NS95]. interacting [PB95]. interaction [CCC94]. interactions [CCL91, Fas99a, Tup99]. interactivity [Kre90]. interchange [BNR90]. Interdependencies [Fas99a, Tup99]. interdisciplinarity [MP98, TM96]. Interdisciplinary [RDC^{+}96, RDQ^{+}98, BZRB99, Hin94, McC98, Mic95, RWG^{+}97]. interest [Eto99a]. interface [Hal92, Hal93]. interim [NO98]. interlibrary [Rey92, vB92]. International [ASR94, Bar91b, BGF^{+}96, BGMS92, BG96, EDM92, GW92, GH94, IM89, KM97, Kre92, KL94, LG94, MDNT91, NBFM92, SB90, Ste98, Van96, Bon94c, CIW97, DG97, DR92, DS95, GD97, GSC99a, GFS99, GSMF99, Hic99, IW99, LFJK93, LTPS93, MJ93, Men92, MO94, NB95, NN92, OMFD92, Oku93, Rab92, Rus95, Rus98, Tei94, Van98, Vin93a, Vog97, Whi92, Spa97]. Internationalization [ZB98, ZB99]. Internet [BI98a, BB96c]. interpretation [Bas99]. interpretations [Luu97]. Introduction [DF90a, Moe99, Van96, WL97, DRC^{+}99, Tur91]. invention [CCS93]. Inventive [Hub99]. investigated [QT90]. Investigation [Kun96, GH94, Qin94, SQRD95, Vin98a]. investment [Lin91b]. invisible [Kre94a, Kre97]. Invited [Mep95]. involvement [WW93]. iron [SPC92]. irregularity [WDB94]. Irreversibilites [Ley92a]. Irvine [Ano97a, Kea91, Pav97]. ISI [MC95, TV95]. isoc [RM94]. Israel [Cza97]. Israeli [CFKS92]. ISSI [Le 95, Le 99]. issue [Bra91]. issues [BB96b, Phe99, RA98, Rou93]. Italian [CMS91, GS96]. item [GSC99b, GSSC99, TS94]. item-by-item [GSC99b, GSSC99].

J [Ano95, GS91a, Le 95, Oku93]. J. [Kea91]. Jan [Tod90]. Japan [GK95, SA91, Ura90, Yam94]. Japanese [Lec98, Yam95]. japonais [Lec98]. JETRO [QZ90]. John [Ano93a, Ano95, Ano97a, Ano99, Bon94b, Ano97a, Pav97]. joint [Rin95]. Journal [Bon90a, Bon90b, MS96a, Ano97b, Ark99, BB99, CIW97, DR96, FU92, HH90, KETP97, Kre90, LC93, Ley94b, McC98, MVR96, MVR99, Pic95, PF95, SRMP99, TV95, WD97, Yit94, GP99b, MVR96, YBT91]. Journals [FU92, Bal96, BZ99, BP98, BBS99, Bur95, Cam95a, Cam95b, CDC95, CH93, DEP97, DCL^{+}96, ER96, GSC99b, GSSC99, GOR97, GSMF99, JD94, Lau95, LC93, MZ92, MS96a, MG91, MS96b, Mic95, NM95, O’N98, OD99, OPMU92, Ove90, Pic93, PF95, Rab92, RRMS99, Rou99a, SSS99, Sch99b, Sig96, Tod92, UMÖ93, Ves95, Vin91b, Vin99, Yam95, Yit97, ZB98, ZB99]. judgment [So98]. June [Le 95]. jurisdiction [WD94].

Karachi [Qur91]. keeping [BG96]. key [CCC94, Hai96, Men96]. key-word [CCC94]. keyword [Ano93d]. keywords [SQRD95]. Kingdom [SA91]. knowledge [CCC94, Fuj98, God98, HF99, NM95, Sch93]. Korea [LOK99]. Kossuth [BG91].

laboratories [MM96b, SRMP99]. laboratory [eKOM^{+}98]. Lajos [BG91]. Lancet [Per91]. landscape [BS91b, Kuk99]. landscapes [Sch98]. Language [Rub92, BY99, ERY99, Lau95, Yit98]. languages [CH93, Ste90]. large [Sma97a, Sma98b, SQRD95, Tri96]. large-scale [Sma98b]. Laser [JG92, GP98, GP99a, GP99b]. last [Men96]. latent [Hai92a, Hai92b]. Latin [Can95, FAM^{+}92, GSMF99, KVKWD95, LFJK93, MC92, MC94, MCRCNB98, MCSAMSL99, NBFM92, NB95, Ves95, Zum97]. latina [DHS99]. Latour [Luu97]. lattice [Cze90]. launched [MZ92]. laureate [KKJ99]. law [QBR^{+}98, Rao98, SH98, Boo94b, CZ98, Egg91, GSS93, GK99, RBBMJCC99b, SR99a, Nor90]. laws [BI97, Egg91]. lead [Avk97, DLV99]. leadership [SS94]. leading [CDC95]. learned [Kat96]. legal [WD94]. lending [Rey92, vB92]. length [Yit94]. less [Lew91, Lew93, San92]. lessons [Kat96]. Letter [Bon94a, Lew96]. letters [Per91]. level [BZ99, BSS97, BB98, Bor99, Dav99, Dic96, FR98, Fay99, Gio93, GC96, GI96, Kne99, Lis99, Man99, Roj99, Sma93, Str99, Tsi99, Wil98]. level-science [Tsi99]. lexical [Zit91, ZB94]. Leximappe [Ley92c, Cou92]. Leydesdorff’s’a [Cou92]. libraries [Cun98, LMRR92]. library [DJD98, Egg99b]. life [BIP99, CL91, MB94, MG91, Mat94, Pes92, Yam94, Yam95, lP94, BGG95b]. line [NM93, Ste93]. lingua [RC94]. link [FQD97, MO94, OMFD92, Bar94]. Linkage [NO98]. linkages [Nag99, RDC^{+}96]. linking [RA98]. links [GD97]. liquid [SA91]. lists [Gai92]. literature [AL91a, AL91b, Bra99, Chu94, CB95, DR92, FR98, GS91b, GP98, GP99b, GS94b, GSK97, Gup98, GSK99, HS98, Hic99, HB92, KETP97, LBR92, LM98, MCRCNB98, MCSAMSL99, Mat94, Nag95a, NHHA94, RS99, RDQ^{+}98, San99, SR99b, SK91, SA91, SRMP99, SH98]. Little [Egg94b, GS94a, Gri94, McG94, Mea94, Miq94, Per94a, Rao94, Sch94b, vR94, vR98]. Local [BGF^{+}96]. Locating [PB95]. logic [BWD96, WDB94, WD97, WD99]. lognormal [Kun90]. Long [Gio93, Tri91, Vin99, WD98]. Long-term [Tri91]. longing [Luu94]. longitudinal [McC98]. look [Bar94]. Looking [PD98]. lost [Luu94]. Lotka [Die92, GK99, SR99a]. low [Bal96]. low-impact [Bal96]. Luukkonen [Oku93].

Macro [Sma93, Bar91a, Bar92, BSS97, RGK95, ZT96]. macro-indicators [ZT96]. Macro-level [Sma93]. macro-strategic [Bar91a, Bar92]. Macroeconomic [Kyr95]. mad [BI97]. made [For92]. Magic [Vin91a]. main [CR98]. Mainstream [Zum97, Nag95a, Pic93, Spa90]. major [BMS92, CR98, De 96, LMHZYYS96, Whi93]. making [Dav99, Fas99a, Kne99, Tup99]. Malaysia [NHHA94]. male [GKA99]. mammal [Bir97]. Mammals [BB99]. management [Dav99, FR98, Kne99]. manifested [Vin94a]. manpower [Ste93]. manuscript [Abt94, Lin91a]. map [Sma94, Sma94]. Mapping [BB98, Bla97, HG96b, LW92, NV97, RKV95, Spa97, TW90, DB95, NV98, NMV99, Sma97a, TV95, CLCL93]. maps [Ley94b, NV97, Sma98b, Tij93]. margin [RV94]. mariage [Lec98]. Marine [BB99, Bir97]. market [CR91]. marriage [Lec98]. Marseille [DQH91]. Martin [Ano97a, BGS91, Kea91, Ley91, Pav97]. mastitis [RWG^{+}97, RDQ^{+}98]. mastitis-related [RWG^{+}97, RDQ^{+}98]. material [Mea91]. materials [Dic96, SPC92]. Mathematical [Egg98, NZ98, BWD96, BDT97, CM90, Eto99b, RT95, RT97, WDB94, WD99]. mathematician [BI98a]. mathematics [BGG95a, Eto99b, Moe96b, WD97]. matrices [Bur95]. matters [KRI99, vR98]. Matthew [BBS97, BBS99]. Max [Kru95]. meaning [Kra94]. means [GS93, Kra92]. measure [BY99, Hog90, Kre90, Lin91b, Mat94, NB95, NM93, Ric91, TS94, WSW90]. Measurement [DL91, Kre93, NM95, Bon96, BW97, Kru99, LTPS93, Oku93, Pfe90, San92, SCQ93, Urb96, ZB98]. Measures [ERY99, Yit98, DRC^{+}99, Eng91, Kos96, Kre90, Lin96b, Nag95b, NH96, Pic95]. Measuring [BZRB99, Fay99, Fin90, GHL92, God98, Her93, Led99, MG99, PN92, RA98, RDM93, RP92, Rou93, SH96, Van97a, VMP98, ARR96, DG97, Egg99b, GD97]. mechanics [FU92]. Medical [lMD91, Bre96a, GOR97, Hai96, Kun96, Lew98, LS91, MC92, OD99, Ste93, Yam91]. medicinal [GSM90]. Medicine [Hai94, IW99, KK99, MM90]. medium [BNR90]. Medline [Hai94]. Meeting [Nal91, Smi95, BB99, God98]. member [Van94]. members [Her93]. memorial [Bra91]. men [Lem92b, Ove90]. mental [Tij93]. merit [Gei96]. Merton [Ano95, Bon95]. Meta [Geo99, Hor99, Pra99]. Meta-evaluation [Geo99, Hor99, Pra99]. Metatheorizing [Cro98]. method [HQD92, LW92, Lin96a, Pic93, SS92, Ste90, VMP98, Zit91, ZB94]. Methodological [NT96, Rao96, GC96, RA98]. Methodologies [NR95, Kra94]. methodology [Cun95, Kuh95]. methods [Air99, Cze90, Far99, HW99, KM90, MS93, NN92, NZ98, Sch97c, Tka99, ZB96, vR96, vR99]. metric [Hai92a, Hai92b]. metricians [Laz94]. metrics [Phi92, Tur94]. Mexican [MCRC97, Rus98, dA92, dAVA99]. Mexico [ARR96, DR92, Le 99, MC95, Rus95]. Michael [Bra91, GS91a, SB92]. micro [BB98, Gio93, RGK95]. micro-level [Gio93]. mid [lP94]. mid-life [lP94]. Middle [Uzu96]. might [Ley90b, Ley90c]. Mike [TW91]. Misjudgments [San92]. misuse [Kos98]. Model [Vin94a, Vin96a, CM90, CCC94, DG97, Egg99b, GS94b, JG92, KH97b, Kre99, KK99, PMA90, Rao98, RT97, RBBMJCC99a, RBBMJCC99b, SH98, Vin96b, Vin98a]. Modeling [HB92, NM93]. Modelling [DG97, Fra91, GD97, GSK95, TS94, GSK97, Hal93, RT95]. Models [SG91, BES90, CZ98, ER92, Hai92a, Hai92b, Hal92, Mil94, Rou94a, vCMvdB96]. modern [BDT97, TM96]. Moed [Ano99, Le 99]. molecular [Her95, eKOM^{+}98]. monitoring [QZ90]. Monotonous [Eng91]. Monte [Cze90]. Moravcsik [Bra91, GS91a, SB92, TW91]. most [BP98]. Motivations [BS91a]. motives [Vin98a]. move [MCGC95]. movement [RT95]. mover [DEP97]. much [KH97b, vR98]. multi [BZ99, Boo94b]. multi-disciplinary [Boo94b]. multi-level [BZ99]. Multidimensional [GS91a, BWD96, Bon94b, NS94]. multidisciplinary [BZRB99, GSC99b, RA93, VT93a]. multilateral [DG97, GD97]. multinational [Egg99a, NM93]. Multinationality [Lew94]. multinationally [NSW91]. multiple [Bri91, Mar96, Pla98]. multipliers [Rao98]. multivariate [OMFD92, ODOM98]. Muslim [AB97]. mutual [Yit98].

N.C.I. [Qur93]. Nalimov [Ano97b]. names [Tod94]. Nanoscience [BSZ97]. Nanotechnology [MP98]. Nanotechnology-interdisciplinarity [MP98]. nanotecnology [BSZ97]. National [DCL^{+}96, FAM^{+}92, GS96, lMD91, ARR96, Bha97, BB96b, BS91b, Fay99, Fra91, GC96, GOR97, LC91, Mel99, MDV95, NP93, NB95, NM93, NV97, NT96, Pap99, Rab92, RDM93, Roj99, RP92, Rus98, Sei99, Str99, BGMS94b, BGMS94a, DR92, Glä96a]. nationality [Lew96]. nations [BGG95a, BGG95b, RR98]. Natural [BG91]. nature [Ark99]. Nebulae [SW94]. need [Aru98, Glä96b]. needs [Bar94]. negative [BH95, LMRR92, Lin96a]. Neglected [HH90, SOM91]. net [VT93b]. network [BH95, Bon90a, CCL91, DL91, PB95, PFK98, Tij93, TR91, TLG94, VT93b]. networking [Mel96]. networks [Cam95a, Cam95b, DB95, Fuj98, GFS99, Kre94a, Kre97, Ley92a, McC98, RBBMJCC99a, RBBMJCC99b, Sch98, Sig96, eKOM^{+}98]. Neural [McC98, Cam95a, Cam95b, PFK98, Tij93, VT93b]. News [Ano96c, Far90a, Far90b, San96]. Newsgroups [BI97]. Nigeria [Ada92]. Nigerian [Ehi90]. Nitrogen [dL94]. Nobel [KKJ99, SL93]. non [Chu92a, Chu92b, JCFA96, MC95, Mau91, Tri91, WW93, MC94]. non-Chinese [Chu92a, Chu92b]. non-English-speaking [JCFA96]. non-ISI [MC95]. non-nuclear [Mau91]. non-prestigious [Tri91]. Non-SCI [MC94]. non-specialist [WW93]. nontextual [SOM91]. Nordic [Hel95, PMDK97]. normalization [Kos97, SB96a]. North [Sko99]. note [MG99]. Notion [Laz94, VMP98]. nuclear [Hal92, Mau91, VT93a]. nuclei [Mil94]. Number [YBT91, AL91b, GKR98, MM90, Plo90, Yit94].

O [Oku93]. objects [McG96]. observations [Sni95, WD95]. observed [Hai96]. Obsolescence [CB95, ERR95, San99, Egg93, Gup98]. occasion [Le 95, Le 99]. occurrence [Lew96]. oceanography [TM99]. offering [vR98]. old [Vin98c]. on-line [NM93, Ste93]. one [Ano97b, Fer98, Ley90b, Ley90c, Luu94, Ley91]. Online [CI96, CIW97, Wil99b]. only [MVR96, Rab92]. opened [MCGC95]. Opening [Gab95, Bun99c, Ein99b, Esc99, Per97, Sca99]. Openness [So90]. operation [MDNT91, NSW91, Van94]. Ophthalmology [DWH99]. opinions [Urb96]. opportunities [Boo94a]. optical [Tri96]. optics [DWH99]. Optimizing [TLG94]. option [VMP98]. oranges [JJ92]. organic [KG99]. organisation [Van94]. organization [Bon96, DEP97]. Organizational [BGS90, CS98]. orientation [GH94]. oriented [RRMS99]. origin [Vin94b, Yam91]. originating [YG99]. ORTEGA [Kre93, Eto91, KM90]. Ost [ZT96]. other [DV97, Eto99b, Pes92, YBT91]. otorhinolaryngological [HN93]. our [Ano97b]. outcome [CR98]. outline [WDB99]. output [Bob91, BG90a, BGMS94b, BGMS94a, BS96, DR96, Her93, Kru99, Lew91, LFJK93, NS94, Nar95, NM93, QAG95, RDM93, SS92, Uzu90, WD97, Zac91]. outputs [Lew93, LD98]. outreach [RWG^{+}97]. overall [MVR96, NT96]. overlap [BI98b, PN92]. overseas [QZ90]. Overview [Kos96, MDV95, TW91]. Own [BY99, ERY99, Yit98]. Own-language [BY99, ERY99, Yit98].

paper [Bur01, HF99]. Papers [Ano96b, Ano96a, BB99, Bre96b, Bri91, EK95, GS92, GSC99b, GSSC99, GSK95, Kre92, Lem92a, Lew96, LS95, NSW91, NO98, NZ98, Plo90, Plo94, Rus98, Sik91, SOM91, TV98, Yit97]. paradigm [Kos97]. parameter [Bon90a]. parameters [Bon90a, Bon90b, Die92]. parametrization [RGK95]. Pareto [Egg91, LL93]. part [MO94, BWD96, Bon90a, Bon90b, BB92, BGMS94b, BGG95a, BGG95b, DG97, GD97, GSS98, Hai92a, Hai92b, Kre99, KK99, RBBMJCC99a, RBBMJCC99b]. Partial [dA92]. Participation [ASR94, GK97, NB95, WD95]. particle [SB96b]. particles [DV97]. particularities [Tro90]. partitioning [SC98]. partners [Ste98]. Pascal [JSBCF91]. past [RA98]. Patent [Kut98, Nar94, Opp97, Vin93a, CCS93, DLV99, Fau90, Fra91, GP98, Raj96b, Sch93]. Patent-related [Kut98]. patenting [Vin93a]. Patents [GSK91, Nar95, Tei94, Gup99, Lem92a, NO98, Vin94b]. pattern [CR98]. Patterns [DS95, Kre97, MC92, Row99, Whi93, Bir97, BBS99, BGMS92, CD97, De 97, FCB99, GK96, HF99, KM97, KG98, KK99, Lem92b, LFJK93, MP98, MOO^{+}95, NP93, Nor93, O’N98, Prp96b, Rus98, SA91, UÖ96]. Paul [BI98a]. Paulo [PEZ98]. pay [Her95]. PD [Vin93a]. Peer [HN90, Roj99, Str99, Bec95, BB99, TW98, Wou97, vR96]. peer-review [Bec95]. peer-reviewed [BB99]. peers [AL91a]. [Kyr95]. 20 [CCT93]. agencies [CR98]. output [Hei95, TR91]. regional [Eto99b]. STIMULATION [DL91]. US [NO98]. per-capita [Tei94]. Percentage [Vin93a]. perception [DRC^{+}99]. perceptions [UH97]. Perestroika [MGCG96]. perfectly [Rou93]. Performance [Kos96, Ale90, ARR96, BG91, BG90a, BGS91, Bro96, CS98, CDC95, DF90b, Fin90, Gil91, Hai96, IM89, Kra92, KL94, Lin91b, Lin96b, Mar91, MGB93, MDV95, Nag95b, NH96, NMMvR93, NV99, Nie96, NMV99, Pes92, Plo94, SPC92, SBA90, SBNA90, Vin98b, dAVA99]. performances [BS91b, BGMS94b, BGMS94a, Glä96a, Vin95]. period [Kla95, MGCG96, MOO^{+}95, PE99, PMR96]. periodical [Can95]. periodicals [Can95, SL92]. periods [RKV95]. peripheral [Can95]. Permuted [Ano93d]. persistent [Fuj98]. Personal [Kat96]. personnel [For92, RT95]. perspective [Bar97, Chu92a, Chu92b, GSK97, KVKWD95]. perspectives [MDNT91, Van97b]. Persson [Oku93]. petroleum [AD92, Hal93]. Ph.D. [KV97]. Pharmaceutical [Spa97, Vin94b]. pharmacologists [BGJB92, BB92]. pharmacology [MG92]. pharmacy [MG92, Spa97]. phase [MB94]. philology [Bon93]. philosophical [Qin94]. physical [Ehi90, MG91]. physicians [MM90]. Physics [Kea92, WDB99, Bha97, BSS97, BGMS92, BGMS94b, BGG95a, Cze90, DV97, Dha98, GSK99, NS94, RGK95, RG98, RDM93, SB96b, SQRD95, TW91, Uzu90, UMÖ93, Uzu96, Vog97, Dha98]. Pinski [Kre90]. Pitfalls [Gil91]. place [Egg91, LLD90, LBR92]. Planck [Kru95]. Planetary [SW94]. planning [Eto99b]. plant [DL97, Pla98, Tho92b, dL94]. plants [DRC^{+}99, GSM90]. point [Wes99]. points [Hor99]. policy [BK90, Dav99, Kne99, RMP^{+}98, RG99, Row99, SR90, Sig96, Spa90, TW98, Tsi91, Van97b]. polish [Ste98, Bon93, Win96]. Political [And99, Cur99, Lin96a, Sch99a, Wil99a]. politics [CDC95, LI99]. polluted [Rou98]. polymer [CCL91]. Pontificia [KVA95]. Popular [Kyv94, AL91a]. population [GK96, Gup97, GKK97, GKR98, GK98, Gup98, KG98]. portfolio [NV97]. portrait [KKJ99, KM95]. position [LF99]. Possibilities [KW90]. possibility [Glä97]. Possible [PMA90, Vin91b]. post [GH95, KRI99, Tsa95]. post-Communist [KRI99]. post-graduate [GH95]. post-industrial [Tsa95]. potential [GSK91, Kra94, MS96a, MF90, NR95, Pfe90]. power [GS91b, GSS93]. PPR [Vin93a]. practical [Vin96c]. practice [Gar98]. practices [Abt92]. Precision [Lin91a, BI98b]. prediction [Ley90a, WD99]. predictions [Glä97]. Preface [GKMS96]. preference [BY99, ERY99, Yit98]. preferred [Yam95]. presence [RBBMJCC99b]. present [DQH91, RA98]. presentation [Le 95, Le 99, Mea91]. Prest [Hil95]. prestigious [Tri91]. Price [Ano95, Ano97a, Ano99, Bon94b, Bon95, Le 95, Le 99, Pav97, Sha90, Sma97b, Tod90, Ano93a, Ste90, WL94]. Prices [Mic95]. Primary [MC95]. Principal [BDT97]. principles [Bon91, Egg92, Win96]. priorities [BSS97, NS95]. Prize [SL93]. probability [GKR98, PN92]. problem [Bri91, GBFM96, Hai92a, Hai92b, JJ92]. Problems [MM96a, MDNT91, MG99, RDM93, Tho92b]. Proceedings [Van96]. proces [Ban98]. process [Lin91a, SAE95]. processes [Cur99, Fas99a, Glä97, Rou92, Rou94a, SPC92, Sch99a, Tup99, Wil99a]. processing [SN90]. produced [MCSAMSL99, SRMP99]. producer [Kea91]. producers [Moe96a]. producing [Sch97c]. Product [SAE95, HS98]. production [CJ92, DAD99, ERR95, Gai92, GSS98, HF99, HG96a, MC94, MCRC97, Men92, Nun99, Pes92, Rou92, Rub92]. Productivity [Ehi90, GK97, Lem92b, Mak98, Tri96, BS98, Coh91, FVWdM97, GP99a, GK96, GK99, GKA99, Hai96, Hub99, Kla95, LLD98, MJ93, MLH^{+}98, MWHG93, MM90, POV91, Prp94, Prp96a, Prp96b, RT95, RT97, Shr97, WD95]. productivity- [RT97]. prof. [Le 95]. professional [Gol98, Prp98]. professions [LCS^{+}91]. Professor [Ano97b]. professors [Plo94, Yam91]. professorship [For92]. Profile [dAVA99, AD92, GKK97, KVA95, LD96, PvR94, Sik97]. profiles [Ale90, KV97, MO94]. Profiling [NV99]. prognosis [For92]. program [GH95, Mau91]. Programme [Led99, Bob91, BZRB99, CC91, Cun95, Kru99, Kuh95, Kyr95, Lew94, Tei91, Fay99, Tsi91]. programmes [BCML^{+}95, Cun97, DL91, GSK91, Lar95]. programming [Eto99b]. programs [Hog90, Kam95, KV97, PS91, RWG^{+}97]. project [CR98]. projections [JG92]. projects [CR98, FAM^{+}92, Kam95]. prolific [Qur91]. prolificacy [Plo90]. prominence [So98]. Promoting [Bon94c]. promotion [Spi90]. properties [CIW97]. proposals [LOK99]. proposition [Raj96a]. proteins [DL97]. proximity [Kat94]. Pseudo [Phi92]. Pseudo-metrics [Phi92]. psychiatric [NZ98, Nie96]. psychology [Kea92, LCS^{+}91, Ove90, San99]. Puberty [lP94]. public [Gei96, LMRR92, Mea91, Urb96]. Publication [Abt92, BB96a, IW99, Luu90, LM98, Qur91, SG91, SA91, UÖ96, Bas99, BG90a, BGMS94b, BGMS94a, CH93, DS95, Gai92, GSM90, GK97, HN93, KM97, KM90, LLD90, LBR92, LL93, MJ93, Men92, MOO^{+}95, MWHG93, NM93, ODOM98, Qur93, SR99a, SBG93, TW91, UPS97, Uzu90, Vin97, WD95, ZB98]. Publication-rate [Qur91, Qur93]. Publications [CDC95, Lew94, Rab92, Raj96a, BGJB92, BB92, BDT97, Can95, DM93, Egg99a, GBL95, Her93, JJ92, Kar90, MADO98, NM95, OW97, Pao91, Pao92, PS91, Per94b, RC94, SH96, SB96b, Spa90, Ste93, UMÖ93, Uzu96, Vin94a, YG99, eKOM^{+}98, lMD91]. publish [Rab92]. published [Ano96a, Bre96b, DR92, GSC99b, GSSC99, MVR96, NHHA94, Ove90]. publishers [Ano97b]. Publishing [JCFA96, Rus98, Can95, DS91, Kyv94, Luu92, RMP^{+}98, Zum97].

Qualitative [Rip97]. Quality [Bun99b, Eto94, Lin99, Nyb99, Pom99, Avk97, Bri91, Buc95, GLMM92, Hem93, HHM^{+}96, Kre94b, SCQ93, UPS97, Van97a]. Quantification [Seg96, Joh95]. quantified [SAE95]. Quantifying [vCMvdB96]. quantitation [Bra99]. Quantitative [Air99, DF90a, Far99, MS94, Mea91, MK99, SR90, Sch90a, Sch90b, Sch91, Sch95, Tka99, Urb96, KM97, Lin96a, Nag95b, RS99, RBBMJCC99a, SBG93, Uzu90, Vin96a, WDB99, vR96, vR98]. Quantity [Kre94b]. Quasi [Nag95b]. Quasi-quantitative [Nag95b]. query [BI98b]. question [Men92, RDQ^{+}98]. questionnaire [Hem93, Tei91]. Questions [Bas99, Rao96].

R [Ano97a, Oku93, Pav97]. R&D [Bob91, Bro96, CdSM93, CMS91, CR98, FR98, GS96, HG96b, Mas91, McC95, MM96b, MG99, NvR94, Pou92, Tei94, Tka99, Tsi91]. Raan [Ano95, Le 95]. radiation [Ste93]. rain [AL91a]. Random [Gar98]. rank [POV91].

Ranking [TM99, Bon90a, Bon90b, So98]. Rankings [Gie90, RH90, TW98]. rate [CH93, KM90, Qur91, Qur93, RDC^{+}96, Vin96b, YG99]. rates [BB96a, ER92, LF99, Rou99a, ZB96]. ratings [ZMM91]. Ratio [Vin99]. rays [BH95, SB96b]. re [Fuj98]. re-evaluation [Fuj98]. reactions [BMSV91, RV94, RKV95]. reality [WSW90]. reason [Lyo90]. Reassessing [BH95]. Reassessment [ZB96, Kea91]. recall [BI98b, ZB96]. recently [MZ92]. reception [Bon93]. Recherche [DCL^{+}96]. recipient [Bon95, Sha90, Sma97b, Tod90]. recipients [Pav97]. recognition [LF99]. recommendations [HH06]. reconsidered [SR99a]. records [HW99]. referee [HH06]. Refereeing [Yam95]. referees [HH90]. Reference [SB93, BH95, GSC99b, GSSC99, ODOM98, SA91, Vin98a]. referenced [Vin94b]. references [Vin99, YBT91]. referencing [Vin98a]. refinement [Pic95]. reflected [GK97, Sch93]. reflecting [Bon90a]. Reflections [BK90]. refractory [SPC92]. regard [GH94]. region [Lew93, Whi93]. regional [GC96, GI96, NR95, O’H95, Tsi99, TR91]. regions [BMS92, Lew91]. Regularity [WDB94]. rehabilitated [Kun90]. rejoinder [LW99]. related [Eto99a, EK95, Kut98, RWG^{+}97, RDQ^{+}98, Vin95]. relatedness [PF95]. Relation [Yit94, Egg99a, Ehi90, GSS93, Mak98, WW93]. Relations [Ley90b, Ley90c, Vin97, PMR96, Sch97a, Vin93a]. Relationship [Eto99b, HS98, DEP97, FVWD98, Hai92a, Hai92b, Lin91b, LM98, Nor90, Pao91, Pao92]. Relationships [Vin96b, De 96, FVWdM97, PF95, TR91]. relative [Abt94, BY99, ERY99, Mar94, Vin91a, Vin96a, Vin97, Yit98]. relevance [DL97, NV99, RG99, RM97, SSS99]. reliability [Bec95, Glä97, Nag95b]. remarks [Bun99a, Bun99c, Bur01, Ein99b, Esc99, Rou99b, Sca99]. repeatedly [Fuj98]. reply [BGS91, Ley91, Ley92b, Mar91]. Report [Abt94]. reports [ARR96, FU92, Nie96, MS96a]. Representation [Vin93a, TLG94]. represented [Vin99, vR91]. Representing [CA92]. Republic [DF90a, Fic90, RH90, Kla95, MF90]. Reputation [BNR90, Mak98]. requirements [Bro96]. Research [BGF^{+}96, Coz95, DF90b, DJD98, GS96, GK95, Hai96, HG96a, Joh95, MS92, NS94, NMMvR93, PMDK97, SST98, SCQ93, UPS97, Vin93b, Yam94, Zac91, lMD91, Air99, AD92, Avk97, BS98, BM91, Bar91a, Bar92, BZ99, Bha97, BSS97, BB98, Bla92, BZ97, BZRB99, BG96, BS96, CCL91, CLCL93, CS98, Coh91, CR91, CDC95, CM90, CC91, Cun98, Cza97, DAD99, DEP97, DV97, DCL^{+}96, De 96, Dha98, Dic96, DS95, DB95, Eto94, Eto99a, FR98, Far99, FCB99, FAM^{+}92, Fin90, GS91b, GSS93, GP99b, GHL92, Gei96, Gil91, Glä96a, Glä96b, Gru90a, GSK91, GH94, GSM90, Gup93, GSK97, HH97, HHM^{+}96, HN93, JG92, JGSK98, Jas92, Kam95, Kar90, KG95, KG99, Kla90, Kor97, Kos95, Kos96, KETP97, Kos98, Kre90, Kun96, KL94, Lau95, LW92]. research [LCS^{+}91, LOK99, Lem92a, LC91, Lew93, LD98, Lin91b, Lin96b, MJ93, Mak98, Mar96, McC98, MP96, Mel99, MG92, MGB93, MDV95, MLH^{+}98, Nag95b, Nar95, NN92, NM95, NV97, NV99, Nie96, NMV99, Nun99, O’N98, OW97, Pao91, Pao92, Pap96, Per90b, Pla98, Plo94, Qin94, Qur91, RG98, RMP^{+}98, Rin95, RWG^{+}97, RM96, RG99, RP92, Rus95, SPC92, SR90, SW99, SS92, Sen93, Sig96, Sma94, Smi95, So98, Ste93, Syl93, Tho92b, TW98, Tij93, TV98, Tro90, TR91, VT93a, VT93b, Van98, Vin98b, Yit97, YPC^{+}93, ZMM91, ZB94, dA92, dAVA99, dHLR94, vCMvdB96]. researchers [And99, CFKS92, DCL^{+}96, Kla95, Ste98]. resources [GH95, LLD98]. respectability [Kea92]. response [RMP^{+}98]. restructuring [KRI99]. result [BB99]. Results [Bar91b, Bro96, Van98, Win96, vCMvdB96]. retrieval [Ada92, QBR^{+}98]. return [Lin91b]. Review [Sch94a, Bec95, Glä90, HN90, LOK99, Ley94a, Lin96a, Lin91a, Mar92, Nun99, Roj99, Rou90, Str99, TW98, Wou97, vR96]. review- [Str99]. reviewed [BB99]. Reviews [Sig96]. revisited [IM89, Lin91a, RZ92, SL93, WD99]. revolution [Eto91, MADO98]. reward [Luu92]. Rise [RF92, Lyo90]. risk [DRC^{+}99, UH97]. River [Le 95]. roadmaps [KETP97]. Robert [Ano95, Bon95]. Role [Cur99, Sch99a, Wil99a, AL91a, BH95, Esc99, Men96, Mey99, Pap99, RRMS99, Rus95, Sei99, vH99]. ROM [Ley94b]. roots [Gol98]. RT [Kuh95, O’H95]. RT&D [Cun95, Lar95]. RTD [Kam95]. Rugjer [Kla90]. rule [CCT93]. Rural [DCL^{+}96]. Russian [GOR97, MGCG96].

S [Bar91b, KM95, Bar91b, Kyr95]. S&T [Bar97, NT96, RA98, Van97b]. S/ [Kyr95]. safety [CLCL93]. samples [Rou93]. Sampling [Rou93, Die92, PN92]. sanctions [Pou92]. sanitation [PN92]. São [PEZ98]. scale [Pou93, Sma98b, SQRD95]. Scandinavian [IW99]. scapegoats [Bar94]. scattering [Rao98, SC98]. scenario [PE99]. schema [Hei95]. scholar [GS91a]. scholarly [Ove90, SSS99]. scholars [So98]. schools [Bon94c, Lew98, Yam91]. Schubert [Ano93a, Mar91]. sci [Moe96a, BGJB92, BB92, LC93, MC94, MS96a, Sma94]. Science [Art96, Bas99, BBS93, DWH99, DHS99, Esc99, Eto91, KW90, Lec98, NS95, Prp98, Van96, ZT96, Abt94, And99, AB97, ASR94, Bec95, Bir97, BK90, BGMS94b, BGMS94a, BGG95a, BGG95b, Col93, Dah99, DF90a, DJD98, Egg94b, Eto94, FQD97, Fer98, GS91a, GP99a, Glä96a, Gru90b, GH95, Hal92, Hal93, Hel99, Hic99, JW99, Kar90, KH96, Kea91, KRI99, KVKWD95, Kre94b, Kru99, Kyv94, LG94, LD96, LI99, Ley90a, Ley90b, Ley90c, Ley91, Ley92a, Lin91b, LM98, Luw99, Lyo90, MS96b, Mar92, Mar94, Men96, MCGC95, Mic95, MO94, MOO^{+}95, Nag99, NT96, NV98, NMV99, OMFD92, PD98, PM96, Pic93, Pla98, Pou93, RA93, Rus95, Sch98, Sch93, Sch97a, Sch90a, Sch90b, Sch91, Sch95, Sch99b, SR99b, SL92, Sik97, Sir99, Sma97a, Sma98a, Sma98b]. science [SOM91, Spa90, SN90, Spi90, TV98, TM96, Tsi99, VT93a, Van97b, Wes99, Whi92, WL94, Wou98, Yam95, ZB96, dMR99, vR98, DL91, Agu96, JSBCF91, Ley94b]. science-industry [Hal92, Hal93]. science-part [MO94]. SCIENCE/STIMULATION [DL91]. Sciences [BG91, Rub92, Abt92, Bar94, BG90a, BS91b, BGG95b, BS96, CM99, DV92, GW92, GS92, Glä96a, HG96b, JGSK98, KM97, KVA95, Kun96, MG91, NV97, NV99, PEZ98, PE99, Pes92, Qin94, RRMS99, SB90, SH96, Tho95, Yam94, dA92, dAVA99, Vin98b]. Scientific [Avk97, BGF^{+}96, GI96, GOR97, GSS98, HH97, Hem93, JSBCF91, LL93, MADO98, MS92, NSW91, PMR96, Prp96b, QAG95, RG99, Rub92, Ada92, BZ99, BG91, BWD96, Bon91, Bon96, BG90b, BGS91, Bra99, Buc95, Cam95a, Cam95b, Cam96, CJ92, CA92, Coh91, CM99, CH93, DM93, DG97, DHS99, DQH91, EDM92, Eto99a, FVWdM97, Gai92, GHL92, GS94b, GD97, GSC99a, God98, Gol98, GFS99, GSK95, GSK97, GK97, Hai96, Hig91, IM89, JD94, Kat94, Kea91, Kla95, KVA95, KV97, Kre99, KK99, LF99, LG94, LLD98, Lew91, LFJK93, Lew96, LMHZYYS96, LTPS93, LM98, MC92, MC94, MS96b, MS93, Mar91, Mea91, MF90, Mil94, MDNT91, MM90, Nag95a, NBFM92, NM93, NvR94, Nun99, Oku93, OPMU92, PS91, Per94b, PB95, Pes92, PV91]. scientific [Pfe90, PA91, Plo90, Plo94, PFK98, POV91, Prp94, RC94, Ric91, RV94, RF92, RT95, RT97, RR98, Rou99a, RBBMJCC99a, RBBMJCC99b, SR99a, SMPK93, San92, SBG93, SH96, Seg96, SS92, Sik91, SOM91, SS94, SBA90, SBNA90, Tod92, Tro90, VGMP^{+}94, Ves95, Vin95, Vin96b, WDB94, WD95, WD99, Whi92, ZB98, ZB99, dS92]. scientific-technical [DHS99, dS92]. scientist [Hem93, Luu92, MWHG93, Prp96a, Prp96b, Prp98, Spa90]. Scientists [Vin98c, AL91a, Chu92a, Chu92b, DR92, Ehi90, Gai92, GKA99, KM90, LBR92, Lem92b, MGCG96, PvR94, Rab92, Rus98, Shr97, Yam95]. Scientometric [Ark99, BG91, BMS92, BGS93, Cze90, GP98, JG92, KKJ99, KM95, Kla90, Kra92, MB94, SGB90, Sch97b, Sch97c, Tro90, WD98, BBG^{+}95, BGG95a, BGG95b, BS97, CG97, FCB99, GS91b, GS93, KV97, Ley91, MS92, RR97, SB92, SB96a, Ste98, Tij93, TV95, Vin96a, Vin96c, Vin97, Vin98b, dL94]. scientometrical [PMA90]. Scientometrics [Boo94a, De 96, GP99a, GP99b, KG99, LV97, Sch99c, Van97c, Vla94, Bon94c, Bon96, BES90, CLCL93, Cou94, Cun97, GS94a, Gri94, GPK97, Gru90a, KH97a, Kra94, McG94, McG96, Mea94, Mén94, Miq94, Per94a, Pou94, Rao94, Rip97, Sch94b, Tur91, Vin94c, WL94, Zit91, vR94, Bon94b, Kre92, Sch96a, Sch96b, Sch96c, Sch96d]. Scientostochastics [Bur94]. SciViz [Sma98b]. scope [ZB98]. score [Egg99a, Pic93]. scores [Tho95]. sea [Eto99a]. sea-related [Eto99a]. search [BI98b]. Searching [RV94]. section [FR98]. sector [DF90b, GS96, Gei96, dL94]. sectors [GBL95]. Secular [Vin98d]. seed [YPC^{+}93]. seeking [Luu92]. seen [RM94]. selected [Glä96a, GKR98, Yit97]. selection [For92, Mea91, Vin96a]. selective [Bon90a]. self [BH95, Bon96, BS91a, BY99, ERY99, Pic95, Rou99a, Yit98]. self-citation [BY99, ERY99, Pic95, Rou99a, Yit98]. self-organization [Bon96]. sell [JCFA96]. Serbia [Kut98]. SERC [JGSK98]. serendipity [Cam96]. serial [AL91b]. serials [TM99]. series [DG97, Hal92, Hal93]. service [Eto94]. session [Ein99a]. set [ER96, LC93]. sets [QT90]. sex [Bre96a, Lem92a]. Shannon [Mat94]. shape [Col94]. shelves [Egg99b]. short [JCFA96, SSS99, Vin99]. shortcomings [MG99, San92]. should [Bal96]. significance [Plo90, SQRD95]. signs [Wou98]. Similarities [Rou94b, Vin99]. simple [Zit91]. simulation [CCC94]. since [DS91, Kea91]. single [Bri91].

Sivertsen [Oku93]. Size [Coh91, SMPK93, Tho95, AL91b, Jas91, Mel99, PA91, Qur91, Qur93, So98, Tei94]. SLAC [DV97]. small [GI96]. Social [Rub92, And99, Bar94, Col93, CM99, CCC94, Eng91, Glä96a, Hic99, IW99, KM97, Kre93, Lew93, Mic95, NV97, NV99, Row99, Tho95, WW93]. social-science [Col93]. Socialist [RA98]. societal [Bar94]. Society [Bri99, Fas99b, KETP97, Obo99, Ska99, VB99, NM95, Tsa95, MVR96]. Socio [Bla92, CdSM93, Gol98, Prp94]. socio-cognitive [Prp94]. Socio-economic [Bla92, CdSM93]. socio-professional [Gol98]. sociological [Luu97]. Sociology [DCL^{+}96, Win96, CG97, De 97, RG98, Yit98]. Solar [GS91b, GSS93]. Solla [Ano93a, Ano95, Ano97a, Ano99, Bon94b, Bon95, Le 95, Le 99, Pav97, Sha90, Sma97b, Tod90]. Some [Bur01, CJ92, GS92, MGB93, SW94, Sko99, Sni95, Spa90, SBA90, SBNA90, Tsi91, Vin95, Vin96c, BGS91, Hor99, Oku93, Per94b, Qur93, ZT96, WD95]. sorrow [Ano97b]. sources [LBR92, WW93]. South [ODOM98, Pou91]. South-East [ODOM98]. Southern [Tho92a]. Soviet [Dic96, MG91, SA91]. space [Bar91b, Bon94b, vR91]. spaces [Sma97a]. Spain [Bla92, DJD98, MG92, Pes92, PMR96]. span [BIP99]. Spanish [BGJB92, BB92, CJ92, FCB99, FAM^{+}92, JCFA96, Kra94, MS92, OPMU92, Pes92, RRMS99, RMP^{+}98, RM94, Rub92]. speaking [JCFA96, Kra94, Rey92]. special [ODOM98, SA91]. specialist [WW93]. specialities [GSK97, SW99]. speciality [GKK97, GK98, Pic93]. Specialization [PA91, Bar91a, Bar92, NP93]. specialties [Eto99b, LC93, PB95]. specialty [Cze90, GSK95, LS91, So98]. specific [BIP99]. spectra [Hal92, Hal93, VMP98]. Speech [Le 95, Le 99, Mep95]. speed [GK97, WD95]. spending [Sen93]. Sponsored [CR98, Kos96, Lin96b]. sponsoring [CR98]. SSCI [GSSC99]. stability [Agu96]. staff [DS95]. standard [MS96b]. standardization [Vin96c]. Standards [BB96b, Bre96a, Glä96b, Kat96, Rao96, SB93]. standing [NN92]. stands [Vin94c]. starts [MZ92]. State [Sen93, Van97c, AB97, Kra92, Ste93, OPMU92, PEZ98]. State-of-the-art [Van97c]. statements [Oku93]. Statistical [LS95, Sik97]. statistics [BMS92, Hub99, Raj96b, SW94, vB92]. status [UMÖ93]. steelmaking [SPC92]. stochastic [GS94b, Glä97]. strategic [Bar91a, Bar92, BMSV91, CC91, HQD92]. strategies [Gai92, Sch98, Ves95, YPC^{+}93]. Strategy [BBS93, Vin97]. stratification [KM90, Kre93]. stratified [Rou93]. strength [Pou91, Vin98a]. Strengths [MK99]. Stride [Tsi91]. Structural [Lar95, Bun99b, CA92, Hel99, Lin99, Nag99, Nyb99, Pom99, Tsi99]. Structure [MF90, MO94, OMFD92, RBBMJCC99a, RBBMJCC99b, Bon90a, Eng91, GFS99, KRI99, Kre90, McC95, MLH^{+}98, Row99, Sma93, SS94]. structured [RT97]. Structures [WD97, HG96b, Luu90, MS94]. Structuring [PV91]. students [FVWD98]. studies [BWD96, DF90a, DR92, Lau95, LC91, Lin96a, Nun99, Sch90a, Sch90b, Sch91, Sch95, Sch97b, Sch97c, vR98]. study [AL91a, BI98b, Cam95a, Cam95b, Cam96, Cou92, CH93, DWH99, DS91, DLV99, Dha98, DQH91, Gai92, GS91b, GSS93, GP98, GD97, GK99, GSS98, Hai94, Hem93, Her95, JSBCF91, JG92, JCFA96, Kor97, LCS^{+}91, Ley92c, MG91, MS94, McG96, Mea91, NMMvR93, NT96, PvR94, Prp98, RGK95, RS99, SSS99, SST98, SC98, Sma94, Ste98, Tho92a, Tho92b, TW98, Tij93, Tsa95, VT93a, YBT91, Yit97, dMR99, lMD91]. Studying [MP96, SS94, MS96a]. styles [HF99]. sub [RGK95]. sub-disciplines [RGK95]. subdisciplinary [SS94]. subfield [SRMP99]. subfields [BGMS92, DM93, Lew99, NS94, Vin91b]. Subject [GSM90, MCRC97, MCSAMSL99, GSC99b, GSSC99, GBFM96, MC94, Wil98, Wil99b]. Subject-based [GSM90]. subsequent [RKV95]. substances [Sch97b, Sch97c]. Substantiating [RG98]. suited [Bon90b]. Summary [BMS92]. Summing [dMR99]. superconductor [Chu92a]. superconductors [Chu92b]. superpower [BG90b]. Superstring [BH91, HB92]. supplementary [Gru90b, SGB90]. supply [PN92]. support [Lew94, Vin98c]. supported [Kra92]. Survey [De 98, CMS91, GHL92, Tei91, dS92]. Swedish [Ano96a, Bre96a, Bre96b]. switching [Bla97]. Synthesis [RDQ^{+}98]. Synthesizing [FCB99]. synthetic [VT93a]. system [Bec95, Bon91, Esc99, Fic90, Fuj98, GPK97, MB94, Sma98b, TLG94, Yam95]. systemic [KH96]. systems [Chu94, CD97, Egg97, Kre90, RWG^{+}97, RP92]. systems-based [RWG^{+}97].

T [Bar91b, Kyr95, Oku93, Bar91b, Chu92b]. T. [KM95]. Taking [RR98, RT95]. Tc [Chu92a]. teacher [Tod97]. team [Ano97a, BZ97, Vin93b]. Technical [DAD99, Coh91, DHS99, KETP97, OPMU92, PFK98, SST98, Whi92, dS92]. technique [FR98]. techniques [Lew98, OMFD92, RM97, TLG94]. Technological [Gup99, CCL91, CMS91, De 98, Fau90, Fra91, Kea91, LMHZYYS96, Mau91, MF90, Mil94, NvR94, Pfe90]. technologie [Lec98]. technologies [Urb96]. technologiges [TV98]. Technology [Van96, AB97, DF90a, DLV99, FQD97, GP99a, Glä96b, GK95, Gru90b, GH94, GH95, Hel99, Kam95, KW90, Kru99, Lec98, Ley92a, MB94, NT96, NV98, PFK98, Pou91, Pou93, QZ90, Rus95, SR90, Sch93, Sch97a, SR99b, Sik97, Sir99, Smi95, Tsi99, VT93a, Wes99, dMR99]. telephone [Bla97]. telescopes [Tri96]. Temporal [Rou99a, SC98]. ten [CL91, Men96]. term [Gio93, Kra94, Tri91, Vin99]. terminology [Laz96]. terms [AL91b, LC93, LMHZYYS96]. terrain [Spa97]. test [LW92]. tested [KM90]. text [Hei95, RM97]. texts [Ley90b, Ley90c, LW99]. Textual [SOM91]. Thailand [YPC^{+}93]. their [AL91a, GK97, Hal92, Hal93, Lew96, LM98, MLH^{+}98, Per94b, TM99, WW93]. thematic [MS94]. themes [CC91]. Theoretical [CCT93, Cze90, GK96, Gup97, GKK97, GKR98, GK98, Gup98, KG98, Van97b, Kre99]. theoretically [LV97]. Theories [Ley98, Egg98, LW99, Wou99]. theorists [vR98]. theory [Aru98, BH91, Cze90, Gar98, HB92, Ley90a, LV97, Luu97, O’N98, Rou92, Rou98, Vin98c, Wou99]. there [GBL95]. theses [Ano96a, Bre96a, Bre96b, GLMM92]. Thinking [Mén94]. Third [OW97, GS91a, KW90, Whi92]. thirteen [BP98]. thoughts [Gar98, Tsi91]. Three [SB92, GW92, Sma98b, Vin91a]. threshold [Vin98a]. throughout [YBT91]. Tijsen [Oku93]. Time [WD97, Hal92, Hal93, RKV95, SC98, TS94, Tod97]. tissue [YPC^{+}93]. title [Yit94]. titles [CCS93, MM90, Yit97]. together [Laz94]. tomography [KETP97]. too [vR98]. tool [CCL91, De 96, DB95, MS96a, MS96b, RR97]. tools [BDT97, MDV95, dL94]. top [PvR94]. top-scientists [PvR94]. topic [BIP99]. topical [Tro90]. topics [CCS93, Eto99a]. topographical [BG90a]. Toshiba [Kam95]. total [MM90]. Tracing [Sch93]. Tracking [BSS97]. trade [JCFA96, NM95]. Traditional [Hai94]. Training [GH95, Tei91]. trans [LC91]. trans-national [LC91]. transactions [Qin94]. transdisciplinary [HG96b]. transfer [BH91, Egg92, NM95, Sch93]. transgenic [DRC^{+}99]. transition [Esc99, Fer98, RM94]. translation [CG97, RBBMJCC99a]. Transnational [Nag99, NS94]. trend [Gup93, LMRR92, Vin98d, ZB94]. trendlines [BMS92]. Trends [Bob91, BDT97, CCS93, DB95, Gup99, HN93, Kun96, ODOM98]. trial [NT96]. trials [Per94b]. triangle [Vin91a]. tribute [SB92]. true [WL94]. Turkey [YG99]. Turkish [Uzu90, UMÖ93, UÖ96]. TWC [OW97]. twilight [RF92]. Two [Bon90a, DRC^{+}99, So90, BZ97, DEP97, De 96, Qur91, Sma98b, Zac91, vCMvdB96]. Type [Nie96]. types [BB96a]. typological [CS98]. Typology [OMFD92, NS95].

U.K. [Qur93]. U.S. [Coz95, RM97]. UK [BGS91, Can95, Cun95, Mar91]. uncitedness [SH91]. understanding [BMSV91]. unfunded [Per90b]. Union [Dic96, GSC99a, SA91]. unit [McG96]. United [BG90b, GH94, SA91]. Universidad [KVA95]. Universities [Bri99, Ehi90, Fas99b, Obo99, Ska99, VB99, Ale90, Bra99, Bun99b, Gie90, Hai96, Kra92, Lin99, MLH^{+}98, Nag95a, Nyb99, PD98, PMDK97, Pom99, Rub92, Tri91, vR99]. University [BG91, DR92, Qur91, Bal95, BZRB99, CDC95, DF90b, DR96, Fic90, Her93, KVA95, Mel96, NN92, NMMvR93, NVBG95, Qur93, SST98, Sni95, Tod97, Kuk99, Zac91]. unresolved [Bri91]. Update [Sma97a]. US-biased [Luw99]. USA [Qur93]. usage [Ano96a, Bre96b]. Use [Gai92, LBR92, NV97, Bob91, Bur95, CCS93, Cza97, ERR95, God98, HHM^{+}96, Kos98, Mar96, Mas91, Pla98, Sni95, Van97b, WW93, Yit98, Zac91, vB92]. used [QBR^{+}98, SPC92]. Usenet [BI97]. users [NM95]. Using [BMSV91, Cam95a, Cam95b, Cam96, Mau91, Wil99b, Bha97, BB98, DR96, DM93, DG97, DLV99, GSC99b, GSSC99, Gup99, Hei95, HW99, Ley90a, LC93, Lin96a, MP96, MO94, OMFD92, Pic93, PN92, RKV95, SBG93, SRMP99, TR91, Zit91]. utility [GSK91, NM95, Sik91]. utilization [ERR95, vCMvdB96]. utilizing [Ste93]. vaccines [GSS98]. Validation [Moe96b, Cou92, Ley92c]. validity [Nag95b, RDQ^{+}98]. value [Per94b, Ves95]. values [Prp98]. Variable [SH98, GK99]. Variables [MWHG93, Col94, Ehi90]. Variation [Yit97, Lew96]. various [Abt92, KM90, Moe96a]. varying [Col93, Wil98]. Vassiliyevich [Ano97b]. Vassily [Ano97b]. Venezuela [Lem92b]. Venezuelan [RF92]. version [Ley94b]. versus [BGMS94b, BGMS94a, Men92, O’N98, So98, Tij93]. veterinary [RWG^{+}97]. via [NM95, TW98]. View [Shr97, KH96, Moe96a, Tsi91, Wes99, ZB99]. viewed [GP99b, KG95]. vigorous [GH95]. vintage [Vin98c]. visibility [CM99, Jas92, MC94]. Visible [Shr97]. Vision [DWH99]. Vlachy [Tod90]. Volume [Ano93e]. vs [Tri91].

W [Oku93]. war [Yam91]. warnings [Sko99]. water [PN92]. weaknesses [MK99]. wealth [RR98]. web [BIP99, GPK97]. Weibull [LMHZYYS96]. weight [BGG95a, BGG95b]. Welcoming [Bun99c, Ein99b, Esc99, Sca99]. well [MZ92]. Were [WD95]. West [GH94, Jas92, KM95]. Which [Bon90b, GBL95, WSW90, Dou94]. Who [BM91]. Wide [GPK97, NV97]. Will [BG96]. William [WD99]. Win [Ano99, Ano95, Ano97a]. wins [Ano93a]. winter [Vin98c]. winter-biased [Vin98c]. wish [Ley90b, Ley90c]. within [BZRB99, CC91, Mak98, MDNT91]. without [MGB93]. witnesses [AL91a]. Wolfgang [Le 99, Ano99]. women [Lem92b, Ove90]. word [BB98, CCL91, CCC94, DL97, LW92, TR91]. Words [Vin94c, DM93, RZ92]. work [LCS^{+}91, Men96, MWHG93]. working [Tod97]. World [BGMS94b, BGMS94a, MOO^{+}95, OW97, AB97, BG90a, GS91a, GS91b, GSK99, KW90, NV97, SR99b, Shr97, Whi92, GPK97]. world-wide [NV97]. worth [Bal96, KH97b].

XXIth [Nal91].

year [Agu96]. Years [Ric91, CLCL93, Men96]. Yogoslav [LI99]. Yugoslavia [Kla90].

Zipf [Egg91, LL93, RZ92, RBBMJCC99b]. Zoology [Qur91].

References

Anwar:1997:CSS

[AB97] M. A. Anwar and A. B. Abu Bakar. Current state of science and technology in the Muslim world. Scientometrics, 40(1): 23–44, September 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

               com/article/10.1007/BF02459260.

Abt:1992:PPV

[Abt92] H. A. Abt. Publication practices in various sciences. Scientometrics, 24(3):441–447, July 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02051040.

Abt:1994:RME

[Abt94] H. Abt. Report on the manuscript entitled “British science in the 1980s — Has the relative decline continued?”. Scientometrics, 29(1):57–58, January 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02018383.

See [Mar94].

Anderson:1992:PCC

[AD92] F. Anderson and R. Dalpé. A profile of Canadian coal and petroleum research communities. Scientometrics, 25(3):447– 463, November 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016931.

Adamson:1992:ARI

[Ada92] I. Adamson. Access and retrieval of information as coordinates of scientific development and achievement in Nigeria. Scientometrics, 23(1):191–199, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020922.

Aguillo:1996:IBY

[Agu96] I. F. Aguillo. Increasing the between-year stability of the impact factor in the Science Citation Index. Scientometrics, 35(2):279–282, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018486.

Airaghi:1999:QMI

[Air99] A. Airaghi. Quantitative methods in industrial research and development. Scientometrics, 45(3):455–457, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457606.

Abdullah:1991:CSP

[AL91a] Szarina Bt. Abdullah and F. W. Lancaster. The contribution of scientists to the popular literature, their role as expert witnesses, and their influence on their peers: A case study in the field of acid rain. Scientometrics, 20(1):55–64, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018143.

Archibald:1991:SGS

[AL91b] G. Archibald and M. B. Line. The size and growth of serial literature 1950–1987, in terms of the number of articles per serial. Scientometrics, 20(1):173–196, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018154.

Alewell:1990:CPP

[Ale90] K. Alewell. Criteria for performance profiles of departments and universities. Scientometrics, 19(5–6):337–347, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020697.

Andersen:1999:PAC

[And99] Heine Andersen. Political attitudes and cognitive convictions among Danish social science researchers. Scientometrics, 46(1):87–108, September 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02766297.

Anonymous:1993:ASW

[Ano93a] Anonymous. András Schubert wins the 1993 Derek John de Solla Price award. Scientometrics, 28(3):233–235, November/ December 1993. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02026509.

Anonymous:1993:AI

[Ano93b] Anonymous. Author index. Scientometrics, 26(3):361–367, March 1993. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02091782.

Anonymous:1993:GCI

[Ano93c] Anonymous. Geographical & corporate index. Scientometrics, 26(3):367–376, March 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02091783.

Anonymous:1993:PKI

[Ano93d] Anonymous. Permuted keyword index. Scientometrics, 26(3): 376–484, March 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02091784.

Anonymous:1993:VC

[Ano93e] Anonymous. Volume contents. Scientometrics, 26(3):315–361, March 1993. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02091781.

Anonymous:1995:RKM

[Ano95] Anonymous. Robert K. Merton and Anthony F. J. van Raan win the 1995 Derek John de Solla Price Award. Scientometrics, 34(2):I–II, October 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02020417.

See comments [Bon95].

Anonymous:1996:AUP

[Ano96a] Anonymous. Authorship on and usage of published papers in current Swedish biomedical theses. Scientometrics, 37(2):383, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/BF02093632.pdf.

Anonymous:1996:CP

[Ano96b] Anonymous. Call for papers. Scientometrics, 36(3):471, July/ August 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02129606.

Anonymous:1996:N

[Ano96c] Anonymous. News. Scientometrics, 37(2):381–382, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02093631.

Anonymous:1997:BCG

[Ano97a] Anonymous. Belver C. Griffith and John Irvine and Ben R. Martin (as a team) win the 1997 Derek John de Solla Price Award. Scientometrics, 40(3):357, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02459283. See comments [Sma97b,

Pav97].

Anonymous:1997:EBP

[Ano97b] Anonymous. The Editorial Board and the publishers of our journal announce with deep sorrow the death of Professor Vassily Vassiliyevich Nalimov one of the editorial advisors of our journal. Scientometrics, 39(2):143–145, June 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457443.

Anonymous:1998:A

[Ano98] Anonymous. Announcement. Scientometrics, 41(1–2):273, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02457983.

Anonymous:1999:WGH

[Ano99] Anonymous. Wolfgang Glänzel and Henk F. Moed win the 1999 Derek John de Solla Price Award. Scientometrics, 46 (1):3–4, September 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02766290.

Arkhipov:1999:SAN

[Ark99] D. B. Arkhipov. Scientometric analysis of nature, the journal. Scientometrics, 46(1):51–72, September 1999. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02766295.

Arvanitis:1996:ENC

[ARR96] R. Arvanitis, J. M. Russell, and A. Ma. Rosas. Experiences with the national citation reports database for measuring national performance: The case of Mexico. Scientometrics, 35(2):247–255, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018482.

Artus:1996:SID

[Art96] H. M. Artus. Science indicators derived from databases. Scientometrics, 37(2):297–311, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02093626.

Arunachalam:1998:CAD

[Aru98] S. Arunachalam. Citation analysis: Do we need a theory? Scientometrics, 43(1):141–142, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02458402.

Arunachalam:1994:ICS

[ASR94] S. Arunachalam, R. Srinivasan, and V. Raman. International collaboration in science: Participation by the Asian giants. Scientometrics, 30(1):7–22, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02017209.

Avkiran:1997:SCF

[Avk97] N. K. Avkiran. Scientific collaboration in finance does not lead to better quality research. Scientometrics, 39(2):173– 184, June 1997. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02457446.

Balaban:1995:CAU

[Bal95] A. T. Balaban. Can the assignment of university chairs be automated? Scientometrics, 32(2):121–122, February 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02016889.

Balaban:1996:HSC

[Bal96] A. T. Balaban. How should citations to articles in highand low-impact journals be evaluated, or what is a citation worth? Scientometrics, 37(3):495–498, November 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019260.

Banerjee:1998:IIP

[Ban98] Parthasarathi Banerjee. Indicators of ‘innovation as a process’. Scientometrics, 43(3):331–357, November 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457403.

Barre:1991:CRF

[Bar91a] R. Barré. Clustering research fields for macro-strategic analysis: A comparative specialization approach. Scientometrics, 22(1):95–112, September 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02019277.

Barre:1991:IEE

[Bar91b] R. Barré. Indicators of the emerging European S & T space: Results of the 1st International Conference on European S & T Indicators. Scientometrics, 21(3):465–469, July/August 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093982.

Barre:1992:CRF

[Bar92] R. Barré. Clustering research fields for macro-strategic analysis: A comparitive specialization approach. Scientometrics, 23 (2):353, February 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/content/pdf/10.1007/BF02017046.pdf.

Barre:1994:DLS

[Bar94] R. Barre. Do not look for scapegoats! Link bibliometrics to social sciences and address societal needs. Scientometrics, 30(2– 3):419–424, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018116.

Barre:1997:EPS

[Bar97] R. Barré. The European perspective on S&T indicators. Scientometrics, 38(1):57–70, January 1997. CODEN SCNTDX.

ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02461123.

Basu:1999:SPI

[Bas99] Aparna Basu. Science publication indicators for India: Questions of interpretation. Scientometrics, 44(3):347–360, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458484.

Bordons:1992:BAPb

[BB92] Maria Bordons and S. Barrigón. Bibliometric analysis of publications of Spanish pharmacologists in the SCI (1984– 89). Part II. Scientometrics, 25(3):425–446, November 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016930.

Bourke:1996:PTC

[BB96a] P. Bourke and Linda Butler. Publication types, citation rates and evaluation. Scientometrics, 37(3):473–494, November 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019259.

Bourke:1996:SIN

[BB96b] P. Bourke and Linda Butler. Standards issues in a national bibliometric database: The Australian case. Scientometrics, 35(2):199–207, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018478.

Breimer:1996:CDC

[BB96c] L. H. Breimer and D. D. Breimer. The CED-le DEC: Common European doctorate, or doctorate Européen commune or dissertations on the Internet. Scientometrics, 35(3):347–353, March 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02016905.

Bhattacharya:1998:MRA

[BB98] Sujit Bhattacharya and Prajit K. Basu. Mapping a research area at the micro level using co-word analysis. Scientometrics, 43(3):359–372, November 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02457404.

Bird:1999:DPR

[BB99] J. E. Bird and M. D. Bird. Do peer-reviewed journal papers result from meeting abstracts of the Biennial Conference on the Biology of Marine Mammals? Scientometrics, 46(2): 287–297, October 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02464779.

Braun:1995:HDB

[BBG^{+}95] T. Braun, M. Brocken, W. Glänzel, E. Rinia, and A. Schubert. “Hyphenation” of databases in building scientometric indicators. Scientometrics, 33(2):131–148, June 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020565.

Bonitz:1993:SSI

[BBS93] M. Bonitz, E. Bruckner, and Andrea Scharnhorst. The science strategy index. Scientometrics, 26(1):37–50, January 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016791.

Bonitz:1997:CIM

[BBS97] M. Bonitz, E. Bruckner, and Andrea Scharnhorst. Characteristics and impact of the Matthew Effect for countries. Scientometrics, 40(3):407–422, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02459289.

Bonitz:1999:MIC

[BBS99] M. Bonitz, E. Bruckner, and Andrea Scharnhorst. The Matthew index — concentration patterns and Matthew core journals. Scientometrics, 44(3):361–378, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458485.

Bach:1995:EEE

[BCML^{+}95] L. Bach, Núria Conde-Molist, M.-J. Ledoux, Mireille Matt, and Véronique Schaeffer. Evaluation of the economic effects of Brite–Euram programmes on the European industry. Scientometrics, 34(3):325–349, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018003.

Budilova:1997:PTM

[BDT97] E. V. Budilova, J. A. Drogalina, and A. T. Teriokhin. Principal trends in modern ecology and its mathematical tools: An analysis of publications. Scientometrics, 39(2):147–157, June 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457444.

Beck:1995:GSF

[Bec95] M. T. Beck. Guardians of science. Fairness and reliability of peer-review system. Scientometrics, 32(1):91–92, January 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02020191.

Bruckner:1990:AEM

[BES90] E. Bruckner, W. Ebeling, and A. Scharnhorst. The application of evolution models in scientometrics. Scientometrics, 18(1– 2):21–41, January 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019160.

Braun:1990:TAW

[BG90a] T. Braun and W. Glänzel. A topographical approach to world publication output and performance in the sciences, 1981–1985. Scientometrics, 19(3–4):159–165, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02095344.

Braun:1990:UGN

[BG90b] T. Braun and W. Glänzel. United Germany: The new scientific superpower? Scientometrics, 19(5–6):513–521, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020712.

Beck:1991:SES

[BG91] M. T. Beck and V. Gáspár. Scientometric evaluation of the scientific performance at the Faculty of Natural Sciences, Kossuth Lajos University, Debrecen, Hungary. Scientometrics, 20 (1):37–54, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018142.

Braun:1996:ICW

[BG96] T. Braun and W. Glänzel. International collaboration: Will it be keeping alive East European research? Scientometrics, 36 (2):247–254, June 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017317.

Bordons:1996:LDI

[BGF^{+}96] Maria Bordons, Isabel Gómez, M. Teresa Fernández, M. Angeles Zulueta, and Aida Méndez. Local, domestic and international scientific collaboration in biomedical research. Scientometrics, 37(2):279–295, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02093625.

Braun:1995:SWNa

[BGG95a] T. Braun, W. Glänzel, and H. Grupp. The scientometric weight of 50 nations in 27 science areas, 1989–1993. Part I. All fields combined, mathematics, engineering, chemistry and physics. Scientometrics, 33(3):263–293, July 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017332.

Braun:1995:SWNb

[BGG95b] T. Braun, W. Glänzel, and H. Grupp. The scientometric weight of 50 nations in 27 science areas, 1989–1993. Part II. Life sciences. Scientometrics, 34(2):207–237, October 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020421.

Bordons:1992:BAPa

[BGJB92] M. Bordons, F. García-Jover, and S. Barrigon. Bibliometric analysis of publications of Spanish pharmacologists in the SCI (1984–89). Scientometrics, 24(1):163–177, May 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02026479.

Braun:1992:ICA

[BGMS92] T. Braun, I. Gómez, A. Méndez, and A. Schubert. International co-authorship patterns in physics and its subfields, 1981–1985. Scientometrics, 24(2):181–200, June 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017908.

Braun:1994:WSEb

[BGMS94a] T. Braun, W. Glänzel, Hajnalka Maczelka, and A. Schubert. World science in the eighties. National performances in publication output and citation impact, 1985–1989 versus 1980–1984. Scientometrics, 31(1):3–30, September 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018099.

Braun:1994:WSEa

[BGMS94b] T. Braun, W. Glänzel, Hajnalka Maczelka, and A. Schubert. World science in the eighties. National performances in publication output and citation impact, 1985–1989 versus 1980–1984. Part I. All science fields combined, physics, and chemistry. Scientometrics, 29(3):299–334, March 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02033443.

Backes-Gellner:1990:OIB

[BGS90] U. Backes-Gellner and D. Sadowski. Organizational implementation of bibliometric indicators. Scientometrics, 19(5–6): 385–395, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020701.

Braun:1991:BAU

[BGS91] T. Braun, W. Glänzel, and A. Schubert. The bibliometric assessment of UK scientific performance — some comments on Martin’s “reply”. Scientometrics, 20(2):359–362, February 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.

1007/BF02017525. See reply [Mar91].

Braun:1993:SID

[BGS93] T. Braun, W. Glänzel, and A. Schubert. Scientometric indicators datafiles. Scientometrics, 28(2):137–150, October 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016896.

Budd:1991:STI

[BH91] J. Budd and C. D. Hurt. Superstring theory: Information transfer in an emerging field. Scientometrics, 21(1):87–98, May 1991. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02019184.

Baldi:1995:RRR

[BH95] Stéphane Baldi and L. L. Hargens. Reassessing the Nrays reference network: The role of self citations and negative citations. Scientometrics, 34(2):239–253, October 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020422.

Bhattacharya:1997:CNC

[Bha97] S. Bhattacharya. Cross-national comparison of frontier areas of research in physics using bibliometric indicators. Scientometrics, 40(3):385–405, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02459288.

Bar-Ilan:1997:MCD

[BI97] Judit Bar-Ilan. The “mad cow disease”, Usenet newsgroups and bibliometric laws. Scientometrics, 39(1):29–55, May 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457429.

Bar-Ilan:1998:MPE

[BI98a] Judit Bar-Ilan. The mathematician, Paul Erdős (1913–1996) in the eyes of the Internet. Scientometrics, 43(2):257–267, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02458410.

Bar-Ilan:1998:OPE

[BI98b] Judit Bar-Ilan. On the overlap, the precision and estimated recall of search engines. A case study of the query “Erdős”. Scientometrics, 42(2):207–228, June 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458356.

Bar-Ilan:1999:LSS

[BIP99] Judit Bar-Ilan and Bluma C. Peritz. The life span of a specific topic on the web. Scientometrics, 46(3):371–382, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459598.

Bird:1997:APM

[Bir97] J. E. Bird. Authorship patterns in marine mammal science, 1985–1993. Scientometrics, 39(1):99–105, May 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457432.

Block:1990:WCR

[BK90] H.-J. Block and W. Krull. What are the consequences? Reflections on the impact of evaluations conducted by a science policy advisory body. Scientometrics, 19(5–6):427–437, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02020705.

Blasco:1992:SEI

[Bla92] P. González Blasco. Socio-economic indicators on research and development in Spain. Scientometrics, 24(1):79–93, May 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02026474.

Blauwhof:1997:MDT

[Bla97] Gertrud Blauwhof. Mapping the dynamics of telephone switching devices. Scientometrics, 38(1):117–140, January 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02461127.

Balmer:1991:WDW

[BM91] B. Balmer and B. R. Martin. Who’s doing what in human genome research? Scientometrics, 22(3):369–377, November 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019769.

Braun:1992:SID

[BMS92] T. Braun, Hajnalka Maczelka, and A. Schubert. Scientometric indicators datafiles. Summary statistics and trendlines of major geopolitical regions, 1980–1989. Scientometrics, 25(2): 211–217, October 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02028082.

Bauin:1991:UBS

[BMSV91] S. Bauin, B. Michelet, M. G. Schweighoffer, and P. Vermeulin. Using bibliometrics in strategic analysis: “understanding chemical reactions” at the CNRS. Scientometrics, 22(1):113–137, September 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02019278.

Baumert:1990:RHC

[BNR90] J. Baumert, J. Naumann, and P. M. Roeder. Reputation — a hard-currency medium of interchange. Scientometrics, 19 (5–6):397–408, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02020702.

Bobe:1991:TUR

[Bob91] B. Bobe. Trends in the use of R&D output indicators in EC programme evaluation. Scientometrics, 21(3):263–282, July/ August 1991. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02093970.

Bonitz:1990:JRDa

[Bon90a] M. Bonitz. Journal ranking by different parameter. Part I. Collectivity and selective collectivity: Two ranking parameters reflecting the structure of a journal network. Scientometrics, 18(1–2):57–73, January 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02019162.

Bonitz:1990:JRDb

[Bon90b] M. Bonitz. Journal ranking by different parameters. Part II. Individual or collective: Which parameters are best suited for journal ranking? Scientometrics, 18(1–2):75–93, January 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019163.

Bonitz:1991:IBP

[Bon91] M. Bonitz. The impact of behavioral principles on the design of the system of scientific communication. Scientometrics, 20(1): 107–111, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018148.

Bonheim:1993:RPP

[Bon93] H. Bonheim. The reception of Polish philology abroad. Scientometrics, 26(2):243–253, February 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016217.

Bonitz:1994:LE

[Bon94a] M. Bonitz. Letter to the editor. Scientometrics, 31(1):123, September 1994. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02018105.

Bonitz:1994:MSS

[Bon94b] M. Bonitz. The multidimensional space of Scientometrics; the Derek John de Solla Price Awards 1984–1993. Scientometrics, 29(1):3–14, January 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018380.

Bonitz:1994:PSI

[Bon94c] M. Bonitz. Promoting scientometrics by international schools. Scientometrics, 30(2–3):451–453, June 1994. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02018123.

Bonitz:1995:CRK

[Bon95] Manfred Bonitz. Comments on Robert K. Merton, recipient of the 1995 Derek de Solla Price Award. Scientometrics, 34(2):III–VI, October 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02020418.

See [Ano95].

Bonitz:1996:CSD

[Bon96] M. Bonitz. The challenge of scientometrics: The development, measurement and self-organization of scientific communications. Scientometrics, 36(2):271–272, June 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02017320.

Bookstein:1994:SNO

[Boo94a] A. Bookstein. Scientometrics: New opportunities. Scientometrics, 30(2–3):455–460, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018124.

Bookstein:1994:TMD

[Boo94b] A. Bookstein. Towards a multi-disciplinary Bradford Law. Scientometrics, 30(1):353–361, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02017233.

Borgman:1999:AFE

[Bor99] J. Borgman. Advisory functions at European level. Scientometrics, 45(3):485–486, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02457613.

Beckmann:1998:TMC

[BP98] M. Beckmann and O. Persson. The thirteen most cited journals in economics. Scientometrics, 42(2):267–271, June 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458360.

Braun:1991:FMM

[Bra91] T. Braun. Foreword to the Michael Moravcsik memorial issue. Scientometrics, 20(1):3–7, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018138.

Braun:1992:E

[Bra92] Tibor Braun. Editorial. Scientometrics, 23(2):265, February 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02017041.

Braun:1994:F

[Bra94] Tibor Braun. Foreword. Scientometrics, 30(2–3):373, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02018106.

Braun:1998:E

[Bra98] Tibor Braun. Editorial. Scientometrics, 43(1):3, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02458390.

Braun:1999:BIE

[Bra99] T. Braun. Bibliometric indicators for the evaluation of universities — intelligence from the quantitation of the scientific literature. Scientometrics, 45(3):425–432, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457602.

Breimer:1996:ASS

[Bre96a] L. H. Breimer. Age, sex and standards of current doctoral theses by Swedish medical graduates. Scientometrics, 37(1):171– 176, September 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02093493.

Breimer:1996:AUP

[Bre96b] L. H. Breimer. Authorship on and usage of published papers in current Swedish biomedical theses. Scientometrics, 36 (2):255–258, June 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017318.

Bridgstock:1991:QSM

[Bri91] M. Bridgstock. The quality of single and multiple authored papers; an unresolved problem. Scientometrics, 21(1):37– 48, May 1991. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02019181.

Bricall:1999:UEI

[Bri99] J. Bricall. Universities — engines of innovation in the information society. Scientometrics, 45(3):551–556, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457627.

Brown:1996:CGR

[Bro96] E. A. Brown. Conforming the government R&D function with the requirements of the government performance and results act. Scientometrics, 36(3):445–470, July/August 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02129605.

Bonzi:1991:MCC

[BS91a] Susan Bonzi and H. W. Snyder. Motivations for citation: A comparison of self citation and citation to others. Scientometrics, 21(2):245–254, June 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017571.

Braun:1991:LNP

[BS91b] T. Braun and A. Schubert. The landscape of national performances in the sciences, 1981–1985. Scientometrics, 20(1): 9–17, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018139.

Braun:1996:IRO

[BS96] T. Braun and A. Schubert. Indicators of research output in the sciences from 5 central European countries, 1990–1994.

Scientometrics, 36(2):145–165, June 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02017311.

Braun:1997:DSI

[BS97] T. Braun and A. Schubert. Dimensions of scientometric indicator datafiles. Scientometrics, 38(1):175–204, January 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02461130.

Babu:1998:DRP

[BS98] A. Ramesh Babu and Y. P. Singh. Determinants of research productivity. Scientometrics, 43(3):309–329, November 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457402.

Bhattacharya:1997:TCR

[BSS97] Sujit Bhattacharya, S. P. Singh, and P. Sudhakar. Tracking changes in research priorities in physics: A macro level analysis. Scientometrics, 40(1):57–82, September 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459262.

Braun:1997:NNB

[BSZ97] T. Braun, A. Schubert, and S. Zsindely. Nanoscience and nanotecnology on the balance. Scientometrics, 38(2):321–325, February 1997. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02457417.

Buchholz:1995:CAS

[Buc95] K. Buchholz. Criteria for the analysis of scientific quality. Scientometrics, 32(2):195–218, February 1995. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02016894.

Bundschuh:1999:CR

[Bun99a] E. Bundschuh. Closing remarks. Scientometrics, 45(3):567, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02457631.

Bundschuh:1999:QAS

[Bun99b] E. Bundschuh. Quality assessment and structural change in universities. Scientometrics, 45(3):359–365, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457590.

Bundschuh:1999:WOR

[Bun99c] E. Bundschuh. Welcoming and opening remarks. Scientometrics, 45(3):335–336, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02457584.

Burrell:1991:BDG

[Bur91] Q. L. Burrell. The Bradford distribution and the Gini index. Scientometrics, 21(2):181–194, June 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02017568.

Burrell:1994:S

[Bur94] Q. L. Burrell. Scientostochastics? Scientometrics, 30(2–3): 505–509, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018131.

Burton:1995:UCM

[Bur95] M. P. Burton. The use of citations matrices to group journals. Scientometrics, 33(2):257–262, June 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020572.

Burrell:2001:SRP

[Bur01] Quentin L. Burrell. Some remarks on a paper by Egghe. Scientometrics, 50(2):345–350, February 2001. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1010586026266.

See [Egg99a].

Bookstein:1997:AM

[BW97] A. Bookstein and B. Wright. Ambiguity in measurement. Scientometrics, 40(3):369–384, November 1997. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02459287.

Berg:1996:MAS

[BWD96] J. Berg and R. Wagner-Döbler. A multidimensional analysis of scientific dynamics. Part I. Case studies of mathematical logic in the 20th century. Scientometrics, 35(3):321–346, March 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016904.

Bookstein:1999:OLP

[BY99] A. Bookstein and M. Yitzhaki. Own-language preference: A new measure of “relative language self-citation”. Scientometrics, 46(2):337–348, October 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02464782.

Bordons:1997:CRT

[BZ97] Maria Bordons and M. A. Zulueta. Comparison of research team activity in two biomedical fields. Scientometrics, 40(3): 423–436, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459290.

Bassecoulard:1999:IRI

[BZ99] E. Bassecoulard and M. Zitt. Indicators in a research institute: A multi-level classification of scientific journals. Scientometrics, 44(3):323–345, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02458483.

Bordons:1999:MIC

[BZRB99] M. Bordons, M. A. Zulueta, F. Romero, and S. Barrigón. Measuring interdisciplinary collaboration within a university: The effects of the multidisciplinary research programme. Scientometrics, 46(3):383–398, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02459599.

Chatelin:1992:RSA

[CA92] Y. Chatelin and R. Arvanitis. Representing scientific activity by structural indicators: The case of Côte d’Ivoire 1884–1968. Scientometrics, 23(1):235–247, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020925.

Campanario:1995:UNNa

[Cam95a] J. M. Campanario. Using neural networks to study networks of scientific journals. Scientometrics, 33(1):23–40, May 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020773.

Campanario:1995:UNNb

[Cam95b] J. M. Campanario. Using neural networks to study networks of scientific journals. Scientometrics, 33(3):423, July 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/BF02017339.pdf.

Campanario:1996:UCC

[Cam96] J. M. Campanario. Using citation classics to study the incidence of serendipity in scientific discovery. Scientometrics, 37(1):3–24, September 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02093482.

Cano:1995:CPI

[Can95] V. Cano. Characteristics of the publishing infrastructure of peripheral countries: A comparison of periodical publications from Latin America with periodicals from the US and the UK. Scientometrics, 34(1):121–138, September 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019177.

Cunningham:1995:OCL

[CB95] Sally Jo Cunningham and D. Bocock. Obsolescence of computing literature. Scientometrics, 34(2):255–262, October 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020423.

Courtial:1991:IIS

[CC91] J. P. Courtial and M. Callon. Indicators for the identification of strategic themes within a research programme. Scientometrics, 21(3):447–457, July/August 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02093980.

Courtial:1994:MSI

[CCC94] J. P. Courtial, T. Cahlik, and M. Callon. A model for social interaction between cognition and action through a keyword simulation of knowledge growth. Scientometrics, 31(2): 173–192, October 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018559.

Callon:1991:CWA

[CCL91] M. Callon, J. P. Courtial, and F. Laville. Co-word analysis as a tool for describing the network of interactions between basic and technological research: The case of polymer chemistry. Scientometrics, 22(1):155–205, September 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019280.

Courtial:1993:UPT

[CCS93] J. P. Courtial, M. Callon, and A. Sigogneau. The use of patent titles for identifying the topics of invention and forecasting trends. Scientometrics, 26(2):231–242, February 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016216.

Chen:1993:TFR

[CCT93] Ye-Sho Chen, P. Pete Chong, and Yueguo Tong. Theoretical foundation of the 80/20 rule. Scientometrics, 28(2):183–204, October 1993. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02016899.

Cunningham:1997:API

[CD97] Sally Jo Cunningham and S. M. Dillon. Authorship patterns in information systems. Scientometrics, 39(1):19–27, May 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457428.

Colman:1995:BER

[CDC95] A. M. Colman, Debra Dhillon, and B. Coulthard. A bibliometric evaluation of the research performance of British university politics departments: Publications in leading journals. Scientometrics, 32(1):49–66, January 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020188.

Caraca:1993:RIS

[CdSM93] J. M. G. Caraça, C. Marciano da Silva, and L. Massimo. R&D indicators and socio-economic cohesion. Scientometrics, 26(2): 293–309, February 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016221.

Czapski:1992:CBI

[CFKS92] G. Czapski, A. Frenkel, D. Kohn, and A. Shoham. Cooperation between Israeli and foreign researchers. Scientometrics, 25(3):381–400, November 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016927.

Courtial:1997:SAA

[CG97] J. P. Courtial and L. Gourdon. A scientometric approach to autism based on translation sociology. Scientometrics, 40(2): 333–355, October 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457442.

Czerwon:1993:IPL

[CH93] H.-J. Czerwon and F. Havemann. Influence of publication languages on the citation rate of scientific articles: A case study of East German journals. Scientometrics, 26(1):51–63, January 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016792.

Chu:1992:CBCa

[Chu92a] H. Chu. Communication between Chinese and non-Chinese scientists in the discovery of high-tc superconductor: I. The formal perspective. Scientometrics, 25(2):229–252, October 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02028084.

Chu:1992:CBCb

[Chu92b] H. Chu. Communication between Chinese and non-Chinese scientists in the discovery of high-T_{c} superconductors: II. The informal perspective. Scientometrics, 25(2):253–277, October 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02028085.

Chung:1994:BDC

[Chu94] Y.-K. Chung. Bradford distribution and core authors in classification systems literature. Scientometrics, 29(2):253–269, February 1994. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02017976.

Christensen:1996:OCA

[CI96] F. Hjortgaard Christensen and P. Ingwersen. Online citation analysis. Scientometrics, 37(1):39–62, September 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093484.

Christensen:1997:ODJ

[CIW97] F. Hjortgaard Christensen, P. Ingwersen, and Irene Wormell. Online determination of the journal impact factor and its international properties. Scientometrics, 40(3):529–540, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459298.

Cano:1992:SIS

[CJ92] F. Cano and S. Julián. Some indicators in Spanish scientific production. Scientometrics, 24(1):43–59, May 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02026472.

Cano:1991:CLC

[CL91] V. Cano and N. C. Lind. Citation life cycles of ten citation classics. Scientometrics, 22(2):297–312, October 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020003.

Cambrosio:1993:HSM

[CLCL93] A. Cambrosio, C. Limoges, J. P. Courtial, and F. Laville. Historical scientometrics? Mapping over 70 years of biological safety research with coword analysis. Scientometrics, 27(2): 119–143, June 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016546.

Courtial:1990:MMD

[CM90] J.-P. Courtial and B. Michelet. A mathematical model of development in a research field. Scientometrics, 19(1–2):127– 141, July 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02130469.

Coronini:1999:ISV

[CM99] R. Coronini and V. Mangematin. From individual scientific visibility to collective competencies: The example of an academic department in the social sciences. Scientometrics, 45 (1):55–80, May 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458468.

Cesaratto:1991:IBI

[CMS91] S. Cesaratto, S. Mangano, and G. Sirilli. The innovative behaviour of Italian firms: A survey on technological innovation and R&D. Scientometrics, 21(1):115–141, May 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019186.

Cohen:1991:SAP

[Coh91] J. E. Cohen. Size, age and productivity of scientific and technical research groups. Scientometrics, 20(3):395–416, March 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019761.

Coleman:1993:BDS

[Col93] S. R. Coleman. Bradford distributions of social-science bibliographies varying in definitional homogeneity. Scientometrics, 27(1):75–91, May 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017756.

Coleman:1994:DVA

[Col94] S. R. Coleman. Disciplinary variables that affect the shape of Bradford’s bibliograph. Scientometrics, 29(1):59–81, January 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018384.

Courtial:1992:CLV

[Cou92] J.-P. Courtial. Comments on Leydesdorff’s’a validation study of Leximappe’. Scientometrics, 25(2):313–316, October 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.

1007/BF02028088. See [Ley92c, Ley92b].

Courtial:1994:CAS

[Cou94] J. P. Courtial. A coword analysis of scientometrics. Scientometrics, 31(3):251–260, November 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016875.

Cozzens:1995:URA

[Coz95] S. Cozzens. U.S. research assessment: Recent developments. Scientometrics, 34(3):351–362, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018004.

Collins:1991:EMC

[CR91] P. M. D. Collins and M. J. Ringe. Europeanization of the market for contract research. Scientometrics, 21(3):283–289, July/ August 1991. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02093971.

Chetal:1998:SRI

[CR98] Rakesh Chetal and Ashok Raj. Sponsored R&D in India: The project sponsoring pattern and main outcome of projects sponsored by major central departments/agencies. Scientometrics, 43(3):393–421, November 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457406.

Cronin:1998:MC

[Cro98] Blaise Cronin. Metatheorizing citation. Scientometrics, 43(1): 45–55, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458393.

Chawla:1998:OEP

[CS98] Anju Chawla and J. P. Singh. Organizational environment and performance of research groups — a typological analysis. Scientometrics, 43(3):373–391, November 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457405.

Cunion:1995:UGD

[Cun95] K. M. Cunion. UK government departments experience of RT&D programme evaluation and methodology. Scientometrics, 34(3):363–374, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018005.

Cunningham:1997:EEP

[Cun97] P. Cunningham. The evaluation of European programmes and the future of scientometrics. Scientometrics, 38(1):71–85, January 1997. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02461124.

Cunningham:1998:ABR

[Cun98] Sally Jo Cunningham. Applications for bibliometric research in the emerging digital libraries. Scientometrics, 43(2):161– 175, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458405.

Curien:1999:REP

[Cur99] H. Curien. Role of experts in political consultancy processes. Scientometrics, 45(3):467–471, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457609.

Chongde:1998:EMB

[CZ98] Wang Chongde and Wang Zhe. Evaluation of the models for Bradford’s Law. Scientometrics, 42(1):89–95, May 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02465014.

Czapski:1997:UDC

[Cza97] G. Czapski. The use of deciles of the citation impact to evaluate different fields of research in Israel. Scientometrics, 40(3): 437–443, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459291.

Czerwon:1990:SIS

[Cze90] H.-J. Czerwon. Scientometric indicators for a specialty in theoretical high-energy physics: Monte Carlo methods in lattice field theory. Scientometrics, 18(1–2):5–20, January 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019159.

deArenas:1992:PAM

[dA92] Judith Licea de Arenas. Partial assessment of Mexican health sciences research 1982–1986. Scientometrics, 23(1):47–55, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02020913.

DaSilvaESouza:1999:TEP

[DAD99] Geraldo Da Silva E. Souza, Eliseu Alves, and Antonio Flávio Dias Ávila. Technical efficiency of production in agricultural research. Scientometrics, 46(1):141–160, September 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02766299.

Dahoun:1999:BAS

[Dah99] A. Maxime Dahoun. Black Africa in the science citation index. Scientometrics, 46(1):11–18, September 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02766292.

Davignon:1999:EMP

[Dav99] Viscount E. Davignon. Evaluation in management and policy making at European level. Scientometrics, 45(3):347–354, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457588.

deArenas:1999:PMH

[dAVA99] Judith Licea de Arenas, J. Valles, and M. Arenas. Profile of the Mexican health sciences elite: A bibliometric analysis of research performance. Scientometrics, 46(3):539–547, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459610.

Duplenko:1995:CAC

[DB95] Y. K. Duplenko and S. G. Burchinsky. Computer-aided clustering of citation networks as a tool of mapping of research trends in biomedicine. Scientometrics, 32(3):247–258, March 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017644.

DeLooze:1996:DCJ

[DCL^{+}96] Marie-Angele De Looze, R. Coronini, Monique Legentil, P. Jeannin, and Marie-Hélene Magri. Determining the core of journals of a research centre: The example of researchers from the Department of Rural Economy and Sociology of the Institut National de la Recherche Agronomique, France. Scientometrics, 36(2):167–183, June 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02017312.

DeLooze:1996:STA

[De 96] Marie-Angèle De Looze. Scientometrics as a tool for analysis of the industrial relationships of two departments in a major French applied research institute, 1988–1992. Scientometrics, 37(2):349–360, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02093629.

DeHaan:1997:APD

[De 97] J. De Haan. Authorship patterns in Dutch sociology. Scientometrics, 39(2):197–208, June 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02457448.

DeOliveiraCabral:1998:STI

[De 98] J. E. De Oliveira Cabral. Survey on technological innovative behavior in the Brazilian food industry. Scientometrics, 42 (2):129–169, June 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458353.

Danell:1997:FMC

[DEP97] R. Danell, L. Engwall, and O. Persson. The first mover and the challenger: The relationship between two journals in organization research. Scientometrics, 40(3):445–453, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459292.

Daniel:1990:IQS

[DF90a] H.-D. Daniel and R. Fisch. Introduction: Quantitative science and technology indicators studies in the Federal Republic of Germany. Scientometrics, 19(5–6):327–329, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020695.

Daniel:1990:RPE

[DF90b] H.-D. Daniel and R. Fisch. Research performance evaluation in the German university sector. Scientometrics, 19(5–6):349– 361, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020698.

DeLange:1997:MMM

[DG97] C. De Lange and W. Glänzel. Modelling and measuring multilateral co-authorship in international scientific collaboration. Part I. Development of a new model using a series expansion approach. Scientometrics, 40(3):593–604, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459303.

Dhawan:1998:CSP

[Dha98] S. M. Dhawan. Comparative study of physics research in India and China based on INSPEC–Physics for 1990 and 1995. Scientometrics, 43(3):423–441, November 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457407.

deHaan:1994:AAD

[dHLR94] J. de Haan, F. L. Leeuw, and C. Remery. Accumulation of advantage and disadvantage in research groups. Scientometrics, 29(2):239–251, February 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017975.

DeMoya-Anegon:1999:SAL

[DHS99] F. De Moya-Anegón and V. Herrero-Solana. Science in America latina: A comparison of bibliometric and scientifictechnical indicators. Scientometrics, 46(2):299–320, October 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02464780.

Dickenson:1996:LRA

[Dic96] R. P. Dickenson. The level of research in advanced composite materials in the countries of the former Soviet union. Scientometrics, 36(1):43–57, May 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02126644.

Dierick:1992:DLP

[Die92] J. C. J. Dierick. Determining the Lotka parameters by sampling. Scientometrics, 25(1):115–148, September 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016851.

DeMoyaAnegon:1998:RFL

[DJD98] F. De Moya Anegón, E. Jiménez Contreras, and Mercedes De La Moneda Corrochano. Research fronts in library and information science in Spain (1985–1994). Scientometrics, 42 (2):229–246, June 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458357.

Dahl:1991:MNE

[DL91] M. Dahl and S. Lahlou. Measurement of network effects from the EC SCIENCE/STIMULATION programmes. Scientometrics, 21(3):325–342, July/August 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02093974.

deLooze:1994:AST

[dL94] Marie-Angèle de Looze. The application of scientometric tools to the analysis of a sector in plant biotechnologies: Nitrogen fixation. Scientometrics, 30(1):23–34, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017210.

DeLooze:1997:CRT

[DL97] Marie-Angèle De Looze and Juliette Lemarié. Corpus relevance through co-word analysis: An application to plant proteins. Scientometrics, 39(3):267–280, July 1997. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02458530.

Debackere:1999:CTL

[DLV99] K. Debackere, M. Luwel, and Reinhilde Veugelers. Can technology lead to a competitive advantage? A case study of Flanders using European patent data. Scientometrics, 44(3): 379–400, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458486.

DeBruin:1993:DSS

[DM93] R. E. De Bruin and H. F. Moed. Delimitation of scientific subfields using cognitive words from corporate addresses in scientific publications. Scientometrics, 26(1):65–80, January 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016793.

deMarchi:1999:SAS

[dMR99] M. de Marchi and M. Rocchi. Summing up approaches to the study of science and technology indicators. Scientometrics, 46(1):39–49, September 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02766294.

Dou:1994:WBW

[Dou94] H. Dou. In which business are we? Scientometrics, 30(2– 3):401–406, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018112.

Dou:1991:SDC

[DQH91] H. Dou, L. Quoniam, and Parina Hassanaly. The scientific dynamics of a city: A study of chemistry in Marseille from 1981 to the present. Scientometrics, 22(1):83–93, September 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019276.

Delgado:1992:ISP

[DR92] H. Delgado and Jane M. Russell. Impact of studies published in the international literature by scientists at the National University of Mexico. Scientometrics, 23(1):75–90, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020915.

Davis:1996:CAU

[DR96] G. Davis and Pamela Royle. A comparison of Australian university output using journal impact factors. Scientometrics, 35 (1):45–58, January 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018232.

DeLooze:1999:TMI

[DRC^{+}99] M.-A. De Looze, A. Roy, R. Coronini, M. Reinert, and O. Jouve. Two measures for identifying the perception of risk associated with the introduction of transgenic plants. Scientometrics, 44(3):401–426, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458487.

Davoust:1991:SPA

[DS91] E. Davoust and L. D. Schmadel. A study of the publishing activity of astronomers since 1969. Scientometrics, 22(1):9– 39, September 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019273.

deSmet:1992:IBS

[dS92] E. de Smet. Information behaviour in a scientific-technical environment: A survey with innovation engineers. Scientometrics, 25(1):101–113, September 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016850.

Dizon:1995:PPS

[DS95] L. B. Dizon and M. S. M. Sadorra. Patterns of publication by the staff of an international fisheries research center. Scientometrics, 32(1):67–75, January 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020189.

Dorban:1992:BAB

[DV92] M. Dorban and A. F. Vandevenne. Bibliometric analysis of bibliographic behaviours in economic sciences. Scientometrics, 25(1):149–165, September 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016852.

Davidse:1997:PIC

[DV97] R. J. Davidse and A. F. J. Van Raan. Out of particles: Impact of CERN, DESY and SLAC research to fields other than physics. Scientometrics, 40(2):171–193, October 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457436.

Davis:1999:OOI

[DWH99] M. Davis, C. S. Wilson, and W. W. Hood. Ophthalmology and optics: An informetric study of Australia’s contribution to fields in the vision science domain, 1991–95. Scientometrics, 46(3):399–416, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02459600.

ElAlami:1992:ISC

[EDM92] J. El Alami, J. C. Dore, and J. F. Miquel. International scientific collaboration in Arab countries. Scientometrics, 23(1): 249–263, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020926.

Egghe:1991:EPZ

[Egg91] L. Egghe. The exact place of Zipf’s and Pareto’s Law amongst the classical informetric laws. Scientometrics, 20(1):93–106, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02018147.

Egghe:1992:GTP

[Egg92] L. Egghe. Generalized transfer principles in econometrics and informetrics. Scientometrics, 25(1):167–191, September 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016853.

Egghe:1993:IGO

[Egg93] L. Egghe. On the influence of growth on obsolescence. Scientometrics, 27(2):195–214, June 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016550.

Egghe:1994:BGC

[Egg94a] L. Egghe. Bridging the gaps: Conceptual discussions on informetrics. Scientometrics, 30(1):35–47, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017211.

Egghe:1994:LSB

[Egg94b] L. Egghe. Little science, big science … and beyond. Scientometrics, 30(2–3):389–392, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018109.

Egghe:1997:FIA

[Egg97] L. Egghe. Fractal and informetric aspects of hypertext systems. Scientometrics, 40(3):455–464, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459293.

Egghe:1998:MTC

[Egg98] L. Egghe. Mathematical theories of citation. Scientometrics, 43(1):57–62, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02458394.

Egghe:1999:ERB

[Egg99a] L. Egghe. An explanation of the relation between the fraction of multinational publications and the fractional score of a country. Scientometrics, 45(2):291–310, June 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.

1007/BF02458438. See remarks [Bur01].

Egghe:1999:MMC

[Egg99b] L. Egghe. A model for measuring the congestion in library shelves. Scientometrics, 46(3):417–430, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459601.

Ehikhamenor:1990:PPS

[Ehi90] F. A. Ehikhamenor. Productivity of physical scientists in Nigerian universities in relation to communication variables. Scientometrics, 18(5–6):437–444, May 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020155.

Einem:1999:CS

[Ein99a] C. Einem. Closing session. Scientometrics, 45(3):571–572, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02457633.

Einem:1999:WOR

[Ein99b] C. Einem. Welcoming and opening remarks. Scientometrics, 45(3):343–346, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457587.

Exner:1995:CHR

[EK95] O. Exner and M. Kunz. Citation histories of related papers in the field of chemical correlation analysis. Scientometrics, 32 (1):3–10, January 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020185.

elKader:1998:HAN

[eKOM^{+}98] M. Abd el Kader, T. Ojasoo, J. F. Miquel, Y. Okubo, and J. C. Doré. Hierarchical author networks: An analysis of European molecular biology laboratory (EMBL) publications. Scientometrics, 42(3):405–421, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458379.

Englisch:1991:MSM

[Eng91] H. Englisch. Monotonous structure measures for social groups. Scientometrics, 21(2):159–169, June 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02017566.

Egghe:1992:CGM

[ER92] L. Egghe and I. K. Ravichandra Rao. Classification of growth models based on growth rates and its applications. Scientometrics, 25(1):5–46, September 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016845.

Egghe:1996:AGI

[ER96] L. Egghe and R. Rousseau. Average and global impact of a set of journals. Scientometrics, 36(1):97–107, May 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02126648.

Egghe:1995:IPU

[ERR95] L. Egghe, I. K. Ravichandra Rao, and R. Rousseau. On the influence of production on utilization functions: Obsolescence or increased use? Scientometrics, 34(2):285–315, October 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020425.

Egghe:1999:OLP

[ERY99] L. Egghe, R. Rousseau, and M. Yitzhaki. The “own-language preference”: Measures of relative language self-citation. Scientometrics, 45(2):217–232, June 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458434.

Escritt:1999:WOR

[Esc99] R. Escritt. Welcoming and opening remarks: Science and the academic system in transition — the role of evaluation. Scientometrics, 45(3):337, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02457585.

Eto:1991:SRO

[Eto91] H. Eto. Science revolution and Ortega Hypothesis in developing countries. Scientometrics, 20(1):283–295, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018159.

Eto:1994:QAS

[Eto94] H. Eto. Quality assessment of science information research and service. Scientometrics, 30(2–3):429–432, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018118.

Eto:1999:ISC

[Eto99a] H. Eto. The interest of scientific communities in searelated research topics. Scientometrics, 45(2):167–183, June 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458431.

Eto:1999:RMP

[Eto99b] H. Eto. Relationship of mathematical programming with mathematics, economic/regional planning and other specialties. Scientometrics, 45(2):311–324, June 1999. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02458439.

Fernandez:1992:CRP

[FAM^{+}92] M. T. Fernández, A. Agis, A. Martín, A. Cabrero, and I. Gómez. Cooperative research projects between the Spanish National Research Council and Latin–American institutions. Scientometrics, 23(1):137–148, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020919.

Farkas:1990:Na

[Far90a] J. Farkas. News. Scientometrics, 18(1–2):169–172, January 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019168.

Farkas:1990:Nb

[Far90b] J. Farkas. News. Scientometrics, 19(1–2):143–149, July 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02130470.

Farge:1999:QMI

[Far99] Y. Farge. Quantitative methods in industrial research and development. Scientometrics, 45(3):463–465, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457608.

Fasella:1999:IIB

[Fas99a] P. M. Fasella. Interdependencies and interactions between evaluation and decision making processes. Scientometrics, 45 (3):405–408, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457598.

Fasella:1999:UEI

[Fas99b] P. M. Fasella. Universities — engines of innovation in the information society. Scientometrics, 45(3):543–545, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457625.

Faust:1990:EIT

[Fau90] K. Faust. Early identification of technological advances on the basis of patent data. Scientometrics, 19(5–6):473–480, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02020708.

Fayl:1999:MIE

[Fay99] G. Fayl. Measuring the impact of EU Framework Programme at national level. Scientometrics, 45(3):507, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02457619.

Fernandez-Cano:1999:SSP

[FCB99] A. Fernández-Cano and A. Bueno. Synthesizing scientometric patterns in Spanish educational research. Scientometrics, 46(2):349–367, October 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02464783.

Fernandez:1998:TIS

[Fer98] J. A. Fernández. The transition from an individual science to a collective one: The case of astronomy. Scientometrics, 42(1):61–74, May 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02465012.

Fichtner:1990:CUS

[Fic90] D. Fichtner. Competition in the university system of the Federal Republic of Germany. Scientometrics, 19(5–6):331–335, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02020696.

Finkenstaedt:1990:MRP

[Fin90] T. Finkenstaedt. Measuring research performance in the humanities. Scientometrics, 19(5–6):409–417, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020703.

Formann:1992:APS

[For92] A. K. Formann. Academic personnel selection: Description and prognosis of the decisions made by the committee for the selection of candidates for a full professorship. Scientometrics, 25(3):401–414, November 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016928.

Faucompre:1997:ELB

[FQD97] P. Faucompré, L. Quoniam, and H. Dou. An effective link between science and technology. Scientometrics, 40(3):465– 480, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459294.

Falkingham:1998:CAT

[FR98] L. T. Falkingham and R. Reeves. Context analysis — a technique for analysing research in a field, applied to literature on the management of R&D at the section level. Scientometrics, 42(2):97–120, June 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458351.

Frame:1991:MNT

[Fra91] J. Frame. Modelling national technological capacity with patent indicators. Scientometrics, 22(3):327–339, November 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019766.

Ferreiro:1992:CMJ

[FU92] L. Ferreiro and S. Ugena. Citation mechanics in journals covered by the journal citation reports. Scientometrics, 24(1):

149–162, May 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02026478.

Fujigaki:1998:CSC

[Fuj98] Yuko Fujigaki. The citation system: Citation networks as repeatedly focusing on difference, continuous re-evaluation, and as persistent knowledge accumulation. Scientometrics, 43(1): 77–85, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458397.

Fonseca:1998:RBA

[FVWD98] Lucia Fonseca, Sancia Velloso, Susana Wofchuk, and L. De Meis. The relationship between advisors and students. Scientometrics, 41(3):299–312, March 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02459047.

Fonseca:1997:IHR

[FVWdM97] Lucia Fonseca, Sancia Velloso, Susana Wofchuk, and L. de Meis. The importance of human relationships in scientific productivity. Scientometrics, 39(2):159–171, June 1997. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02457445.

Gabolde:1995:OA

[Gab95] I. Gabolde. Opening address. Scientometrics, 34(3):317–320, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02018001.

Gabolde:1996:E

[Gab96] I. Gabolde. Erratum. Scientometrics, 36(1):143, May 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/BF02126652.pdf.

Gaillard:1992:UPL

[Gai92] J. Gaillard. Use of publication lists to study scientific production and strategies of scientists in developing countries. Scientometrics, 23(1):57–73, January 1992. CODEN SCNTDX.

ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020914.

Garfield:1998:RTC

[Gar98] E. Garfield. Random thoughts on citationology its theory and practice. Scientometrics, 43(1):69–76, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458396.

Gomez:1996:CPS

[GBFM96] Isabel Gómez, Maria Bordons, M. T. Fernández, and Aida Méndez. Coping with the problem of subject classification diversity. Scientometrics, 35(2):223–235, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018480.

Godin:1995:BAP

[GBL95] B. Godin, R. S. Barker, and M. Landry. Besides academic publications: Which sectors compete, or are there competitors? Scientometrics, 33(1):3–12, May 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020771.

Glanzel:1996:NMA

[GC96] W. Glänzel and H. J. Czerwon. A new methodological approach to bibliographic coupling and its application to the national, regional and institutional level. Scientometrics, 37(2): 195–221, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02093621.

Glanzel:1997:MMM

[GD97] W. Glänzel and C. De Lange. Modelling and measuring multilateral co-authorship in international scientific collaboration. Part II. A comparative study on the extent and change of international scientific collaboration links. Scientometrics, 40(3): 605–626, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459304.

Geisler:1996:IFM

[Gei96] E. Geisler. Integrated figure of merit of public sector research evaluation. Scientometrics, 36(3):379–395, July/ August 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02129601.

Georghiou:1999:MEE

[Geo99] L. Georghiou. Meta-evaluation: Evaluation of evaluations. Scientometrics, 45(3):523–530, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457622.

Gomez:1999:ASI

[GFS99] Isabel Gómez, María Teresa Fernández, and Jesús Sebastián. Analysis of the structure of international scientific cooperation networks through bibliometric indicators. Scientometrics, 44 (3):441–457, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458489.

Grupp:1994:IOE

[GH94] H. Grupp and Sybille Hinze. International orientation, efficiency of and regard for research in East and West Germany: A bibliometric investigation of aspects of technology genesis in the United Germany. Scientometrics, 29(1):83–113, January 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018385.

Guimaraes:1995:THR

[GH95] J. A. Guimarães and Marta C. Humann. Training of human resources in science and technology in Brazil: The importance of a vigorous post-graduate program and its impact on the development of the country. Scientometrics, 34(1):101–119, September 1995. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02019176.

Garrison:1992:MCS

[GHL92] H. H. Garrison, S. S. Herman, and J. A. Lipton. Measuring characteristics of scientific research: A comparison of bibliographic and survey data. Scientometrics, 24(2):359–370, June 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017915.

Godin:1996:SCR

[GI96] B. Godin and M.-P. Ippersiel. Scientific collaboration at the regional level: The case of a small country. Scientometrics, 36(1):59–68, May 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02126645.

Giese:1990:RUF

[Gie90] E. Giese. Rankings of universities in the FRG. Scientometrics, 19(5–6):363–375, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02020699.

Gillett:1991:PAR

[Gil91] R. Gillett. Pitfalls in assessing research performance by grant income. Scientometrics, 22(2):253–263, October 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020000.

Giorgi:1993:LTA

[Gio93] Eleonora P. Giorgi. Long term analysis of citation counts at the micro-level. Scientometrics, 28(3):375–386, November/ December 1993. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02026516.

Gonda:1995:RTD

[GK95] Kinji Gonda and Fumihiko Kakizaki. Research, technology and development evaluation; developments in Japan. Scientometrics, 34(3):375–389, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018006.

Gupta:1996:APP

[GK96] B. M. Gupta and C. R. Karisiddappa. Author productivity patterns in theoretical population genetics (1900–1980). Scientometrics, 36(1):19–41, May 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02126643.

Gupta:1997:PAR

[GK97] B. M. Gupta and C. R. Karisiddappa. Productivity of authors as reflected by duration of their scientific participation and speed of publication. Scientometrics, 39(3):281–291, July 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458531.

Gupta:1998:CTP

[GK98] B. M. Gupta and C. R. Karisiddappa. Collaboration in theoretical population genetics speciality. Scientometrics, 42(3): 349–376, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458377.

Gupta:1999:CAP

[GK99] B. M. Gupta and C. R. Karisiddappa. Collaboration and author productivity: A study with a new variable in Lotka’s Law. Scientometrics, 44(1):129–134, January 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458481.

Gupta:1999:CPM

[GKA99] B. M. Gupta, Suresh Kumar, and B. S. Aggarwal. A comparision of productivity of male and female scientists of CSIR. Scientometrics, 45(2):269–289, June 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458437.

Gupta:1997:CPT

[GKK97] B. M. Gupta, Suresh Kumar, and C. R. Karisiddappa. Collaboration profile of theoretical population genetics speciality. Scientometrics, 39(3):293–314, July 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458532.

Glanzel:1996:P

[GKMS96] W. Glänzel, S. Katz, H. Moed, and U. Schoepflin. Preface. Scientometrics, 35(2):165–166, February 1996. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02018474.

Gupta:1998:ASP

[GKR98] B. M. Gupta, Suresh Kumar, and R. Rousseau. Applicability of selected probability distributions to the number of authors per article in theoretical population genetics. Scientometrics, 42(3):325–334, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458375.

Glanzel:1990:BR

[Glä90] Wolfgang Glänzel. Book review. Scientometrics, 19(5–6):523– 524, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02020713.

Glanzel:1996:BAS

[Glä96a] W. Glänzel. A bibliometric approach to social sciences. National research performances in 6 selected social science areas, 1990–1992. Scientometrics, 35(3):291–307, March 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016902.

Glanzel:1996:NSB

[Glä96b] W. Glänzel. The need for standards in bibliometric research and technology. Scientometrics, 35(2):167–176, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018475.

Glanzel:1997:PRP

[Glä97] W. Glänzel. On the possibility and reliability of predictions based on stochastic citation processes. Scientometrics, 40(3): 481–492, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459295.

Granovsky:1992:IBE

[GLMM92] Yu. V. Granovsky, T. N. Luibimova, T. I. Murashova, and V. D. Myatlev. Information-based evaluation of the quality of doctoral theses. Scientometrics, 23(3):361–376, March 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02029804.

Godin:1998:MKF

[God98] B. Godin. Measuring knowledge flows between countries: The use of scientific meeting data. Scientometrics, 42(3):313–323, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02458374.

Golub:1998:CSE

[Gol98] Branka Golub. The Croatian scientific elite and its socioprofessional roots. Scientometrics, 43(2):207–229, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458408.

Goldberg:1997:SAN

[GOR97] A. I. Goldberg, Luydmilla Oigenblick, and Ami-Hai E. Rubin. Scientific articles and national medical cultures: A comparison of Russian and American medical journals. Scientometrics, 39(1):57–75, May 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457430.

Garg:1998:SSL

[GP98] K. C. Garg and P. Padhi. Scientometric study of laser patent literature. Scientometrics, 43(3):443–446, November 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457408.

Garg:1999:SIP

[GP99a] K. C. Garg and P. Padhi. Scientometrics of institutional productivity of laser science and technology. Scientometrics, 46(1):19–38, September 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02766293.

Garg:1999:SLR

[GP99b] K. C. Garg and P. Padhi. Scientometrics of laser research literature as viewed through the Journal of Current Laser Abstracts. Scientometrics, 45(2):251–268, June 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458436.

Grivel:1997:CSB

[GPK97] L. Grivel, X. Polanco, and A. Kaplan. A computer system for big scientometrics at the age of the World Wide Web. Scientometrics, 40(3):493–506, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02459296.

Griffith:1994:LSL

[Gri94] B. C. Griffith. Little scientometrics, little scientometrics, little scientometrics, little scientometrics, … and so on and so on. Scientometrics, 30(2–3):487–493, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018129.

Grupp:1990:CES

[Gru90a] H. Grupp. The concept of entropy in scientometrics and innovation research. Scientometrics, 18(3–4):219–239, March 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017763.

Grupp:1990:SFS

[Gru90b] H. Grupp. On the supplementary functions of science and technology indicators. Scientometrics, 19(5–6):447–472, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020707.

Garfield:1991:MJM

[GS91a] E. Garfield and H. Small. Michael J. Moravcsik: Multidimensional scholar and hero of third world science. Scientometrics, 20(1):19–23, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018140.

Garg:1991:SPR

[GS91b] K. C. Garg and Praveen Sharma. Solar power research: A scientometric study of world literature. Scientometrics, 21 (2):147–157, June 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017565.

Glanzel:1992:SFF

[GS92] W. Glänzel and A. Schubert. Some facts and figures on highly cited papers in the sciences, 1981–1985. Scientometrics, 25(3): 373–380, November 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016926.

Glanzel:1993:CSD

[GS93] W. Glänzel and A. Schubert. A characterization of scientometric distributions based on harmonic means. Scientometrics, 26 (1):81–96, January 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016794.

Glanzel:1994:LSB

[GS94a] W. Glänzel and U. Schoepflin. Little scientometrics, big scientometrics … and beyond? Scientometrics, 30(2–3):375– 384, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02018107.

Glanzel:1994:SMA

[GS94b] W. Glänzel and U. Schoepflin. A stochastic model for the ageing of scientific literature. Scientometrics, 30(1):49–64, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017212.

Galante:1996:REI

[GS96] E. Galante and C. Sala. R&D evaluation at the Italian National Research Council: The agricultural sector. Scientometrics, 36(2):207–222, June 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017314.

Glanzel:1999:BAI

[GSC99a] W. Glänzel, A. Schubert, and H.-J. Czerwon. A bibliometric analysis of international scientific cooperation of the European Union (1985–1995). Scientometrics, 45(2):185–202, June 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458432.

Glanzel:1999:IISa

[GSC99b] W. Glänzel, A. Schubert, and H.-J. Czerwon. An item-by-item subject classification of papers published in multidisciplinary and general journals using reference analysis. Scientometrics, 44(3):427–439, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02458488.

Grupp:1991:PPI

[GSK91] H. Grupp, U. Schmoch, and U. Kuntze. Patents as potential indicators of the utility of EC research programmes. Scientometrics, 21(3):417–445, July/August 1991. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02093979.

Gupta:1995:MGP

[GSK95] B. M. Gupta, Lalita Sharma, and C. R. Karisiddappa. Modelling the growth of papers in a scientific specialty. Scientometrics, 33(2):187–201, June 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02020568.

Gupta:1997:GRL

[GSK97] B. M. Gupta, Praveen Sharma, and C. R. Karisiddappa. Growth of research literature in scientific specialities. A modelling perspective. Scientometrics, 40(3):507–528, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459297.

Gupta:1999:GWI

[GSK99] B. M. Gupta, Praveen Sharma, and Suresh Kumar. Growth of world and Indian physics literature. Scientometrics, 44 (1):5–16, January 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458474.

Gupta:1990:SBP

[GSM90] B. M. Gupta, S. C. Sharma, and N. N. Mehrotra. Subjectbased publication activity indicators for medicinal & aromatic plants research. Scientometrics, 18(5–6):341–361, May 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020150.

Gomez:1999:ILA

[GSMF99] Isabel Gómez, Rosa Sancho, Luz Moreno, and María Teresa Fernández. Influence of Latin American journals coverage by international databases. Scientometrics, 46(3):443–456, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02459603.

Garg:1993:BLR

[GSS93] K. C. Garg, Praveen Sharma, and Lalita Sharma. Bradford’s Law in relation to the evolution of a field. A case study of solar power research. Scientometrics, 27(2):145–156, June 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016547.

Guzman:1998:BSV

[GSS98] M. V. Guzman, E. Sanz, and G. Sotolongo. Bibliometric study on vaccines (1990–1995). Part I: Scientific production in Iberian–American countries. Scientometrics, 43(2):189–205, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02458407.

Glanzel:1999:IISb

[GSSC99] W. Glänzel, A. Schubert, U. Schoepflin, and H. J. Czerwon. An item-by-item subject classification of papers published in journals covered by the SSCI database using reference analysis. Scientometrics, 46(3):431–441, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459602.

Gupta:1993:CRT

[Gup93] D. K. Gupta. Collaborative research trend in exploration geophysics. Scientometrics, 28(3):287–296, November/December 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02026511.

Gupta:1997:ADA

[Gup97] B. M. Gupta. Analysis of distribution of the age of citations in theoretical population genetics. Scientometrics, 40(1):139– 162, September 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459265.

Gupta:1998:GOL

[Gup98] B. M. Gupta. Growth and obsolescence of literature in theoretical population genetics. Scientometrics, 42(3):335–347, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02458376.

Gupta:1999:TTA

[Gup99] V. K. Gupta. Technological trends in the area of fullerenes using bibliometric analysis of patents. Scientometrics, 44(1): 17–31, January 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458475.

Glanzel:1992:ICT

[GW92] W. Glänzel and M. Winterhager. International collaboration of three east European countries with Germany in the sciences, 1980–1989. Scientometrics, 25(2):219–227, October 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02028083.

Haitun:1992:PILa

[Hai92a] S. D. Haitun. The problem of indicator-latent relationship in metric models. Part 1. Scientometrics, 23(2):335–351, February 1992. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02017045.

Haitun:1992:PILb

[Hai92b] S. D. Haitun. The problem of indicator-latent relationship in metric models. Part 2. Scientometrics, 24(2):221–235, June 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017910.

Haiqi:1994:BSM

[Hai94] Zhang Haiqi. A bibliometric study on medicine Chinese traditional in Medline database. Scientometrics, 31(3):241–250, November 1994. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02016874.

Haiqi:1996:RPK

[Hai96] Zhang Haiqi. Research performance in key medical universities in China observed from the scientific productivity. Scientometrics, 37(1):177–190, September 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02093494.

Hall:1992:SII

[Hal92] D. H. Hall. The science-industry interface: Correlation of time series of indicators and their spectra, and growth models in the nuclear fuels industry. Scientometrics, 24(2):237–280, June 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017911.

Hall:1993:SII

[Hal93] D. H. Hall. The science-industry interface in the petroleum industry: Correlation of time series of indicators and their spectra, and growth modelling. Scientometrics, 28(3):237– 286, November/December 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02026510.

Hurt:1992:MLS

[HB92] C. D. Hurt and J. M. Budd. Modeling the literature of superstring theory: A case of fast literature. Scientometrics, 24 (3):471–480, July 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02051042.

Heine:1995:CTC

[Hei95] M. H. Heine. The characterisation of text corpora using an input/output schema for citations. Scientometrics, 32(2):177– 194, February 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016893.

Helander:1995:EAN

[Hel95] Elisabeth Helander. Evaluation activities in the Nordic countries. Scientometrics, 34(3):391–400, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018007.

Helander:1999:EST

[Hel99] Elisabeth Helander. Evaluation of science and technology in the structural funds. Scientometrics, 45(3):517–521, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457621.

Hemlin:1993:SQE

[Hem93] S. Hemlin. Scientific quality in the eyes of the scientist. A questionnaire study. Scientometrics, 27(1):3–18, May 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017752.

Herbstein:1993:MPO

[Her93] F. H. Herbstein. Measuring “publications output” and “publications impact” of faculty members of a university chemistry department. Scientometrics, 28(3):349–373, November/ December 1993. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02026515.

Herbertz:1995:DIP

[Her95] H. Herbertz. Does it pay to cooperate? A bibliometric case study in molecular biology. Scientometrics, 33(1):117– 122, May 1995. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02020777.

Hayashi:1999:DKP

[HF99] Takayuki Hayashi and Yuko Fujigaki. Differences in knowledge production between disciplines based on analysis of paper styles and citation patterns. Scientometrics, 46(1):73–86, September 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02766296.

Hemlin:1996:RPA

[HG96a] S. Hemlin and M. Gustafsson. Research production in the arts and humanities. Scientometrics, 37(3):417–432, November 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019256.

Hinze:1996:MRS

[HG96b] Sybille Hinze and H. Grupp. Mapping of R&D structures in transdisciplinary areas: New biotechnology in food sciences. Scientometrics, 37(2):313–335, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093627.

Hargens:1990:NCA

[HH90] L. L. Hargens and J. R. Herting. Neglected considerations in the analysis of agreement among journal referees. Scientometrics, 19(1–2):91–106, July 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02130467. See [HH06]

and comments [Lin91a].

Harsanyi:1993:EE

[HH93] Martha A. Harsanyi and S. P. Harter. Ecclesiastes effects. Scientometrics, 27(1):93–96, May 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02017757.

Haiqi:1997:SRC

[HH97] Zhang Haiqi and Guo Hong. Scientific research collaboration in China. Scientometrics, 38(2):309–319, February 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457416.

Hargens:2006:AAB

[HH06] Lowell L. Hargens and Jerald R. Herting. Analyzing the association between referees’ recommendations and editors’ decisions. Scientometrics, 67(1):15–26, April 2006. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-006-0048-6. See [HH90] and comments

[Lin91a].

Hodges:1996:UAA

[HHM^{+}96] Sally Hodges, B. Hodges, A. J. Meadows, Micheline Beaulieu, and D. Law. The use of an algorithmic approach for the assessment of research quality. Scientometrics, 35(1):3–13, January 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018229.

Hicks:1999:DAF

[Hic99] Diana Hicks. The difficulty of achieving full coverage of international social science literature and the bibliometric consequences. Scientometrics, 44(2):193–215, February 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457380.

Higgins:1991:IES

[Hig91] T. Higgins. Indicators of European scientific cohesion. Scientometrics, 21(3):367–381, July/August 1991. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02093976.

Hills:1995:PEE

[Hil95] P. Hills. Prest’s experience of evaluation. Scientometrics, 34(3):401–414, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018008.

Hinze:1994:BCE

[Hin94] Sybille Hinze. Bibliographical cartography of an emerging interdisciplinary discipline: The case of bioelectronics. Scientometrics, 29(3):353–376, March 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02033445.

Hinze:1999:CCD

[Hin99] Sybille Hinze. Collaboration and cross-disciplinarity in autoimmune diseases. Scientometrics, 46(3):457–471, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459604.

Hartmann:1990:PRD

[HN90] I. Hartmann and F. Neidhardt. Peer review at the Deutsche Forschungsgemeinschaft. Scientometrics, 19(5–6):419–425, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02020704.

Hussain:1993:BOR

[HN93] S. S. M. Hussain and D. A. Nunez. British otorhinolaryngological research. An analysis of publication trends. Scientometrics, 26(2):255–262, February 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016218.

Hogan:1990:MAF

[Hog90] T. J. Hogan. A measure of accounting faculties and doctoral programs. Scientometrics, 19(3–4):207–221, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02095348.

Horvat:1999:MEE

[Hor99] M. Horvat. Meta-evaluation: Evaluation of evaluations some points for discussion. Scientometrics, 45(3):533–542, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457624.

Huot:1992:NMA

[HQD92] Ch. Huot, L. Quoniam, and H. Dou. A new method for analysing downloaded data for strategic decision. Scientometrics, 25(2):279–294, October 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02028086.

Hart:1998:RBG

[HS98] P. W. Hart and J. T. Sommerfeld. Relationship between growth in gross domestic product (GDP) and growth in the chemical engineering literature in five different countries. Scientometrics, 42(3):299–311, October 1998. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02458373.

Huber:1999:IPS

[Hub99] J. C. Huber. Inventive productivity and the statistics of exceedances. Scientometrics, 45(1):33–53, May 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458467.

Hood:1999:DBR

[HW99] W. W. Hood and Concepción S. Wilson. The distribution of bibliographic records in databases using different counting methods for duplicate records. Scientometrics, 46(3):473–486, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02459605.

Irvine:1989:ICS

[IM89] J. Irvine and B. R. Martin. International comparisons of scientific performance revisited. Scientometrics, 15(5–6):369– 392, May 1989. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02017060. See reply [Kea91].

Ingwersen:1999:PBI

[IW99] P. Ingwersen and Irene Wormell. Publication behaviour and international impact: Scandinavian clinical and social medicine, 1988–96. Scientometrics, 46(3):487–499, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459606.

Jaschek:1991:SAC

[Jas91] C. Jaschek. The size of the astronomical community. Scientometrics, 22(2):265–282, October 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020001.

Jaschek:1992:VWE

[Jas92] C. Jaschek. The “visibility” of West European astronomical research. Scientometrics, 23(3):377–393, March 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02029805.

Jimenez-Contreras:1996:PAF

[JCFA96] E. Jiménez-Contreras and L. Ferreiro-Aláez. Publishing abroad: Fair trade or short sell for non-English-speaking authors? A Spanish study. Scientometrics, 36(1):81–95, May 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02126647.

Jeannin:1994:TDA

[JD94] P. Jeannin and J. Devillard. Towards a demographic approach to scientific journals. Scientometrics, 30(1):83–95, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017214.

Jain:1992:LRI

[JG92] Ashok Jain and K. C. Garg. Laser research in India: Scientometric study and model projections. Scientometrics, 23 (3):395–415, March 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02029806.

Jain:1998:ISF

[JGSK98] Ashok Jain, K. C. Garg, Praveen Sharma, and Suresh Kumar. Impact of SERC’s funding on research in chemical sciences. Scientometrics, 41(3):357–370, March 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459051.

Johnes:1992:AOA

[JJ92] Geraint Johnes and Jill Johnes. Apples and oranges: The aggregation problem in publications analysis. Scientometrics, 25(2):353–365, October 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02028091.

Johnston:1995:RIQ

[Joh95] R. Johnston. Research impact quantification. Scientometrics, 34(3):415–426, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018009.

Jagodzinski-Sigogneau:1991:SIB

[JSBCF91] M. Jagodzinski-Sigogneau, S. Bauin, J. P. Courtial, and H. Feillet. Scientific innovation in bibliographical databases: A comparative study of the Science Citation Index and the Pascal database. Scientometrics, 22(1):65–82, September 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019275.

Jin:1999:CSC

[JW99] Bihui Jin and Bing Wang. Chinese science citation database: Its construction and application. Scientometrics, 45(2):325– 332, June 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02458440.

Kameoka:1995:ERP

[Kam95] A. Kameoka. Evaluating research projects at Toshiba. Designing a conceptual framework of evaluating research and technology development (RTD) programs. Scientometrics, 34(3): 427–439, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018010.

Karki:1990:ESR

[Kar90] M. M. S. Karki. Environmental science research in India: An analysis of publications. Scientometrics, 18(5–6):363– 373, May 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02020151.

Katz:1994:GPS

[Kat94] J. S. Katz. Geographical proximity and scientific collaboration. Scientometrics, 31(1):31–43, September 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018100.

Katz:1996:BSP

[Kat96] J. S. Katz. Bibliometric standards: Personal experience and lessons learned. Scientometrics, 35(2):193–197, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018477.

Kealey:1991:GFA

[Kea91] T. Kealey. Government-funded academic science is a consumer good, not a producer good: A comparative reassessment of Britain’s scientific and technological achievements since 1794 and a comment on the bibliometry of B. Martin and J. Irvine. Scientometrics, 20(2):369–394, February 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017527. See [IM89].

Keay:1992:PPR

[Kea92] C. S. L. Keay. Physics, psychology and respectability. Scientometrics, 23(3):355–359, March 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02029803.

Kostoff:1997:DTT

[KETP97] R. N. Kostoff, H. J. Eberhart, D. R. Toothman, and R. Pellenbarg. Database tomography for technical intelligence: Comparative roadmaps of the research impact assessment literature and the journal of the American chemical society. Scientometrics, 40(1):103–138, September 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02459264.

Karki:1995:IRI

[KG95] M. M. S. Karki and K. C. Garg. Industrial research in India as viewed through research & industry. Scientometrics, 33 (3):315–328, July 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017334.

Kretschmer:1998:CPT

[KG98] Hildrun Kretschmer and B. M. Gupta. Collaboration patterns in theoretical population genetics. Scientometrics, 43(3):455– 462, November 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457409.

Karki:1999:SIO

[KG99] M. M. S. Karki and K. C. Garg. Scientometrics of Indian organic chemistry research. Scientometrics, 45(1):107–116, May 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458471.

Katz:1996:SVB

[KH96] J. S. Katz and D. M. Hicks. A systemic view of British science. Scientometrics, 35(1):133–154, January 1996. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02018237.

Katz:1997:DS

[KH97a] J. S. Katz and Diana Hicks. Desktop scientometrics. Scientometrics, 38(1):141–153, January 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02461128.

Katz:1997:HMC

[KH97b] J. S. Katz and Diana Hicks. How much is a collaboration worth? A calibrated bibliometric model. Scientometrics, 40(3):541–554, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02459299.

Kundra:1999:NMS

[KK99] Ramesh Kundra and Hildrun Kretschmer. A new model of scientific collaboration. Part 2. Collaboration patterns in Indian medicine. Scientometrics, 46(3):519–528, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459608.

Kademani:1999:SPN

[KKJ99] B. S. Kademani, V. L. Kalyane, and Suresh Jange. Scientometric portrait of Nobel laureate Dorothy Crowfoot Hodgkin. Scientometrics, 45(2):233–250, June 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458435.

Kyvik:1994:ICR

[KL94] S. Kyvik and I. Marheim Larsen. International contact and research performance. Scientometrics, 29(1):161–172, January 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018388.

Klaic:1990:SAR

[Kla90] B. Klaić. Scientometric analysis of the research activities of chemists from the “Rugjer Bosković” Institute (Yugoslavia), 1976–1985. Scientometrics, 19(1–2):11–24, July 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02130462.

Klaic:1995:ASP

[Kla95] B. Klaić. Analysis of the scientific productivity of researchers from the Republic of Croatia for the period 1990–1992. Scientometrics, 32(2):133–152, February 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016891.

Kretschmer:1990:CDO

[KM90] Hildrun Kretschmer and Renate Müller. A contribution to the dispute on the Ortega Hypothesis: Connection between publication rate and stratification of scientists, tested by various methods. Scientometrics, 18(1–2):43–56, January 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019161.

Kalyane:1995:SPW

[KM95] V. L. Kalyane and S. S. Munnolli. Scientometric portrait of T. S. West. Scientometrics, 33(2):233–256, June 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020571.

Kishida:1997:IPP

[KM97] Kazuaki Kishida and Sachiko Matsui. International publication patterns in social sciences: A quantitative analysis of the IBSS file. Scientometrics, 40(2):277–298, October 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457440.

Kneucker:1999:EMP

[Kne99] R. F. Kneucker. Evaluation in management and policy making at European level. Scientometrics, 45(3):355–357, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457589.

Kortelainen:1997:ACD

[Kor97] Terttu A. M. Kortelainen. Applying concepts of diffusion research in an informetric study. Scientometrics, 40(3):555–568, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02459300.

Kostoff:1995:FRI

[Kos95] R. N. Kostoff. Federal research impact assessment: Axioms, approaches, applications. Scientometrics, 34(2):163–206, October 1995. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02020420.

Kostoff:1996:PMG

[Kos96] R. N. Kostoff. Performance measures for governmentsponsored research: Overview and background. Scientometrics, 36(3):281–292, July/August 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02129595.

Kostoff:1997:CAC

[Kos97] R. N. Kostoff. Citation analysis cross-field normalization: A new paradigm. Scientometrics, 39(3):225–230, July 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458527.

Kostoff:1998:UMC

[Kos98] R. N. Kostoff. The use and misuse of citation analysis in research evaluation. Scientometrics, 43(1):27–43, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458392.

Krauskopf:1992:SIM

[Kra92] M. Krauskopf. Scientometric indicators as a means to assess the performance of state supported universities in developing countries: The Chilean case. Scientometrics, 23(1):105–121, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02020917.

Krauskopff:1994:ETC

[Kra94] M. Krauskopff. Epistemometria, a term contributing to express the meaning and potential methodologies of scientometrics in Spanish speaking countries. Scientometrics, 30(2–3): 425–428, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018117.

Kretschmer:1990:PCB

[Kre90] Hildrun Kretschmer. Pinski’s citation based measures of research interactivity and the application of a complex structure measure to journal systems. Scientometrics, 18(1–2):95–122, January 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02019164.

Kretschmer:1992:CPF

[Kre92] Hildrun Kretschmer. Call for papers Fourth International Conference on Bibliometrics, Informetrics, and Scientometrics. Scientometrics, 25(3):465, November 1992. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02016932.

Kretschmer:1993:MSS

[Kre93] Hildrun Kretschmer. Measurement of social stratification. A contribution to the dispute on the ORTEGA hypothesis. Scientometrics, 26(1):97–113, January 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016795.

Kretschmer:1994:CNI

[Kre94a] Hildrun Kretschmer. Coauthorship networks of invisible colleges and institutionalized communities. Scientometrics, 30 (1):363–369, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017234.

Kretschmer:1994:QQS

[Kre94b] Hildrun Kretschmer. Quantity and quality in ‘science of science’. Scientometrics, 30(2–3):533–537, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018136.

Kretschmer:1997:PBC

[Kre97] Hildrun Kretschmer. Patterns of behaviour in coauthorship networks of invisible colleges. Scientometrics, 40(3):579–591, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02459302.

Kretschmer:1999:NMS

[Kre99] Hildrun Kretschmer. A new model of scientific collaboration. Part 1. Theoretical approach. Scientometrics, 46(3):501–518, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02459607.

Kozlowski:1999:HMI

[KRI99] J. Kozlowski, S. Radosevic, and D. Ircha. History matters: The inherited disciplinary structure of the post-Communist science in countries of Central and Eastern Europe and its restructuring. Scientometrics, 45(1):137–166, May 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458473.

Krull:1995:MPE

[Kru95] W. Krull. The max Planck experience of evaluation. Scientometrics, 34(3):441–450, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018011.

Krull:1999:COM

[Kru99] W. Krull. Challenges to output measurement in science and technology programme evaluation. Scientometrics, 45(3):501– 504, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02457617.

Kuhlmann:1995:GGD

[Kuh95] S. Kuhlmann. German government Department’s experience of RT & d programme evaluation and methodology. Scientometrics, 34(3):461–471, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018013.

Kuklinski:1999:EUG

[Kuk99] A. Kuklinski. The European University in the global landscape of the 21^{st} century. Scientometrics, 45(3):561–564, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457629.

Kunz:1990:CLD

[Kun90] M. Kunz. Can the lognormal distribution be rehabilitated? Scientometrics, 18(3–4):179–191, March 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017760.

Kundra:1996:ICR

[Kun96] R. Kundra. Investigation of collaborative research trends in Indian medical sciences: 1900–1945. Scientometrics, 36(1):69– 80, May 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02126646.

Kutlaca:1998:PRA

[Kut98] – D. G. Kutlača. Patent-related activities in Serbia from 1921 to 1995. Scientometrics, 42(2):171–193, June 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458354.

Krauskopf:1997:ASP

[KV97] M. Krauskopf and María Inés Vera. Assessment of scientific profiles and capabilities of Ph.D. programs in Chile: A scientometric approach. Scientometrics, 40(3):569–577, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459301.

Krauskopf:1995:AUS

[KVA95] M. Krauskopf, María Inés Vera, and R. Albertini. Assessment of a university’s scientific capabilities and profile: The case of the faculty of biological sciences of the Pontificia Universidad Catolica de Chile. Scientometrics, 34(1):87–100, September 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019175.

Krauskopf:1995:CPS

[KVKWD95] M. Krauskopf, María Inés Vera, Vania Krauskopf, and A. Welljams-Dorof. A citationist perspective on science in Latin America and the Caribbean, 1981–1993. Scientometrics, 34(1):3–25, September 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02019169.

Konrad:1990:STD

[KW90] N. Konrad and D. Wahl. Science, technology and development indicators for third world countries — possibilities for analysis and grouping. Scientometrics, 19(3–4):245–270, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02095351.

Kyriakou:1995:MAS

[Kyr95] D. Kyriakou. Macroeconomic aspects of S/ t programme evaluation. Scientometrics, 34(3):451–459, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018012.

Kyvik:1994:PSP

[Kyv94] S. Kyvik. Popular science publishing. Scientometrics, 31(2): 143–153, October 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018557.

Laredo:1995:SEE

[Lar95] P. Larédo. Structural effects of EC RT&D programmes. Scientometrics, 34(3):473–487, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018014.

Lau:1995:CCS

[Lau95] Tuen-Yu Lau. Chinese communication studies: A citation analysis of Chinese communication research in Englishlanguage journals. Scientometrics, 33(1):65–91, May 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020775.

Lazarev:1994:NDC

[Laz94] V. S. Lazarev. Notion of a document: A center of “gravity attraction” for getting metricians together. Scientometrics, 30(2–3):511–516, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018132.

Lazarev:1996:CBT

[Laz96] V. S. Lazarev. On chaos in bibliometric terminology. Scientometrics, 35(2):271–277, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018485.

Lancaster:1992:ULE

[LBR92] F. W. Lancaster, R. H. Burger, and B. M. Rauchfuss. Use of literature by East European scientists: What influences place of publication of sources cited? Scientometrics, 24(3):419– 439, July 1992. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02051039.

Lewison:1991:BSE

[LC91] G. Lewison and P. Cunningham. Bibliometric studies for the evaluation of trans-national research. Scientometrics, 21(2): 223–244, June 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017570.

Leydesdorff:1993:DST

[LC93] L. Leydesdorff and Susan E. Cozzens. The delineation of specialties in terms of journals using the dynamic journal set of the SCI. Scientometrics, 26(1):135–156, January 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016797.

LeMarc:1991:DRP

[LCS^{+}91] M. Le Marc, J.-P. Courtial, E. Drozda Senkovska, J.-P. Petard, and Y. Py. The dynamics of research in the psychology of work from 1973 to 1987: From the study of companies to the study of professions. Scientometrics, 21(1):69–86, May 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019183.

Leta:1996:PSB

[LD96] Jacqueline Leta and L. De Meis. A profile of science in Brazil. Scientometrics, 35(1):33–44, January 1996. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02018231.

Lewison:1998:EFO

[LD98] G. Lewison and G. Dawson. The effect of funding on the outputs of biomedical research. Scientometrics, 41(1–2):17–27, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02457963.

LePair:1995:SOP

[Le 95] Cornelius Le Pair. Speech on the occasion of the presentation of the 1995 Derek de Solla Price Award to prof. Dr. A. F. J. Van Raan at the ISSI Conference held at River Forest, Illinois, on June 10, 1995. Scientometrics, 34(2):VII–X, October 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02020419.

LePair:1999:SOP

[Le 99] C. Le Pair. Speech on the occasion of the presentation of the 1999 Derek de Solla Price Award to Drs. Wolfgang Glänzel and Henk Moed at the ISSI conference held at Colima, Mexico. Scientometrics, 46(1):5–9, September 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02766291.

Leclerc:1998:STM

[Lec98] Michel Leclerc. Science et technologie — le mariage japonais. (French) [science and technology — the Japanese marriage]. Scientometrics, 42(3):429–434, October 1998. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02458381.

Ledoux:1999:MIE

[Led99] M. J. Ledoux. Measuring the impact of the EU framework programme. Scientometrics, 45(3):505–506, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02457618.

Lemoine:1992:FDR

[Lem92a] W. Lemoine. The frequency distribution of research papers and patents according to sex. The case of CSIR. India. Scientometrics, 24(3):449–469, July 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02051041.

Lemoine:1992:PPM

[Lem92b] W. Lemoine. Productivity patterns of men and women scientists in Venezuela. Scientometrics, 24(2):281–295, June 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017912.

Lewison:1991:SOE

[Lew91] G. Lewison. The scientific output of the EC’s less favoured regions. Scientometrics, 21(3):383–402, July/August 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093977.

Lewison:1993:CEC

[Lew93] G. Lewison. The contribution of European community less favoured region research outputs to economic and social cohesion. Scientometrics, 28(2):217–229, October 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016901.

Lewison:1994:PEC

[Lew94] G. Lewison. Publications from the European community’s biotechnology action programme (BAP): Multinationality, acknowledgement of support, and citations. Scientometrics, 31 (2):125–142, October 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018556.

Lewison:1996:FOS

[Lew96] G. Lewison. The frequencies of occurrence of scientific papers with authors of each initial letter and their variation with nationality. Scientometrics, 37(3):401–416, November 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019255.

Lewison:1998:NBT

[Lew98] G. Lewison. New bibliometric techniques for the evaluation of medical schools. Scientometrics, 41(1–2):5–16, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457962.

Lewison:1999:DCB

[Lew99] G. Lewison. The definition and calibration of biomedical subfields. Scientometrics, 46(3):529–537, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459609.

Leydesdorff:1990:PSI

[Ley90a] L. Leydesdorff. The prediction of science indicators using information theory. Scientometrics, 19(3–4):297–324, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02095353.

Leydesdorff:1990:RASa

[Ley90b] L. Leydesdorff. Relations among science indicators or more generally among anything one might wish to count about texts. Scientometrics, 18(3–4):281–307, March 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017766.

Leydesdorff:1990:RASb

[Ley90c] L. Leydesdorff. Relations among science indicators or more generally among anything one might wish to count about texts. Scientometrics, 19(3–4):271–296, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02095352.

Leydesdorff:1991:SDB

[Ley91] L. Leydesdorff. On the “scientometric decline” of British science. One additional graph in reply to Ben Martin. Scientometrics, 20(2):363–367, February 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/BF02017526. See

[Mar91].

Leydesdorff:1992:IST

[Ley92a] L. Leydesdorff. Irreversibilites in science and technology networks: An empirical and analytical approach. Scientometrics, 24(2):321–357, June 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017914.

Leydesdorff:1992:RCC

[Ley92b] L. Leydesdorff. A reply to Courtial’s comments. Scientometrics, 25(2):317–319, October 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/BF02028089. See

[Ley92c, Cou92].

Leydesdorff:1992:VSL

[Ley92c] L. Leydesdorff. A validation study of “Leximappe”. Scientometrics, 25(2):295–312, October 1992. CODEN SCNTDX.

ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/BF02028087. See

comments [Cou92] and reply [Ley92b].

Leydesdorff:1994:BR

[Ley94a] L. Leydesdorff. Book review. Scientometrics, 30(2–3):539– 541, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02018137.

Leydesdorff:1994:GAJ

[Ley94b] L. Leydesdorff. The generation of aggregated journal–journal citation maps on the basis of the CD–ROM version of the Science Citation Index. Scientometrics, 31(1):59–84, September 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018102.

Leydesdorff:1998:TC

[Ley98] L. Leydesdorff. Theories of citation? Scientometrics, 43(1): 5–25, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458391.

Lange:1999:GSR

[LF99] Lydia L. Lange and P. A. Frensch. Gaining scientific recognition by position: Does editorship increase citation rates? Scientometrics, 44(3):459–486, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458490.

Lewison:1993:LAS

[LFJK93] G. Lewison, A. Fawcett-Jones, and C. Kessler. Latin American scientific output 1986–91 and international coauthorship patterns. Scientometrics, 27(3):317–336, July 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016945.

Leclerc:1994:ISC

[LG94] M. Leclerc and J. Gagné. International scientific cooperation: The continentalization of science. Scientometrics, 31(3):261– 292, November 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016876.

Lewison:1999:YPE

[LI99] G. Lewison and R. Igic. Yogoslav politics, “ethnic cleansing” and co-authorship in science. Scientometrics, 44(2):183–192, February 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02457379.

Lindsey:1991:PMR

[Lin91a] D. Lindsey. Precision in the manuscript review process: Hargens and Herting revisited. Scientometrics, 22(2):313–325, October 1991. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02020004. See [HH90].

Lindsey:1991:RBP

[Lin91b] D. Lindsey. The relationship between performance indicators for academic research and funding: Developing a measure of return on investment in science. Scientometrics, 20(1): 221–234, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018156.

Lin:1996:ESN

[Lin96a] Yang Lin. Empirical studies of negative political advertising: A quantitative review using a method of combined citation and content analysis. Scientometrics, 37(3):385–399, November 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019254.

Link:1996:EPM

[Lin96b] A. N. Link. Economic performance measures for evaluating government-sponsored research. Scientometrics, 36(3):325– 342, July/August 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02129598.

Lindqvist:1999:QAS

[Lin99] O. V. Lindqvist. Quality assessment and structural change in universities. Scientometrics, 45(3):367–370, July 1999.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457591.

List:1999:AFE

[Lis99] H. List. Advisory functions at European level. Scientometrics, 45(3):479–484, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457612.

Liming:1993:SPA

[LL93] L. Liming and L. Lihua. Scientific publication activities of 32 countries — Zipf–Pareto distribution. Scientometrics, 26(2): 263–273, February 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016219.

Lancaster:1990:DPP

[LLD90] F. W. Lancaster, Sun-Yoon Kim Lee, and Catalina Diluvio. Does place of publication influence citation behavior? Scientometrics, 19(3–4):239–244, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02095350.

Leta:1998:HRS

[LLD98] Jacqueline Leta, Denise Lannes, and L. De Meis. Human resources and scientific productivity in Brazil. Scientometrics, 41(3):313–324, March 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02459048.

Luwel:1998:PDS

[LM98] M. Luwel and H. F. Moed. Publication delays in the science field and their relationship to the ageing of scientific literature. Scientometrics, 41(1–2):29–40, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457964.

leMinor:1991:BSP

[lMD91] Sylvaine le Minor and Paulette Dostatni. A bibliometric study of the publications of the French National Institute for Health and Medical Research (INSERM). Scientometrics, 22(1):41– 63, September 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019274.

Li-Ming:1996:DMS

[LMHZYYS96] Liang Li-Ming, Zhao Hong-Zhou, Wang Yuan, and Wu Yi-Shan. Distribution of major scientific and technological achievements in terms of age group-Weibull distribution. Scientometrics, 36(1):3–18, May 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02126642.

Leemans:1992:NBD

[LMRR92] Marie-Jeanne Leemans, Marleen Maes, R. Rousseau, and Christel Ruts. The negative binomial distribution as a trend distribution for circulation data in Flemish public libraries. Scientometrics, 25(1):47–57, September 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016846.

Lee:1999:BRR

[LOK99] M. Lee, K. Om, and J. Koh. Blind review of research proposals in Korea: Its effectiveness and factors affecting applicant detection. Scientometrics, 45(1):17–31, May 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458466.

lePair:1994:PML

[lP94] C. le Pair. Puberty or mid-life crisis. Scientometrics, 30(2– 3):517–520, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018133.

Logan:1991:BAC

[LS91] Elisabeth L. Logan and W. M. Shaw, Jr. A bibliometric analysis of collaboration in a medical specialty. Scientometrics, 20 (3):417–426, March 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019762.

Liu:1995:SAA

[LS95] Jinming Liu and Sizhu Shu. Statistical analysis of astronomical papers of China during 1986–1990. Scientometrics, 32(3): 237–245, March 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017643.

Luukkonen:1993:MIS

[LTPS93] Terttu Luukkonen, R. J. W. Tijssen, O. Persson, and G. Sivertsen. The measurement of international scientific collaboration. Scientometrics, 28(1):15–36, September 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.

1007/BF02016282. See comments [Oku93].

Luukkonen:1990:PSA

[Luu90] Terttu Luukkonen. Publication structures and accumulative advantages. Scientometrics, 19(3–4):167–184, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02095345.

Luukkonen:1992:SPB

[Luu92] Terttu Luukkonen. Is scientists’ publishing behaviour reward seeking? Scientometrics, 24(2):297–319, June 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017913.

Luukkonen:1994:WLG

[Luu94] Terttu Luukkonen. Are we longing for the golden era lost or for the one to come? Scientometrics, 30(2–3):481–485, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018128.

Luukkonen:1997:WLT

[Luu97] Terttu Luukkonen. Why has Latour’s theory of citations been ignored by the bibliometric community? Discussion of sociological interpretations of citation analysis. Scientometrics, 38 (1):27–37, January 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02461121.

Luwel:1999:SCI

[Luw99] M. Luwel. Is the science citation index US-biased? Scientometrics, 46(3):549–562, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02459611.

Leydesdorff:1997:SCT

[LV97] L. Leydesdorff and P. Van Den Besselaar. Scientometrics and communication theory: Towards theoretically informed indicators. Scientometrics, 38(1):155–174, January 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02461129.

Law:1992:MAR

[LW92] J. Law and J. Whittaker. Mapping acidification research: A test of the co-word method. Scientometrics, 23(3):417–461, March 1992. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02029807.

Leydesdorff:1994:CC

[LW94] L. Leydesdorff and P. Wouters. Crisis or critique? Scientometrics, 30(2–3):433–437, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018119.

Leydesdorff:1999:BTC

[LW99] L. Leydesdorff and P. Wouters. Between texts and contexts: Advances in theories of citation? (a rejoinder). Scientometrics, 44(2):169–182, February 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02457378.

Lyon:1990:DRC

[Lyo90] W. S. Lyon. The dreams of reason. The computer and the rise of the science of complexity. Scientometrics, 18(3–4): 321–322, March 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02017768.

Mely:1998:SPC

[MADO98] B. Mély, M. Abd El Kader, G. Dudognon, and Y. Okubo. Scientific publications of China in 1994: Evolution or revolution? Scientometrics, 42(1):3–16, May 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02465009.

Makino:1998:PRG

[Mak98] Junichiro Makino. Productivity of research groups-relation between citation analysis and reputation within research communities. Scientometrics, 43(1):87–93, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458398.

Mannerkoski:1999:AFE

[Man99] M. Mannerkoski. Advisory functions at European level. Scientometrics, 45(3):487–489, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457614.

Martin:1991:BAU

[Mar91] B. R. Martin. The bibliometric assessment of UK scientific performance: a reply to Braun, Glänzel and Schubert. Scientometrics, 20(2):333–357, February 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com/article/10.1007/BF02017524. See

[Ley91].

Martin:1992:BHB

[Mar92] Ben R. Martin. ‘Big history’ for big science? Critical review of history of CERN. Scientometrics, 24(2):371, June 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02017916.

Martin:1994:BSR

[Mar94] B. R. Martin. British science in the 1980s — has the relative decline continued? Scientometrics, 29(1):27–56, January 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.

1007/BF02018382. See report [Abt94].

Martin:1996:UMI

[Mar96] B. R. Martin. The use of multiple indicators in the assessment of basic research. Scientometrics, 36(3):343–362, July/August 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02129599.

Massimo:1991:UIR

[Mas91] L. Massimo. The use of indicators in the R&D evaluation activity of the European communities. Scientometrics, 21 (3):255–262, July/August 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02093969.

Matricciani:1994:SEM

[Mat94] E. Matricciani. Shannon’s entropy as a measure of the “life” of the literature of a discipline. Scientometrics, 30(1):129– 145, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02017218.

Mauguin:1991:UCD

[Mau91] P. Mauguin. Using a contracts database for evaluating the dynamics of a technological program: The case of the European ‘non-nuclear energy’ program. Scientometrics, 22(1):207–228, September 1991. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02019281.

Makovetskaya:1994:SII

[MB94] O. Makovetskaya and V. Bernadsky. Scientometric indicators for identification of technology system life cycle phase. Scientometrics, 30(1):105–116, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02017216.

Macias-Chapula:1992:PSC

[MC92] C. A. Macías-Chapula. Patterns of scientific communication among Latin American countries, in the field of medical education. Scientometrics, 23(1):123–135, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020918.

Macias-Chapula:1994:NSS

[MC94] C. A. Macias-Chapula. Non-SCI subject visibility of the Latin American scientific production in the health field. Scientometrics, 30(1):97–104, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017215.

Macias-Chapula:1995:PHC

[MC95] C. A. Macias-Chapula. Primary health care in Mexico: A “non-ISI” bibliometric analysis. Scientometrics, 34(1):63–71, September 1995. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02019173.

McCain:1995:SBR

[McC95] Katherine W. McCain. The structure of biotechnology R&D. Scientometrics, 32(2):153–175, February 1995. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02016892.

McCain:1998:NNR

[McC98] Katherine W. McCain. Neural networks research in context: A longitudinal journal cocitation analysis of an emerging interdisciplinary field. Scientometrics, 41(3):389–410, March 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459053.

McGrath:1994:LSB

[McG94] W. E. McGrath. Little scientometrics, big scientometrics … and beyond. Scientometrics, 30(2–3):439–442, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018120.

McGrath:1996:UAO

[McG96] W. E. McGrath. The unit of analysis (objects of study) in bibliometrics and scientometrics. Scientometrics, 35(2):257– 264, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018483.

Meyer:1995:IOC

[MCGC95] J.-B. Meyer, J. Charum, J. Granés, and Y. Chatelin. Is it opened or closed? Colombian science on the move. Scientometrics, 34(1):73–86, September 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02019174.

Macias-Chapula:1997:SCM

[MCRC97] C. A. Macías-Chapula and I. P. Rodea-Castro. Subject content of the Mexican production on health and the environment (1982–1993). Scientometrics, 38(2):295–308, February 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457415.

Macias-Chapula:1998:BAA

[MCRCNB98] C. A. Macias-Chapula, Irma P. Rodea-Castro, and Nora Narvaez-Berthelemot. Bibliometric analysis of aids literature in Latin America and the Caribbean. Scientometrics, 41(1– 2):41–49, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457965.

Macias-Chapula:1999:SCA

[MCSAMSL99] C. A. Macias-Chapula, G. R. Sotolongo-Aguilar, B. Magde, and J. Solorio-Lagunas. Subject content analysis of aids literature, as produced in Latin America and the Caribbean. Scientometrics, 46(3):563–574, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02459612.

Moed:1991:ISC

[MDNT91] H. F. Moed, R. E. De Bruin, A. J. Nederhof, and R. J. W. Tijssen. International scientific co-operation and awareness within the European community: Problems and perspectives. Scientometrics, 21(3):291–311, July/August 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093972.

Moed:1995:NBT

[MDV95] H. F. Moed, R. E. De Bruin, and Th. N. Van Leeuwen. New bibliometric tools for the assessment of national research performance: Database description, overview of indicators and first applications. Scientometrics, 33(3):381–422, July 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017338.

Meadows:1991:QSF

[Mea91] A. J. Meadows. Quantitative study of factors affecting the selection and presentation of scientific material to the general public. Scientometrics, 20(1):113–119, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018149.

Meadows:1994:LSB

[Mea94] A. J. Meadows. Little scientometrics, big scientometrics … and beyond. Scientometrics, 30(2–3):447–449, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018122.

Melin:1996:NU

[Mel96] G. Melin. The networking university. Scientometrics, 35(1): 15–31, January 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018230.

Melin:1999:INS

[Mel99] G. Melin. Impact of national size on research collaboration. Scientometrics, 46(1):161–170, September 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02766300.

Meneghini:1992:BPB

[Men92] R. Meneghini. Brazilian production in biochemistry. The question of international versus domestic publication. Scientometrics, 23(1):21–30, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02020911.

Mendez:1994:TAS

[Mén94] Aida Méndez. Thinking about scientometrics. Scientometrics, 30(2–3):393–395, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018110.

Meneghini:1996:KRC

[Men96] R. Meneghini. The key role of collaborative work in the growth of Brazilian science in the last ten years. Scientometrics, 35 (3):367–373, March 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016907.

Mep:1995:IS

[Mep95] Carlos Robles Piquer Mep. Invited speech. Scientometrics, 34(3):321–323, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018002.

Meyer:1999:RAB

[Mey99] Verena Meyer. The role of advisory bodies in evaluation. Scientometrics, 45(3):381–385, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457594.

Meske:1990:SDS

[MF90] W. Meske and M. C. Fernández De Alaiza. Structure and development of the scientific and technological potential in the Republic of Cuba. Scientometrics, 18(1–2):137–155, January 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019166.

Markusova:1991:HCS

[MG91] Valentina A. Markusova and B. C. Griffith. Highly cited Soviet journals in the physical and life sciences: A study of the function of journals. Scientometrics, 21(1):99–113, May 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019185.

Mendez:1992:CRS

[MG92] Aida Méndez and Isabel Gómez. Collaborative research in Spain in the field of pharmacy and pharmacology. Scientometrics, 24(1):137–147, May 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02026477.

Milman:1993:ACC

[MG93] B. L. Milman and Yu. A. Gavrilova. Analysis of citation and co-citation in chemical engineering. Scientometrics, 27(1):53– 74, May 1993. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02017755.

Molas-Gallart:1999:MDR

[MG99] J. Molas-Gallart. Measuring defence R&D: A note on problems and shortcomings. Scientometrics, 45(1):3–16, May 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458465.

Mendez:1993:SIA

[MGB93] Aida Méndez, Isabel Gómez, and Maria Bordons. Some indicators for assessing research performance without citations. Scientometrics, 26(1):157–167, January 1993. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02016798.

Markusova:1996:IBR

[MGCG96] Valentina A. Markusova, R. S. Gilyarevskii, A. I. Chernyi, and Belver C. Griffith. Information behavior of Russian scientists in the “perestroika” period. Scientometrics, 37(2):361–380, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02093630.

Marton:1998:IBE

[MHZ98] J. Marton, H. Hulesch, and I. Zallár. Intensity breeds effectivity. Scientometrics, 41(3):411–415, March 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459054.

Michalos:1995:PIF

[Mic95] A. C. Michalos. Prices and impact factors of interdisciplinary social science journals. Scientometrics, 32(3):259–261, March 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017645.

Milman:1994:ICC

[Mil94] B. L. Milman. Individual co-citation clusters as nuclei of complete and dynamic informetric models of scientific and technological areas. Scientometrics, 31(1):45–57, September 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018101.

Miquel:1994:LSB

[Miq94] J. F. Miquel. Little scientometrics, big scientometrics … and beyond. Scientometrics, 30(2–3):443–445, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018121.

Maclean:1993:PPI

[MJ93] J. Maclean and Carmela Janagap. The publication productivity of international agricultural research centers. Scientometrics, 28(3):329–348, November/December 1993. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02026514.

Meyer-Krahmer:1999:QAS

[MK99] F. Meyer-Krahmer. Quantitative approaches — strengths and weaknesses. Scientometrics, 45(3):413–415, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457600.

Moed:1998:ECF

[MLH^{+}98] H. F. Moed, M. Luwel, J. A. Houben, E. Spruyt, and H. Van Den Berghe. The effects of changes in the funding structure of the Flemish universities on their research capacity, productivity and impact during the 1980’s and early 1990’s. Scientometrics, 43(2):231–255, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458409.

Mychko-Megrin:1990:EAT

[MM90] A. Y. Mychko-Megrin. Estimates of the annual total number of titles on medicine and its disciplines and scientific productivity of physicians. Scientometrics, 18(5–6):375–388, May 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020152.

MacRoberts:1996:PCA

[MM96a] M. H. MacRoberts and Barbara R. MacRoberts. Problems of citation analysis. Scientometrics, 36(3):435–444, July/August 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02129604.

Miller:1996:BIC

[MM96b] R. Miller and A. Manseau. Bibliometric indicators and the competitive environment of R&D laboratories. Scientometrics, 36(3):421–433, July/August 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02129603.

Miquel:1994:SIC

[MO94] J. F. Miquel and Y. Okubo. Structure of international collaboration in science-part II: Comparisons of profiles in countries using a link indicator. Scientometrics, 29(2):271–297, February 1994. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02017977.

Moed:1996:DCS

[Moe96a] H. F. Moed. Differences in the construction of sci based bibliometric indicators among various producers: A first over view. Scientometrics, 35(2):177–191, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018476.

Moed:1996:VBI

[Moe96b] J. C. Korevaar H. F. Moed. Validation of bibliometric indicators in the field of mathematics. Scientometrics, 37(1):117– 130, September 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02093488.

Moed:1999:I

[Moe99] H. F. Moed. Introduction. Scientometrics, 44(3):319–321, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02458482.

Miquel:1995:WSD

[MOO^{+}95] J. F. Miquel, T. Ojasoo, Y. Okubo, A. Paul, and J. C. Doré. World science in 18 disciplinary areas: Comparative evaluation of the publication patterns of 48 countries over the period 1981–1992. Scientometrics, 33(2):149–167, June 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020566.

Melin:1996:SRC

[MP96] G. Melin and O. Persson. Studying research collaboration using co-authorships. Scientometrics, 36(3):363–377, July/ August 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02129600.

Meyer:1998:NIP

[MP98] M. Meyer and O. Persson. Nanotechnology-interdisciplinarity, patterns of collaboration and differences in application. Scientometrics, 42(2):195–205, June 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458355.

Mendez:1992:ASI

[MS92] Aida Méndez and P. Salvador. The application of scientometric indicators to the Spanish Scientific Research Council. Scientometrics, 24(1):61–78, May 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02026473.

Martens:1993:CCB

[MS93] B. Martens and T. Saretzki. Conferences and courses on biotechnology. Describing scientific communication by exploratory methods. Scientometrics, 27(3):237–260, July 1993.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016941.

Martens:1994:QAT

[MS94] B. Martens and T. Saretzki. Quantitative analysis of thematic structures in the field of biotechnology: A study on the basis of conference data. Scientometrics, 30(1):117–128, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017217.

Magri:1996:SJC

[MS96a] Marie-Hélène Magri and Aline Solari. The SCI Journal Citation Reports: A potential tool for studying journals? Scientometrics, 35(1):93–117, January 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018235.

Marshakova-Shaikevich:1996:SIF

[MS96b] Irina Marshakova-Shaikevich. The standard impact factor as an evaluation tool of science fields and scientific journals. Scientometrics, 35(2):283–290, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018487.

Moed:1996:CAJ

[MVR96] H. F. Moed, Th. N. Van Leeuwen, and J. Reedijk. A critical analysis of the journal impact factors of Angewandte Chemie and the Journal of the American Chemical Society inaccuracies in published impact factors based on overall citations only. Scientometrics, 37(1):105–116, September 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093487.

Moed:1999:TAI

[MVR99] H. F. Moed, Th. N. Van Leeuwen, and J. Reedijk. Towards appropriate indicators of journal impact. Scientometrics, 46(3): 575–589, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459613.

Mulford:1993:VAA

[MWHG93] C. L. Mulford, Lisa Waldner-Haugrud, and Hemchandra Gajbhiye. Variables associated with agricultural scientists’ work alienation and publication productivity. Scientometrics, 27 (3):261–282, July 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016942.

Maczelka:1992:AWI

[MZ92] Hajnalka Maczelka and S. Zsindely. All well if starts well? Citation infancy of recently launched chemistry journals. Scientometrics, 25(2):367–372, October 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02028092.

Nagpaul:1995:CIU

[Nag95a] P. S. Nagpaul. Contribution of Indian universities to the mainstream scientific literature: A bibliometric assessment. Scientometrics, 32(1):11–36, January 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020186.

Nagpaul:1995:QQM

[Nag95b] P. S. Nagpaul. Quasi-quantitative measures of research performance: An assessment of construct validity and reliability. Scientometrics, 33(2):169–185, June 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020567.

Nagpaul:1999:TLI

[Nag99] P. S. Nagpaul. Transnational linkages of Indian science: A structural analysis. Scientometrics, 46(1):109–140, September 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02766298.

Nalimov:1991:MXC

[Nal91] V. V. Nalimov. Meeting the XXIth century. Scientometrics, 20 (1):65–69, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018144.

Narin:1994:PB

[Nar94] F. Narin. Patent bibliometrics. Scientometrics, 30(1):147– 155, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02017219.

Narin:1995:PIE

[Nar95] F. Narin. Patents as indicators for the evaluation of industrial research output. Scientometrics, 34(3):489–496, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018015.

Narvaez-Berthelemot:1995:IMI

[NB95] N. Narvaez-Berthelemot. An index to measure the international collaboration of developing countries based on the participation of national institutions: The case of Latin America. Scientometrics, 34(1):37–44, September 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019171.

Narvaez-Berthelemot:1992:ISC

[NBFM92] N. Narvaez-Berthelemot, L. P. Frigoletto, and J. F. Miquel. International scientific collaboration in Latin America. Scientometrics, 24(3):373–392, July 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02051036.

Narin:1996:BPM

[NH96] F. Narin and Kimberly S. Hamilton. Bibliometric performance measures. Scientometrics, 36(3):293–310, July/August 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02129596.

Nasir:1994:BEA

[NHHA94] A. M. Nasir, H. Hassan, K. A. Hamid, and S. S. Agha. Bibliometric evaluation of agricultural literature published in Malaysia. Scientometrics, 29(2):191–217, February 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017973.

Nieminen:1996:TER

[Nie96] P. Nieminen. Type of empirical research reports as an explanatory factor in citation performance of psychiatric research. Scientometrics, 35(3):309–320, March 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016903.

Nederhof:1993:MMP

[NM93] A. J. Nederhof and H. F. Moed. Modeling multinational publication: Development of an on-line fractionation approach to measure national scientific output. Scientometrics, 27(1):39– 52, May 1993. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02017754.

Nederhof:1995:DBI

[NM95] A. J. Nederhof and R. F. Meijer. Development of bibliometric indicators for utility of research to users in society: Measurement of external knowledge transfer via publications in trade journals. Scientometrics, 32(1):37–48, January 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020187.

Nederhof:1993:RPI

[NMMvR93] A. J. Nederhof, R. F. Meijer, H. F. Moed, and A. F. J. van Raan. Research performance indicators for university departments: A study of an agricultural university. Scientometrics, 27(2):157–178, June 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016548.

Noyons:1999:IRP

[NMV99] E. C. M. Noyons, H. F. Moed, and A. F. J. Van Raan. Integrating research performance analysis and science mapping. Scientometrics, 46(3):591–604, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02459614.

Nederhof:1992:AIS

[NN92] A. J. Nederhof and E. C. M. Noyons. Assessment of the international standing of university departments’ research: A comparison of bibliometric methods. Scientometrics, 24(3): 393–404, July 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02051037.

Narin:1998:LBP

[NO98] F. Narin and D. Olivastro. Linkage between patents and papers: An interim EPO/US comparison. Scientometrics, 41(1– 2):51–59, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457966.

Nordstrom:1990:BLR

[Nor90] L. O. Nordstrom. “Bradford’s law” and the relationship between ecology and biogeography. Scientometrics, 18(3–4):193– 203, March 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02017761.

Norris:1993:APC

[Nor93] R. P. Norris. Authorship patterns in CJNR: 1970–1991. Scientometrics, 28(2):151–158, October 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016897.

Nagpaul:1993:CNA

[NP93] P. S. Nagpaul and Neeraj Pant. Cross-national assessment of specialization patterns in chemistry. Scientometrics, 27(2): 215–235, June 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016551.

Nauwelaers:1995:MER

[NR95] Claire Nauwelaers and A. Reid. Methodologies for the evaluation of regional innovation potential. Scientometrics, 34(3): 497–511, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018016.

Nagpaul:1994:ROT

[NS94] P. S. Nagpaul and Lalita Sharma. Research output and transnational cooperation in physics subfields: A multidimensional analysis. Scientometrics, 31(1):97–122, September 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018104.

Nagpaul:1995:SET

[NS95] P. S. Nagpaul and Lalita Sharma. Science in the eighties: A typology of countries based on inter-field priorities. Scientometrics, 34(2):263–283, October 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020424.

Narin:1991:SCO

[NSW91] F. Narin, K. Stevens, and Edith S. Whitlow. Scientific cooperation in Europe and the citation of multinationally authored papers. Scientometrics, 21(3):313–323, July/August 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093973.

Niwa:1996:TGI

[NT96] F. Niwa and H. Tomizawa. A trial of general indicator of science and technology: Methodological study of overall estimation of national S&T activity. Scientometrics, 37(2):245–265, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02093623.

Nunes:1999:RRS

[Nun99] E. D. Nunes. A review of research studies conducted on scientific production in collective health in Brazil. Scientometrics, 44(2):157–167, February 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02457377.

Nederhof:1997:MSB

[NV97] A. J. Nederhof and E. Van Wijk. Mapping the social and behavioral sciences world-wide: Use of maps in portfolio analysis of national research efforts. Scientometrics, 40(2):237–276, October 1997. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02457439.

Noyons:1998:AMS

[NV98] E. C. M. Noyons and A. F. J. Van Raan. Advanced mapping of science and technology. Scientometrics, 41(1–2):61–67, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02457967.

Nederhof:1999:PII

[NV99] A. J. Nederhof and E. Van Wijk. Profiling institutes: Identifying high research performance and social relevance in the social and behavioral sciences. Scientometrics, 44(3):487–506, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458491.

Nicolini:1995:CAU

[NVBG95] C. Nicolini, S. Vakula, M. Italo Balla, and E. Gandini. Can the assignment of university chairs be automated? Scientometrics, 32(2):93–107, February 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016886.

Noyons:1994:BCS

[NvR94] E. C. M. Noyons and A. F. J. van Raan. Bibliometric cartography of scientific and technological developments of an R&D field. Scientometrics, 30(1):157–173, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017220.

Nybom:1999:QAS

[Nyb99] T. Nybom. Quality assessment and structural change in universities. Scientometrics, 45(3):371–377, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457592.

Nemtsov:1998:MMP

[NZ98] A. V. Nemtsov and N. A. Zorin. Mathematical methods in psychiatric papers. Scientometrics, 42(2):121–128, June 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458352.

Oborne:1999:UEI

[Obo99] M. Oborne. Universities — engines of innovation in the information society. Scientometrics, 45(3):557–560, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457628.

Ojasoo:1999:CBM

[OD99] T. Ojasoo and J. C. Doré. Citation bias in medical journals. Scientometrics, 45(1):81–94, May 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458469.

Okubo:1998:MAP

[ODOM98] Y. Okubo, J. C. Dore, T. Ojasoo, and J. F. Miquel. A multivariate analysis of publication trends in the 1980s with special reference to South-East Asia. Scientometrics, 41(3):273–289, March 1998. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02459045.

OHerlihy:1995:RDR

[O’H95] J. O’Herlihy. RT & d, regional development and evaluation. Scientometrics, 34(3):513–518, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018017.

Okubo:1993:CSS

[Oku93] Y. Okubo. Comments on some of the statements in the article: The measurement of international scientific collaboration by T. Luukkonen, R. J. W. Tijsen, O. Persson, G. Sivertsen. Scientometrics, 28(1):37–39, September 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016283. See [LTPS93].

Okubo:1992:SIC

[OMFD92] Y. Okubo, J. F. Miquel, L. Frigoletto, and J. C. Doré. Structure of international collaboration in science: Typology of countries through multivariate techniques using a link indicator. Scientometrics, 25(2):321–351, October 1992.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02028090.

ONeill:1998:APT

[O’N98] G. Patrick O’Neill. Authorship patterns in theory based versus research based journals. Scientometrics, 41(3):291–298, March 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459046.

Ortega:1992:SST

[OPMU92] Concepción Ortega, L. M. Plaza, M. J. Martin, and M. C. Urdin. Spanish scientific and technical journals. State of the art. Scientometrics, 24(1):21–42, May 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02026471.

Oppenheim:1997:PCA

[Opp97] C. Oppenheim. Patent citation analysis. Scientometrics, 39 (1):141, May 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02457435.

Over:1990:SIA

[Ove90] R. Over. The scholarly impact of articles published by men and women in psychology journals. Scientometrics, 18(5–6): 331–340, May 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020149.

Osareh:1997:TWC

[OW97] Farideh Osareh and Concepción S. Wilson. Third world countries (TWC) research publications by disciplines: A countryby-country citation analysis. Scientometrics, 39(3):253–266, July 1997. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02458529.

Pianta:1991:SSS

[PA91] M. Pianta and Daniele Archibugi. Specialization and size of scientific activities: A bibliometric analysis of advanced countries. Scientometrics, 22(3):341–358, November 1991.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019767.

Pao:1991:RFR

[Pao91] Miranda Lee Pao. On the relationship of funding and research publications. Scientometrics, 20(1):257–281, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018158.

Pao:1992:RFR

[Pao92] Miranda Lee Pao. On the relationship of funding and research publications. Scientometrics, 24(1):179, May 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/BF02026480.pdf.

Papadopoulos:1996:EIR

[Pap96] S. Papadopoulos. Evaluation of industrial research. Scientometrics, 37(1):143–151, September 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02093490.

Papon:1999:RNA

[Pap99] P. Papon. The role of national agencies in evaluation. Scientometrics, 45(3):391–399, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02457596.

Pavitt:1997:CJI

[Pav97] Keith Pavitt. Comments on John Irvine and Ben R. Martin, recipients of the 1997 Derek de Solla Price Award. Scientometrics, 40(3):363–366, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02459285. See [Ano97a].

Persson:1995:LNI

[PB95] O. Persson and M. Beckmann. Locating the network of interacting authors in scientific specialties. Scientometrics, 33 (3):351–366, July 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017336.

PereiraFriedrich:1998:LSB

[PD98] Margarete Pereira Friedrich and P. Dos Santos Rodrigues. Looking at science in Brazilian universities: The case of the instituto de Biofísica carlos chagas filho. Scientometrics, 42 (2):247–258, June 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458358.

Pereira:1999:SBH

[PE99] Julio C. R. Pereira and Maria Mercedes L. Escuder. The scenario of Brazilian health sciences in the period of 1981 to 1995. Scientometrics, 45(1):95–105, May 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458470.

Peritz:1990:BDB

[Per90a] Bluma C. Peritz. A Bradford distribution for bibliometrics. Scientometrics, 18(5–6):323–329, May 1990. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02020148.

Peritz:1990:CIF

[Per90b] Bluma C. Peritz. The citation impact of funded and unfunded research in economics. Scientometrics, 19(3–4):199– 206, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02095347.

Peritz:1991:CIL

[Per91] Bluma C. Peritz. The citation impact of letters to the editor; the case of lancet. Scientometrics, 20(1):121–129, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018150.

Peritz:1994:LSB

[Per94a] Bluma C. Peritz. Little scientometrics, big scientometrics … and beyond. Scientometrics, 30(2–3):415–418, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018115.

Peritz:1994:HVS

[Per94b] Bluma C. Peritz. On the heuristic value of scientific publications and their design; a citation analysis of some clinical trials. Scientometrics, 30(1):175–186, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02017221.

Peritz:1997:OAC

[Per97] Bluma Peritz. From the opening address of the conference. Scientometrics, 40(3):367–368, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02459286.

Pestana:1992:SPL

[Pes92] Angel Pestaña. Spanish performance in life sciences. A comparative appraisal of the scientific production of Spain and five other European countries in 1989. Scientometrics, 24 (1):95–114, May 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02026475.

Pereira:1998:BSG

[PEZ98] Julio C. R. Pereira, Maria Mercedes L. Escuder, and Dirce Maria Trevisan Zanetta. Brazilian sciences and government funding at the State of São Paulo. Scientometrics, 43(2): 177–188, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458406.

Pudovkin:1995:IJC

[PF95] A. I. Pudovkin and Elizabeth A. Fuseler. Indices of journal citation relatedness and citation relationships among aquatic biology journals. Scientometrics, 32(3):227–236, March 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017642.

Pfetsch:1990:MCS

[Pfe90] F. R. Pfetsch. The measurement of a country’s scientific and technological potential. Scientometrics, 19(5–6):495–504, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02020710.

Polanco:1998:ANN

[PFK98] X. Polanco, Claire François, and J.-P. Keim. Artificial neural network technology for the classification and cartography of scientific and technical information. Scientometrics, 41(1–2): 69–82, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457968.

Phelan:1999:CIC

[Phe99] T. J. Phelan. A compendium of issues for citation analysis. Scientometrics, 45(1):117–136, May 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458472.

Philips:1992:PMB

[Phi92] R. Philips. Pseudo-metrics on bibliographic entities. Scientometrics, 25(1):193–199, September 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016854.

Pichappan:1993:IMJ

[Pic93] P. Pichappan. Identification of mainstream journals of science speciality: A method using the discipline-contribution score. Scientometrics, 27(2):179–193, June 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016549.

Pichappan:1995:DRJ

[Pic95] P. Pichappan. A dual refinement of journal self-citation measures. Scientometrics, 33(1):13–21, May 1995. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02020772.

Plaza:1998:UMD

[Pla98] Luis M. Plaza. The use of multiple databases in the assesment of research. An application in the field of plant science. Scientometrics, 43(2):299–304, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458412.

Plomp:1990:SNH

[Plo90] R. Plomp. The significance of the number of highly cited papers as an indicator of scientific prolificacy. Scientometrics, 19(3–4):185–197, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02095346.

Plomp:1994:HCP

[Plo94] R. Plomp. The highly cited papers of professors as an indicator of a research group’s scientific performance. Scientometrics, 29 (3):377–393, March 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02033446.

Persson:1996:EGI

[PM96] O. Persson and G. Melin. Equalization, growth and integration of science. Scientometrics, 37(1):153–157, September 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093491.

Peschel:1990:EGM

[PMA90] M. Peschel, W. Mende, and K. F. Albrecht. The evolon growth model: Possible scientometrical evaluations. Scientometrics, 18(1–2):123–136, January 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02019165.

Persson:1997:RCN

[PMDK97] O. Persson, G. Melin, R. Danell, and A. Kaloudis. Research collaboration at Nordic universities. Scientometrics, 39(2): 209–223, June 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457449.

Plaza:1996:SRB

[PMR96] L. M. Plaza, M. J. Martín, and J. Rey. Scientific relations between Spain and Central–Eastern European countries for the period 1982–1992. Scientometrics, 37(1):131–142, September 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093489.

Provost:1992:MOD

[PN92] F. Provost and P. Nieuwenhuysen. Measuring overlap of data bases in water supply and sanitation using sampling and the binomial probability distribution. Scientometrics, 25(1):201– 209, September 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016855.

Pompidou:1999:QAS

[Pom99] A. Pompidou. Quality assessment and structural change in universities. Scientometrics, 45(3):379, July 1999. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02457593.

Pouris:1991:IAS

[Pou91] A. Pouris. Identifying areas of strength in South African technology. Scientometrics, 21(1):23–35, May 1991. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02019180.

Pouris:1992:ESR

[Pou92] A. Pouris. Economic sanctions and R&D. Scientometrics, 25(3):415–424, November 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016929.

Pouris:1993:ESS

[Pou93] A. Pouris. Economies of scale in science & technology agencies. Scientometrics, 28(3):401–406, November/December 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02026518.

Pouris:1994:SC

[Pou94] A. Pouris. Is scientometrics in a crisis? Scientometrics, 30(2– 3):397–399, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018111.

Pravdic:1991:DSP

[POV91] Nevenka Pravdic and Vesna Oluic-Vukovic. Distribution of scientific productivity: Ambiguities in the assignment of author rank. Scientometrics, 20(1):131–144, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018151.

Praestgaard:1999:MEE

[Pra99] E. Praestgaard. Meta-evaluation: Evaluation of evaluations. Scientometrics, 45(3):531–532, July 1999. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02457623.

Prpic:1994:SCF

[Prp94] Katarina Prpić. The socio-cognitive frameworks of scientific productivity. Scientometrics, 31(3):293–311, November 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016877.

Prpic:1996:CDE

[Prp96a] Katarina Prpić. Characteristics and determinants of eminent scientists’ productivity. Scientometrics, 36(2):185–206, June 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017313.

Prpic:1996:SFE

[Prp96b] Katarina Prpić. Scientific fields and eminent scientists’ productivity patterns and factors. Scientometrics, 37(3):445–471, November 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02019258.

Prpic:1998:SES

[Prp98] Katarina Prpić. Science ethics: A study of eminent scientists’ professional values. Scientometrics, 43(2):269–298, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458411.

Parthey:1991:DPI

[PS91] H. Parthey and W. Schuetze. Distribution of publications as an indicator for the evaluation of scientific programs. Scientometrics, 21(3):459–464, July/August 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093981.

Peters:1991:SSA

[PV91] H. P. F. Peters and A. F. J. Van Raan. Structuring scientific activities by co-author analysis. Scientometrics, 20(1): 235–255, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018157.

Peters:1994:BPT

[PvR94] H. P. F. Peters and A. F. J. van Raan. A bibliometric profile of top-scientists a case study in chemical engineering. Scientometrics, 29(1):115–136, January 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018386.

Quesada-Allue:1995:SOA

[QAG95] L. A. Quesada-Allué and D. S. Gitlin. Scientific output in Argentina 1966–1983. Scientometrics, 34(1):27–35, September 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019170.

Quoniam:1998:BLU

[QBR^{+}98] L. Quoniam, F. Balme, H. Rostaing, E. Giraud, and J. M. Dou. Bibliometric law used for information retrieval. Scientometrics, 41(1–2):83–91, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457969.

Qin:1994:IRC

[Qin94] J. Qin. An investigation of research collaboration in the sciences through the philosophical transactions 1901– 1991. Scientometrics, 29(2):219–238, February 1994. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02017974.

Qiu:1990:CID

[QT90] Liwen Qiu and Jean Tague. Complete or incomplete data sets. The Groos droop investigated. Scientometrics, 19(3–4):223– 237, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02095349.

Qurashi:1991:PRS

[Qur91] M. M. Qurashi. Publication-rate and size of two prolific research groups in Departments of Inorganic Chemistry at Dacca University (1944–1965) and Zoology at Karachi University (1966–84). Scientometrics, 20(1):79–92, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018146.

Qurashi:1993:DPR

[Qur93] M. M. Qurashi. Dependence of publication-rate on size of some university groups and departments in U.K. and Greece in comparison with N.C.I., USA. Scientometrics, 27(1):19–38, May 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017753.

Qihao:1990:AJO

[QZ90] Miao Qihao and Zhang Zuozhi. Anatomy of JETRO’s overseas technology monitoring: Bibliometrical and content analysis. Scientometrics, 19(1–2):75–90, July 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02130466.

Rogers:1993:NAD

[RA93] L. A. Rogers and J. Anderson. A new approach to defining a multidisciplinary field of science: The case of cardiovascular biology. Scientometrics, 28(1):61–77, September 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016285.

Radosevic:1998:MSA

[RA98] S. Radosevic and Laudeline Auriol. Measuring S&T activities in the former socialist economies of Central and Eastern Europe: Conceptual and methodological issues in linking past with present. Scientometrics, 42(3):273–297, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458372.

Rabinovich:1992:PSD

[Rab92] J. E. Rabinovich. Publications of scientists of developing countries in international journals: Are they channels to the international circuit for colleagues that only publish in national journals? Scientometrics, 23(1):91–103, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020916.

Raj:1996:PCP

[Raj96a] C. B. Chidambara Raj. Publications, cost and a proposition. Scientometrics, 37(1):87–103, September 1996. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02093486.

Rajeswari:1996:IPS

[Raj96b] A. R. Rajeswari. Indian patent statistics — an analysis. Scientometrics, 36(1):109–130, May 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02126649.

Rao:1994:LSB

[Rao94] I. K. Ravichandra Rao. Little scientometrics, big scientometrics … and beyond. Scientometrics, 30(2–3):465–469, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018126.

Rao:1996:MCQ

[Rao96] I. K. Ravichandra Rao. Methodological and conceptual questions of bibliometric standards. Scientometrics, 35(2):265–270, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02018484.

Rao:1998:ABM

[Rao98] I. K. Ravichandra Rao. An analysis of Bradford multipliers and a model to explain law of scattering. Scientometrics, 41(1– 2):93–100, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457970.

Ruiz-Banos:1999:SDSa

[RBBMJCC99a] R. Ruiz-Baños, R. Bailón-Moreno, E. Jiménez-Contreras, and J.-P. Courtial. Structure and dynamics of scientific networks. Part I: Fundamentals of the quantitative model of translation. Scientometrics, 44(2):217–234, February 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457381.

Ruiz-Banos:1999:SDSb

[RBBMJCC99b] R. Ruiz-Baños, R. Bailón-Moreno, E. Jiménez-Contreras, and J.-P. Courtial. Structure and dynamics of scientific networks. Part II: The new Zipf’s Law, the clusters of co-citations and the model of the descriptor presence. Scientometrics, 44(2): 235–265, February 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457382.

Reguant:1994:ELF

[RC94] S. Reguant and J. Casadellà. English as lingua franca in geological scientific publications. A bibliometric analysis. Scientometrics, 29(3):335–351, March 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02033444.

Rivas:1996:IDD

[RDC^{+}96] A. L. Rivas, D. Deshler, R. D. Colle, R. N. González, and F. W. Quimby. Indicators of disciplinary differentiation: Interdisciplinary linkages and adoption rate of biological innovations. Scientometrics, 37(1):63–86, September 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093485.

Rinia:1993:MNO

[RDM93] E. J. Rinia, C. De Lange, and H. F. Moed. Measuring national output in physics: Delimitation problems. Scientometrics, 28(1):89–110, September 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016287.

Rivas:1998:IQG

[RDQ^{+}98] A. L. Rivas, J. D. Deshler, F. W. Quimby, H. O. Mohammed, D. J. Wilson, R. N. Gonzalez, D. H. Lein, and P. Bruso. Interdisciplinary question generation: Synthesis and validity analysis of the 1993–1997 bovine mastitis-related literature. Scientometrics, 42(3):377–403, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458378.

Reyniers:1992:FFI

[Rey92] P. Reyniers. Facts and figures on interlibrary lending in Dutch-speaking Belgium in 1989. Scientometrics, 25(1):59– 76, September 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016847.

Roche:1992:RTV

[RF92] M. Roche and Y. Freites. Rise and twilight of the Venezuelan scientific community. Scientometrics, 23(2):267–289, February 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017042.

Raina:1998:FAI

[RG98] Dhruv Raina and B. M. Gupta. Four aspects of the institutionalization of physics research in India (1990–1950): Substantiating the claims of histortical sociology through bibliometrics. Scientometrics, 42(1):17–40, May 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02465010.

Roosendaal:1999:SCR

[RG99] H. E. Roosendaal and P. A. Th. M. Geurts. Scientific communication and its relevance to research policy. Scientometrics, 44(3):507–519, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02458492.

Raina:1995:CIP

[RGK95] Dhruv Raina, B. M. Gupta, and Rohit Kandhari. Collaboration in Indian physics: A case study of the macro and micro parametrization of sub-disciplines (1800–1950). Scientometrics, 33(3):295–314, July 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017333.

Rau:1990:RED

[RH90] E. Rau and T. Hummel. Rankings of economics departments in the Federal Republic of Germany. Scientometrics, 19(5–6): 377–384, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020700.

Richards:1991:YCA

[Ric91] J. M. Richards, Jr. Years cited: An alternative measure of scientific accomplishment. Scientometrics, 20(3):427–438, March 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019763.

Rinaldini:1995:ERE

[Rin95] C. Rinaldini. Experience on research evaluation at the joint research centre of the European commission. Scientometrics, 34(3):519–525, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018018.

Rip:1997:QCS

[Rip97] A. Rip. Qualitative conditions of scientometrics: The new challenges. Scientometrics, 38(1):7–26, January 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02461120.

Rikken:1995:MDA

[RKV95] F. Rikken, H. A. L. Kiers, and R. Vos. Mapping the dynamics of adverse drug reactions in subsequent time periods using INDSCAL. Scientometrics, 33(3):367–380, July 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017337.

Roman:1994:STD

[RM94] Adelaida Román and Aida Méndez. The Spanish transition to democracy seen through the Spanish database isoc. Scientometrics, 30(1):201–211, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017223.

Rodriguez:1996:GDR

[RM96] Ketty Rodríguez and J. A. Moreiro. The growth and development of research in the field of ecology. Scientometrics, 35 (1):59–70, January 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018233.

Rotto:1997:EEB

[RM97] E. Rotto and R. P. Morgan. An exploration of expert-based text analysis techniques for assessing industrial relevance in U.S. engineering dissertation abstracts. Scientometrics, 40(1): 83–102, September 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459263.

Rey:1998:CPB

[RMP^{+}98] J. Rey, M.-J. Martín, L. Plaza, J. J. Ibáñez, and Isabel Méndez. Changes on publishing behavior in response to research policy guidelines. The case of the Spanish Research Council in the field of agronomy. Scientometrics, 41(1–2): 101–111, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457971.

Rojo:1999:PRE

[Roj99] J. M. Rojo. Peer review — experiences at national and European level. Scientometrics, 45(3):497–500, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457616.

Rousseau:1990:BR

[Rou90] R. Rousseau. Book review. Scientometrics, 19(3–4):325–326, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02095354.

Rousseau:1992:CTI

[Rou92] R. Rousseau. Category theory and informetrics: Information production processes. Scientometrics, 25(1):77–87, September 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016848.

Rousseau:1993:MCS

[Rou93] R. Rousseau. Measuring concentration: Sampling design issues, as illustrated by the case of perfectly stratified samples. Scientometrics, 28(1):3–14, September 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016281.

Rousseau:1994:DEM

[Rou94a] R. Rousseau. Double exponential models for first-citation processes. Scientometrics, 30(1):213–227, May 1994. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02017224.

Rousseau:1994:SBI

[Rou94b] R. Rousseau. Similarities between informetrics and econometrics. Scientometrics, 30(2–3):385–387, June 1994. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02018108.

Rousseau:1998:CAT

[Rou98] R. Rousseau. Citation analysis as a theory of friction of polluted air? Scientometrics, 43(1):63–67, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458395.

Rousseau:1999:TDS

[Rou99a] R. Rousseau. Temporal differences in self-citation rates of scientific journals. Scientometrics, 44(3):521–531, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458493.

Routti:1999:CR

[Rou99b] J. Routti. Closing remarks. Scientometrics, 45(3):569–570, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02457632.

Rowlands:1999:PAC

[Row99] I. Rowlands. Patterns of author cocitation in information policy: Evidence of social, collaborative and cognitive structure. Scientometrics, 44(3):533–546, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458494.

Roseboom:1992:MDN

[RP92] J. Roseboom and P. G. Pardey. Measuring the development of national agricultural research systems. Scientometrics, 23(1): 169–190, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020921.

Rousseau:1997:DEA

[RR97] Sandra Rousseau and R. Rousseau. Data envelopment analysis as a tool for constructing scientometric indicators. Scientometrics, 40(1):45–56, September 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02459261.

Rousseau:1998:SWE

[RR98] Sandra Rousseau and R. Rousseau. The scientific wealth of European nations: Taking effectiveness into account. Scientometrics, 42(1):75–87, May 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02465013.

Rey-Rocha:1999:RDJ

[RRMS99] J. Rey-Rocha and M. J. Martín-Sempere. The role of domestic journals in geographically-oriented disciplines: The case of Spanish journals on earth sciences. Scientometrics, 45(2): 203–216, June 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458433.

Rao:1999:QSI

[RS99] I. K. Ravichandra Rao and Ms. P. Suma. A quantitative study of Indian engineering literature. Scientometrics, 46(3):605– 619, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459615.

Romanov:1995:MMS

[RT95] A. K. Romanov and A. I. Terekhov. The mathematical modelling of the scientific personnel movement taking into account the productivity factor. Scientometrics, 33(2):221–231, June 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020570.

Romanov:1997:MMP

[RT97] A. K. Romanov and A. I. Terekhov. The mathematical model of productivity- and age-structured scientific community evolution. Scientometrics, 39(1):3–17, May 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457427.

Rubio:1992:SPS

[Rub92] A. Villagrá Rubio. Scientific production of Spanish universities in the fields of social sciences and language. Scientometrics, 24(1):3–19, May 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02026470.

Russell:1994:BFI

[Rus94] Jane M. Russell. Back to the future for informetrics. Scientometrics, 30(2–3):407–410, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018113.

Russell:1995:IRI

[Rus95] Jane M. Russell. The increasing role of international cooperation in science and technology research in Mexico. Scientometrics, 34(1):45–61, September 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02019172.

Russell:1998:PPM

[Rus98] Jane M. Russell. Publishing patterns of Mexican scientists: Differences between national and international papers. Scientometrics, 41(1–2):113–124, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457972.

Rikken:1994:SAD

[RV94] Floor Rikken and R. Vos. Searching for adverse drug reactions at the margin of scientific fields. Scientometrics, 30 (1):187–199, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017222.

Rivas:1997:ISB

[RWG^{+}97] A. L. Rivas, D. J. Wilson, R. N. Gonzalez, H. O. Mohammed, F. W. Quimby, D. H. Lein, R. A. Milligan, R. D. Colle, J. D. Deshler, and W. M. K. Trochim. An interdisciplinary and systems-based evaluation of academic programs: Bovine mastitis-related veterinary research, education and outreach. Scientometrics, 40(2):195–213, October 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457437.

Rousseau:1992:ZDF

[RZ92] R. Rousseau and Qiaoqiao Zhang. Zipf’s data on the frequency of Chinese words revisited. Scientometrics, 24(2):201– 220, June 1992. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02017909.

Singh:1991:PCP

[SA91] U. N. Singh and S. Arunachalam. Publication and citation patterns in the literature of liquid crystals with special reference to the contribution of India, Canada, Japan, United Kingdom and the Soviet Union. Scientometrics, 20(1):197–220, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018155.

Simonetti:1995:PPI

[SAE95] R. Simonetti, D. Archibugi, and R. Evangelista. Product and process innovations: How are they defined? How are they quantified? Scientometrics, 32(1):77–89, January 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020190.

Sampson:1995:AC

[Sam95] Zora J. Sampson. Authorship counts. Scientometrics, 32(2): 219–226, February 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016895.

Sancho:1992:MSM

[San92] Rosa Sancho. Misjudgments and shortcomings in the measurement of scientific activities in less developed countries. Scientometrics, 23(1):221–233, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020924.

Sangam:1996:N

[San96] S. L. Sangam. News. Scientometrics, 36(1):141–142, May 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02126651.

Sangam:1999:OLF

[San99] S. L. Sangam. Obsolescence of literature in the field of psychology. Scientometrics, 44(1):33–46, January 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458476.

Schubert:1990:ICS

[SB90] A. Schubert and T. Braun. International collaboration in the sciences 1981–1985. Scientometrics, 19(1–2):3–10, July 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02130461.

Schubert:1992:TSE

[SB92] A. Schubert and T. Braun. Three scientometric études on developing countries as a tribute to Michael Moravcsik. Scientometrics, 23(1):3–19, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020910.

Schubert:1993:RSC

[SB93] A. Schubert and T. Braun. Reference standards for citation based assessments. Scientometrics, 26(1):21–35, January 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016790.

Schubert:1996:CFN

[SB96a] A. Schubert and T. Braun. Cross-field normalization of scientometric indicators. Scientometrics, 36(3):311–324, July/ August 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02129597.

Six:1996:BAP

[SB96b] J. Six and Martha Cecilia Bustamante. Bibliometric analysis of publications in experimental particle physics on cosmic rays and with accelerators. Scientometrics, 37(1):25–37, September 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093483.

Spangenberg:1990:SICa

[SBA90] J. F. A. Spangenberg, W. Buijink, and W. Alfenaar. Some incentives and constraints of scientific performance in departments of economics. Scientometrics, 18(3–4):241–268, March 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017764.

Sancho:1993:ACS

[SBG93] Rosa Sancho, G. Bernal, and L. Gálvez. Approach to the Cuban scientific activity by using publication based quantitative indicators (1985–1989). Scientometrics, 28(3):297– 312, November/December 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02026512.

Spangenberg:1990:SICb

[SBNA90] J. F. A. Spangenberg, B. Breemhaar, F. Nijhuis, and W. Alfenaar. Some incentives and contraints of scientific performance in economics. Scientometrics, 18(3–4):269–279, March 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017765.

Sen:1998:BST

[SC98] S. K. Sen and S. K. Chatterjee. Bibliographic scattering and time: An empirical study through temporal partitioning of bibliographies. Scientometrics, 41(1–2):135–154, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457974.

Scapagnini:1999:WOR

[Sca99] U. Scapagnini. Welcoming and opening remarks. Scientometrics, 45(3):339–341, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02457586.

Schubert:1990:QSSa

[Sch90a] A. Schubert. Quantitative studies of science a current bibliography. Scientometrics, 18(1–2):157–168, January 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019167.

Schubert:1990:QSSb

[Sch90b] A. Schubert. Quantitative studies of science a current bibliography. Scientometrics, 18(5–6):445–463, May 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020156.

Schubert:1991:QSS

[Sch91] A. Schubert. Quantitative studies of science: a current bibliography. Scientometrics, 20(3):451–457, March 1991.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019765.

Schmoch:1993:TKT

[Sch93] U. Schmoch. Tracing the knowledge transfer from science to technology as reflected in patent indicators. Scientometrics, 26(1):193–211, January 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016800.

Schubert:1994:BR

[Sch94a] A. Schubert. Book review. Scientometrics, 29(1):173–177, January 1994. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02018389.

Schubert:1994:LSB

[Sch94b] A. Schubert. Little scientometrics, big scientometrics — and beyond. Scientometrics, 30(2–3):411–413, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018114.

Schubert:1995:QSS

[Sch95] A. Schubert. Quantitative studies of science in the eighties indexes to “current bibliographies” 1–17. Scientometrics, 32 (3):263–367, March 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017646.

Schubert:1996:BCBa

[Sch96a] A. Schubert. Scientometrics: A citation based bibliography, 1990. Scientometrics, 35(1):155–163, January 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018238.

Schubert:1996:BCBb

[Sch96b] A. Schubert. Scientometrics: A citation based bibliography, 1991. Scientometrics, 35(3):393–399, March 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016910.

Schubert:1996:BCBc

[Sch96c] A. Schubert. Scientometrics: A citation based bibliography, 1992. Scientometrics, 36(1):131–140, May 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02126650.

Schubert:1996:BCBd

[Sch96d] A. Schubert. Scientometrics: A citation based bibliography, 1993. Scientometrics, 36(2):273–280, June 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017321.

Schmoch:1997:IRB

[Sch97a] U. Schmoch. Indicators and the relations between science and technology. Scientometrics, 38(1):103–116, January 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02461126.

Schummer:1997:SSCa

[Sch97b] J. Schummer. Scientometric studies on chemistry i: The exponential growth of chemical substances, 1800–1995. Scientometrics, 39(1):107–123, May 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02457433.

Schummer:1997:SSCb

[Sch97c] J. Schummer. Scientometric studies on chemistry II: Aims and methods of producing new chemical substances. Scientometrics, 39(1):125–140, May 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02457434.

Scharnhorst:1998:CNS

[Sch98] Andrea Scharnhorst. Citation-networks, science landscapes and evolutionary strategies. Scientometrics, 43(1):95–106, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02458399.

Schmidt:1999:REP

[Sch99a] A. Schmidt. Role of experts in political consultancy processes. Scientometrics, 45(3):475–477, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457611.

Schubert:1999:SJS

[Sch99b] A. Schubert. On science journals in science journals 1980– 1988. Scientometrics, 46(1):171–212, September 1999. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02766301.

Schubert:1999:SCB

[Sch99c] A. Schubert. Scientometrics: A citation based bibliography 1994–1996. Scientometrics, 44(2):267–315, February 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457383.

Souza:1993:MAQ

[SCQ93] G. S. Souza, E. R. Cruz, and T. R. Quirino. The measurement and assessment of quality in agricultural research institutions. Scientometrics, 28(2):159–182, October 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016898.

Seglen:1996:QSA

[Seg96] Per O. Seglen. Quantification of scientific article contents. Scientometrics, 35(3):355–366, March 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016906.

Seidel:1999:RNA

[Sei99] H. Seidel. The role of national agencies in evaluation. Scientometrics, 45(3):401–403, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02457597.

Senter:1993:FAS

[Sen93] R. Senter, Jr. Factors in American state government spending on research and development. Scientometrics, 28(3):313– 327, November/December 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02026513.

Schubert:1991:PDM

[SG91] A. Schubert and W. Glänzel. Publication dynamics: Models and indicators. Scientometrics, 20(1):317–331, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018161.

Schubert:1990:SDS

[SGB90] A. Schubert, W. Glänzel, and T. Braun. Scientometric datafiles supplementary indicators on 96 countries 1981– 1985. Scientometrics, 18(3–4):173–177, March 1990. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02017759.

Sengupta:1991:CUC

[SH91] I. N. Sengupta and R. G. Henzler. Citedness and uncitedness of cancer articles. Scientometrics, 22(2):283–296, October 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020002.

Schwartz:1996:MIS

[SH96] S. Schwartz and J. Lopez Hellin. Measuring the impact of scientific publications. The case of the biomedical sciences. Scientometrics, 35(1):119–132, January 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02018236.

Su:1998:NLG

[SH98] Yi Su and Li-Feng Han. A new literature growth model: Variable exponential growth law of literature. Scientometrics, 42 (2):259–265, June 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458359.

Shaw:1990:CBC

[Sha90] Alexandra Shaw. Comments on Bertram C. Brookes, recipient of the 1989 Derek de Solla Price Award. Scientometrics, 19 (3–4):153–155, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02095342.

Shrum:1997:VAV

[Shr97] Wesley Shrum. View from afar: ‘visible’ productivity of scientists in the developing world. Scientometrics, 40(2):215–235, October 1997. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02457438.

Sigogneau:1996:BPC

[Sig96] Anne Sigogneau. Between policy categories and research activities: Reviews and journals to describe “environment” networks. Scientometrics, 37(2):337–348, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093628.

Sikorav:1991:USP

[Sik91] J.-L. Sikorav. The utility of scientific papers. Scientometrics, 21(1):49–68, May 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019182.

Sikka:1997:SPS

[Sik97] Pawan Sikka. Statistical profile of science and technology in India and Brazil. Scientometrics, 39(2):185–195, June 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457447.

Sirilli:1999:IIS

[Sir99] G. Sirilli. Innovation indicators in science and technology evaluation. Scientometrics, 45(3):439–443, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457604.

Sengupta:1991:BAA

[SK91] I. N. Sengupta and Lalita Kumari. Bibliometric analysis of AIDS literature. Scientometrics, 20(1):297–315, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018160.

Skalicky:1999:UEI

[Ska99] P. Skalicky. Universities — engines of innovation in the information society. Scientometrics, 45(3):565–566, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02457630.

Skoie:1999:BSW

[Sko99] H. Skoie. Bibliometrics — some warnings from the North. Scientometrics, 45(3):433–437, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457603.

Sen:1992:IPS

[SL92] B. K. Sen and V. V. Lakshmi. Indian periodicals in the science citation index. Scientometrics, 23(2):291–318, February 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017043.

Stephan:1993:ANP

[SL93] Paula E. Stephan and Sharon G. Levin. Age and the Nobel Prize revisited. Scientometrics, 28(3):387–399, November/ December 1993. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02026517.

Small:1993:MLC

[Sma93] H. Small. Macro-level changes in the structure of co-citation clusters: 1983–1989. Scientometrics, 26(1):5–20, January 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016789.

Small:1994:SMC

[Sma94] H. Small. A SCI–Map case study: Building a map of AIDS research. Scientometrics, 30(1):229–241, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017225.

Small:1997:USM

[Sma97a] H. Small. Update on science mapping: Creating large document spaces. Scientometrics, 38(2):275–293, February 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457414.

Small:1997:CBC

[Sma97b] Henry Small. Comments on Belver C. Griffith, recipient of the 1997 Derek de Solla Price Award. Scientometrics, 40(3): 359–362, November 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459284. See [Ano97a].

Small:1998:CCS

[Sma98a] H. Small. Citations and consilience in science. Scientometrics, 43(1):143–148, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02458403.

Small:1998:GFC

[Sma98b] H. Small. A general framework for creating large-scale maps of science in two or three dimensions: The SciViz system. Scientometrics, 41(1–2):125–133, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457973.

Smith:1995:ERT

[Smi95] W. A. Smith. Evaluating research, technology and development in Canadian industry: Meeting the challenges of industrial innovation. Scientometrics, 34(3):527–539, November 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018019.

Saavedra:1993:SAS

[SMPK93] F. Saavedra, Mary Rose Mackenzie, R. Pessot, and M. Krauskopf. Size and ageing of the scientific community in Chile. Scientometrics, 27(2):105–117, June 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016545.

Spangenberg:1990:HIP

[SN90] J. F. A. Spangenberg and F. J. N. Nijhuis. Human information processing in science. Scientometrics, 18(5–6):389–407, May 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020153.

Snizek:1995:SOU

[Sni95] W. E. Snizek. Some observations on the use of bibliometric indicators in the assignment of university chairs. Scientometrics, 32(2):117–120, February 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016888.

So:1990:OIA

[So90] C. Y. K. So. Openness index and affinity index: Two new citation indicators. Scientometrics, 19(1–2):25–34, July 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02130463.

So:1998:CRV

[So98] C. Y. K. So. Citation ranking versus expert judgment in evaluating communication scholars: Effects of research specialty size and individual prominence. Scientometrics, 41(3): 325–333, March 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459049.

Snizek:1991:TNC

[SOM91] W. E. Snizek, K. Oehler, and N. C. Mullins. Textual and nontextual characteristics of scientific papers: Neglected science indicators. Scientometrics, 20(1):25–35, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018141.

Spagnolo:1990:BSP

[Spa90] F. Spagnolo. Brazilian scientists’ publications and mainstream science: Some policy implications. Scientometrics, 18(3–4): 205–218, March 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017762.

Spasser:1997:MTP

[Spa97] M. A. Spasser. Mapping the terrain of pharmacy: Coclassification analysis of the International Pharmaceutical Abstracts database. Scientometrics, 39(1):77–97, May 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457431.

Sancho:1992:BAR

[SPC92] Rosa Sancho, A. Pastor, and E. Criado. Bibliometric approach to research performance in the field of refractory materials used in iron and steelmaking processes. Scientometrics, 24 (1):115–136, May 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02026476.

Spiegel:1990:IPS

[Spi90] H.-R. Spiegel. Initiatives for the promotion of science of science. Scientometrics, 19(5–6):505–512, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020711.

Suraud:1995:SDB

[SQRD95] M. G. Suraud, L. Quoniam, H. Rostaing, and H. Dou. On the significance of data bases keywords for a large scale bibliometric investigation in fundamental physics. Scientometrics, 33(1):41–63, May 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020774.

Schlie-Roosen:1990:QIF

[SR90] F. Schlie-Roosen. Quantitative indicators for federal government research and technology policy. Scientometrics, 19(5–6): 439–445, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020706.

Saam:1999:LLR

[SR99a] Nicole J. Saam and L. Reiter. Lotka’s Law reconsidered: The evolution of publication and citation distributions in scientific fields. Scientometrics, 44(2):135–155, February 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457376.

Seetharam:1999:GFS

[SR99b] Geetha Seetharam and I. K. Ravichandra Rao. Growth of food science and technology literature: A comparison of CFTRI, India and the world. Scientometrics, 44(1):59–79, January 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458478.

Srinivasan:1999:AIJ

[SRMP99] R. Srinivasan, Vidyalakshmi Raman, N. Meyyappan, and P. Pichappan. Assessment of the impact of the journal literature produced by Indian CSIR laboratories using subfield corrected impact. Scientometrics, 44(1):81–92, January 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458479.

Sen:1992:ERS

[SS92] B. K. Sen and K. Shailendra. Evaluation of recent scientific research output by a bibliometric method. Scientometrics, 23 (1):31–46, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020912.

Soderqvist:1994:SLS

[SS94] T. Söderqvist and A. M. Silverstein. Studying leadership and subdisciplinary structure of scientific disciplines. Scientometrics, 30(1):243–258, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017226.

Satyanarayana:1999:RSC

[SSS99] K. Satyanarayana, D. Srivastava, and V. Sreenivas. The relevance of short communication in scholarly journals: An empirical study. Scientometrics, 44(1):47–58, January 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458477.

Schwarz:1998:RRI

[SST98] A. Winkel Schwarz, S. Schwarz, and R. J. W. Tijssen. Research and research impact of a technical university — a bibliometric study. Scientometrics, 41(3):371–388, March 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459052.

Stevens:1990:FIB

[Ste90] G. Stevens. The flow of information between languages: An application of Price’s method. Scientometrics, 19(1–2):107– 126, July 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02130468.

Steinberg:1993:SBR

[Ste93] J. J. Steinberg. The state of biomedical radiation research as demonstrated by publications, funding and manpower activity: An analytical example of utilizing on-line medical informatics. Scientometrics, 27(3):283–294, July 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016943.

Stefaniak:1998:ICP

[Ste98] Barbara Stefaniak. International cooperation of polish researchers with partners from abroad: A scientometric study. Scientometrics, 41(1–2):155–167, January 1998. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02457975.

Stromholm:1999:PRE

[Str99] S. Strömholm. Peer review- experience at national and European level. Scientometrics, 45(3):491–495, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457615.

Saurer:1994:PNS

[SW94] W. Saurer and R. Weinberger. Planetary nebulae: Some statistics on a continuously growing field and its contributors. Scientometrics, 31(1):85–95, September 1994.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018103.

Schwechheimer:1999:HDS

[SW99] H. Schwechheimer and M. Winterhager. Highly dynamic specialities in climate research. Scientometrics, 44(3):547–560, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02458495.

Sylvain:1993:CRA

[Syl93] C. Sylvain. Canadian research activity in aquaculture: A bibliometric analysis. Scientometrics, 27(3):295–316, July 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016944.

Teichler:1991:EET

[Tei91] U. Teichler. Evaluation of the EC training fellowship programme based on a fellows’ questionnaire survey. Scientometrics, 21(3):343–365, July/August 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02093975.

Teitel:1994:PRE

[Tei94] S. Teitel. Patents, R&D expenditures, country size, and percapita income: An international comparison. Scientometrics, 29(1):137–159, January 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018387.

Thomas:1992:DEC

[Tho92a] Katharine S. Thomas. The development of eponymy; a case study of the Southern blot. Scientometrics, 24(3):405–417, July 1992. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02051038.

Thomas:1992:EPB

[Tho92b] Sandra M. Thomas. The evaluation of plant biomass research: A case study of the problems inherent in bibliometric indicators. Scientometrics, 23(1):149–167, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020920.

Thomas:1995:SEA

[Tho95] P. R. Thomas. Size effects in the assessment of disciplinecontribution scores: An example from the social sciences. Scientometrics, 33(2):203–220, June 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020569.

Tijssen:1993:SCS

[Tij93] R. J. W. Tijssen. A scientometric cognitive study of neural network research: Expert mental maps versus bibliometric maps. Scientometrics, 28(1):111–136, September 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016288.

Tkamp:1999:QMI

[Tka99] E. Vel Tkamp. Quantitative methods in industrial R&D. Scientometrics, 45(3):459–462, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457607.

Turner:1994:GOD

[TLG94] W. A. Turner, A. Lelu, and A. Georgel. Geode: Optimizing data flow representation techniques in a network information system. Scientometrics, 30(1):269–281, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017228.

Tomov:1996:CII

[TM96] D. T. Tomov and H. G. Mutafov. Comparative indicators of interdisciplinarity in modern science. Scientometrics, 37(2): 267–278, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02093624.

Tapaswi:1999:RSO

[TM99] M. P. Tapaswi and B. S. Maheswarappa. Ranking serials in oceanography: An analysis based on the Indian contributions and their citations. Scientometrics, 44(1):93–127, January 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458480.

Todorov:1990:CJV

[Tod90] R. Todorov. Comments on Jan Vlachy, recipient of the 1989 Derek de Solla Price Award. Scientometrics, 19(3–4):157–158, September 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02095343.

Todorov:1992:DCS

[Tod92] R. Todorov. Displaying content of scientific journals: A coheading analysis. Scientometrics, 23(2):319–334, February 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017044.

Todorov:1994:FIN

[Tod94] R. Todorov. Facts or imposed names on facts? Scientometrics, 30(2–3):461–464, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018125.

Todorovsky:1997:WTB

[Tod97] D. Todorovsky. On the working time budget of the university teacher. Scientometrics, 40(1):13–21, September 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459259.

Turner:1991:EIO

[TR91] W. A. Turner and F. Rojouan. Evaluating input/output relationships in a regional research network using co-word analysis. Scientometrics, 22(1):139–154, September 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019279.

Trimble:1991:LTC

[Tri91] Virginia Trimble. Long-term careers of astronomers with doctoral degrees from prestigious vs. non-prestigious universities. Scientometrics, 20(1):71–77, January 1991. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02018145.

Trimble:1996:PIL

[Tri96] Virginia Trimble. Productivity and impact of large optical telescopes. Scientometrics, 36(2):237–246, June 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017316.

Trofimenko:1990:SAT

[Tro90] A. P. Trofimenko. Scientometric analysis of the topical content of scientific research and its particularities. Scientometrics, 18 (5–6):409–435, May 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02020154.

Tague-Sutcliffe:1994:MFC

[TS94] Jean Tague-Sutcliffe. Modelling and forecasting contact time as a measure of item informativeness. Scientometrics, 30(1): 259–267, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017227.

Tsay:1995:ICP

[Tsa95] Ming-Yueh Tsay. The impact of the concept of post-industrial society and information society: A citation analysis study. Scientometrics, 33(3):329–350, July 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02017335.

Tsipouri:1991:EER

[Tsi91] L. J. Tsipouri. Effects of EC R&D policy on Greece: Some thoughts in view of the Stride Programme. Scientometrics, 21(3):403–416, July/August 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02093978.

Tsipouri:1999:ERL

[Tsi99] Lena Tsipouri. Evaluation at regional level-science and technology in structural funds. Scientometrics, 45(3):509–515, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02457620.

Tuppy:1999:IIB

[Tup99] H. Tuppy. Interdependencies and interactions between evaluation and decision making processes. Scientometrics, 45(3): 409–412, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457599.

Turner:1991:ISF

[Tur91] W. A. Turner. An introduction to scientometrics in France. Scientometrics, 22(1):5–8, September 1991. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02019272.

Turner:1994:WIM

[Tur94] W. A. Turner. What’s in an R: Infor metrics or infometrics? Scientometrics, 30(2–3):471–480, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018127.

Tijssen:1995:GSJ

[TV95] R. J. W. Tijssen and Th. N. Van Leeuwen. On generalising scientometric journal mapping beyond ISI’s journal and citation databases. Scientometrics, 33(1):93–116, May 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020776.

Tijssen:1998:GSB

[TV98] R. J. W. Tijssen and E. Van Wijk. The global science base of information and communication technologiges: Bibliometric analysis of ICT research papers. Scientometrics, 42(1):41– 60, May 1998. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02465011.

Todorov:1990:MAG

[TW90] R. Todorov and M. Winterhager. Mapping Australian geophysics: A co-heading analysis. Scientometrics, 19(1–2):35– 56, July 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02130464.

Todorov:1991:OMM

[TW91] R. Todorov and M. Winterhager. An overview of Mike Moravcsik’s publication activity in physics. Scientometrics, 20(1): 163–172, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018153.

Thomas:1998:IRR

[TW98] P. R. Thomas and D. S. Watkins. Institutional research rankings via bibliometric analysis and direct peer review: A comparative case study with policy implications. Scientometrics, 41(3):335–355, March 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02459050.

Urban:1997:CDR

[UH97] D. Urban and T. J. Hoban. Cognitive determinants of risk perceptions associated with biotechnology. Scientometrics, 40(2): 299–331, October 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457441.

Uzun:1993:CST

[UMÖ93] A. Uzun, A. Menard, and M. E. Özel. Citation status of Turkish physics publications in foreign journals: A global analysis. Scientometrics, 28(1):79–87, September 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016286.

Uzun:1996:PPT

[UÖ96] A. Uzun and M. E. Özel. Publication patterns of Turkish astronomers. Scientometrics, 37(1):159–169, September 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093492.

Ugolini:1997:APQ

[UPS97] Donatella Ugolini, S. Parodi, and L. Santi. Analysis of publication quality in a Cancer Research Institute. Scientometrics, 38(2):265–274, February 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02457413.

Urata:1990:IFA

[Ura90] Hiroaki Urata. Information flows among academic disciplines in Japan. Scientometrics, 18(3–4):309–319, March 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017767.

Urban:1996:QMP

[Urb96] D. Urban. Quantitative measurement of public opinions on new technologies. Scientometrics, 35(1):71–92, January 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018234.

Uzun:1990:QAT

[Uzu90] A. Uzun. A quantitative analysis of Turkish publication output in physics between 1938–1987. Scientometrics, 19(1–2): 57–73, July 1990. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02130465.

Uzun:1996:BAP

[Uzu96] A. Uzun. A bibliometric analysis of physics publications from Middle Eastern countries. Scientometrics, 36(2):259– 269, June 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02017319.

VanDerKruit:1994:CAF

[Van94] P. C. Van Der Kruit. A comparison of astronomy in fifteen member countries of the organisation for economic cooperation and development. Scientometrics, 31(2):155–172, October 1994. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02018558.

VanRaan:1996:IPF

[Van96] A. F. J. Van Raan. Introduction to the Proceedings of the Fourth International Conference on Science and Technology Indicators. Scientometrics, 37(2):191–193, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093620.

VanDalen:1997:MGD

[Van97a] H. P. Van Dalen. Measuring giants and dwarfs: Assessing the quality of economists. Scientometrics, 38(2):231–252, February 1997. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02457411.

VanderMeulen:1997:USI

[Van97b] B. J. R. Van der Meulen. The use of S&T indicators in science policy: Dutch experiences and theoretical perspectives from policy analysis. Scientometrics, 38(1):87–101, January 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02461125.

VanRaan:1997:SSA

[Van97c] A. F. J. Van Raan. Scientometrics: State-of-the-art. Scientometrics, 38(1):205–218, January 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02461131.

VanRaan:1998:IIC

[Van98] A. F. J. Van Raan. The influence of international collaboration on the impact of research results. Scientometrics, 42(3): 423–428, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02458380.

vanBorm:1992:ILS

[vB92] J. van Borm. From interlibrary lending statistics to clearinghouse: The use of ILL statistics in Belgium. Scientometrics, 25(1):89–100, September 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016849.

Viana-Baptista:1999:UEI

[VB99] J. Viana-Baptista. Universities — engines of innovation in the information society. Scientometrics, 45(3):547–550, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457626.

vanCaulil:1996:QUR

[vCMvdB96] G. F. van Caulil, C. A. M. Mombers, and F. C. H. D. van den Beemt. Quantifying the utilization of research: The difficulties and two models to evaluate the utilization of research results. Scientometrics, 37(3):433–444, November 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019257.

Vessuri:1995:RSA

[Ves95] Hebe Vessuri. Recent strategies for adding value to scientific journals in Latin America. Scientometrics, 34(1):139– 161, September 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019178.

VanHooydonk:1994:BAI

[VGMP^{+}94] G. Van Hooydonk, R. Gevaert, G. Milis-Proost, H. Van De Sompel, and K. Debackere. A bibliotheconomic analysis of the impact factors of scientific disciplines. Scientometrics, 30 (1):65–81, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02017213.

vanHeeringen:1999:RAB

[vH99] A. van Heeringen. The role of advisory bodies in evaluation. Scientometrics, 45(3):387–389, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457595.

Vinkler:1991:MTT

[Vin91a] P. Vinkler. Magic triangle for three relative impact indicators. Scientometrics, 21(1):143–146, May 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02019187.

Vinkler:1991:PCD

[Vin91b] P. Vinkler. Possible causes of differences in information impact of journals from different subfields. Scientometrics, 20(1): 145–161, January 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018152.

Vinkler:1993:PPR

[Vin93a] P. Vinkler. Percentage patent representation (PPR) bilateral patent balance (BPB) and patent dominancy (PD) indicators characterizing international patenting relations. Scientometrics, 27(1):97–103, May 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02017758.

Vinkler:1993:RCA

[Vin93b] P. Vinkler. Research contribution, authorship and team cooperativeness. Scientometrics, 26(1):213–230, January 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016801.

Vinkler:1994:MMC

[Vin94a] P. Vinkler. Model of manifested communication through publications. Scientometrics, 31(3):223–239, November 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016873.

Vinkler:1994:OFI

[Vin94b] P. Vinkler. The origin and features of information referenced in pharmaceutical patents. Scientometrics, 30(1):283–302, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017229.

Vinkler:1994:WIS

[Vin94c] P. Vinkler. Words and indicators. As scientometrics stands. Scientometrics, 30(2–3):495–504, June 1994. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02018130.

Vinkler:1995:SAE

[Vin95] P. Vinkler. Some aspects of the evaluation of scientific and related performances of individuals. Scientometrics, 32(2): 109–116, February 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016887.

Vinkler:1996:MQS

[Vin96a] P. Vinkler. Model for quantitative selection of relative scientometric impact indicators. Scientometrics, 36(2):223–236, June 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017315.

Vinkler:1996:RBR

[Vin96b] P. Vinkler. Relationships between the rate of scientific development and citations. The chance for citedness model. Scientometrics, 35(3):375–386, March 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016908.

Vinkler:1996:SPA

[Vin96c] P. Vinkler. Some practical aspects of the standardization of scientometric indicators. Scientometrics, 35(2):237–245, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02018481.

Vinkler:1997:RRS

[Vin97] P. Vinkler. Relations of relative scientometric impact indicators. The relative publication strategy index. Scientometrics, 40(1):163–169, September 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02459266.

Vinkler:1998:CIF

[Vin98a] P. Vinkler. Comparative investigation of frequency and strength of motives toward referencing. The reference threshold model. Scientometrics, 43(1):107–127, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458400.

Vinkler:1998:GPI

[Vin98b] P. Vinkler. General performance indexes calculated for research institutes of the Hungarian Academy of Sciences based on scientometric indicators. Scientometrics, 41(1–2):185–200, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02457977.

Vinogradov:1998:SOV

[Vin98c] A. E. Vinogradov. Scientists of old vintage support a ‘winterbiased birthday’ theory. Scientometrics, 41(3):417–420, March 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459055.

Vinogradov:1998:STA

[Vin98d] A. E. Vinogradov. Secular trend of academician aging. Scientometrics, 43(2):149–160, October 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02458404.

Vinkler:1999:RST

[Vin99] P. Vinkler. Ratio of short term and long term impact factors and similarities of chemistry journals represented by references. Scientometrics, 46(3):621–633, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02459616.

Vlachy:1994:SWD

[Vla94] J. Vlachý. Scientometrics — what to do? Scientometrics, 30(2–3):521–527, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02018134.

VanHooydonk:1998:MIF

[VMP98] G. Van Hooydonk and Greta Milis-Proost. Measuring impact by a full option method and the notion of bibliometric spectra. Scientometrics, 41(1–2):169–183, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457976.

Vogel:1997:IFI

[Vog97] E. E. Vogel. Impact factor and international collaboration in Chilean physics: 1987–1994. Scientometrics, 38(2):253–263, February 1997. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02457412.

vanRaan:1991:FGI

[vR91] A. F. J. van Raan. Fractal geometry of information space as represented by co-citation clustering. Scientometrics, 20(3): 439–449, March 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019764.

vanRaan:1994:LSB

[vR94] A. F. J. van Raan. Little scientometrics, big scientometrics … and beyond. Scientometrics, 30(2–3):529–531, June 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02018135.

vanRaan:1996:ABM

[vR96] A. F. J. van Raan. Advanced bibliometric methods as quantitative core of peer review based evaluation and foresight exercises. Scientometrics, 36(3):397–420, July/August 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02129602.

vanRaan:1998:MQS

[vR98] A. F. J. van Raan. In matters of quantitative studies of science the fault of theorists is offering too little and asking too much. Scientometrics, 43(1):129–139, September 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458401.

vanRaan:1999:ABM

[vR99] A. van Raan. Advanced bibliometric methods for the evaluation of universities. Scientometrics, 45(3):417–423, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457601.

VanLeeuwen:1993:AMA

[VT93a] Th. N. Van Leeuwen and R. J. W. Tijssen. Assessing multidisciplinary areas of science and technology: A synthetic bibliometric study of Dutch nuclear energy research. Scientometrics, 26(1):115–133, January 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02016796.

VanRaan:1993:NNN

[VT93b] A. F. J. Van Raan and R. J. W. Tijssen. The neural net of neural network research. Scientometrics, 26(1):169–192, January 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016799.

Wagner-Dobler:1994:FDL

[WD94] R. Wagner-Döbler. The frequency distribution of legal decision citations in the German jurisdiction. Scientometrics, 29(1): 15–26, January 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018381.

Wagner-Dobler:1995:WCA

[WD95] R. Wagner-Döbler. Were has the cumulative advantage gone? Some observations about the frequency distribution of scientific productivity, of duration of scientific participation, and of speed of publication. Scientometrics, 32(2):123–132, February 1995. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02016890.

Wagner-Dobler:1997:TDB

[WD97] R. Wagner-Döbler. Time dependencies of Bradford distributions: Structures of journal output in 20th-century logic and 19th-century mathematics. Scientometrics, 39(3):231– 252, July 1997. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02458528.

Wagner-Dobler:1998:SEE

[WD98] R. Wagner-Döbler. Scientometric evidence for the existence of long economic growth cycles in Europe 1500–1900. Scientometrics, 41(1–2):201–208, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02457978.

Wagner-Dobler:1999:WGB

[WD99] R. Wagner-Döbler. William Goffman’s “Mathematical approach to the prediction of scientific discovery” and its application to logic, revisited. Scientometrics, 46(3):635–645, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02459617.

Wagner-Dobler:1994:RID

[WDB94] R. Wagner-Döbler and J. Berg. Regularity and irregularity in the development of scientific disciplines: The case of mathematical logic. Scientometrics, 30(1):303–319, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017230.

Wagner-Dobler:1999:PQO

[WDB99] R. Wagner-Döbler and J. Berg. Physics 1800–1900: A quantitative outline. Scientometrics, 46(2):213–285, October 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02464778.

Westerheijden:1999:IIS

[Wes99] D. F. Westerheijden. Innovation indicators in science and technology evaluation: Comments from a higher education point of view. Scientometrics, 45(3):445–453, July 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457605.

Whitney:1992:ATW

[Whi92] Gretchen Whitney. Access to third world science in international scientific and technical bibliographic databases. Scientometrics, 23(1):201–219, January 1992. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02020923.

Whitney:1993:PAM

[Whi93] Gretchen Whitney. Patterns of authorship in major bibliographic databases: The European region. Scientometrics, 26(2):275–292, February 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02016220.

Wilson:1998:DSC

[Wil98] Concepción S. Wilson. Defining subject collections for informetric analyses: The effect of varying the subject aboutness level. Scientometrics, 41(1–2):209–223, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457979.

Wilhjelm:1999:REP

[Wil99a] N. Wilhjelm. Role of experts in political consultancy processes. Scientometrics, 45(3):473–474, July 1999. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02457610.

Wilson:1999:UOD

[Wil99b] Concepción S. Wilson. Using online databases to form subject collections for informetric analyses. Scientometrics, 46(3):647– 667, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459618.

Winclawska:1996:PSC

[Win96] Berenika M. Winclawska. Polish sociology citation index (principles for creation and the first results). Scientometrics, 35(3): 387–391, March 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016909.

Wouters:1994:PDC

[WL94] P. Wouters and L. Leydesdorff. Has Price’s dream come true: Is scientometrics a hard science? Scientometrics, 31(2):193– 222, October 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02018560.

Wouters:1997:I

[WL97] P. Wouters and L. Leydesdorff. Introduction. Scientometrics, 38(1):3–5, January 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02461119.

Wouters:1997:CCP

[Wou97] P. Wouters. Citation cycles and peer review cycles. Scientometrics, 38(1):39–55, January 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02461122.

Wouters:1998:SS

[Wou98] P. Wouters. The signs of science. Scientometrics, 41(1–2): 225–241, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02457980.

Wouters:1999:BHG

[Wou99] P. Wouters. Beyond the Holy Grail: From citation theory to indicator theories. Scientometrics, 44(3):561–580, March 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02458496.

Weingart:1990:WRD

[WSW90] P. Weingart, R. Sehringer, and M. Winterhager. Which reality do we measure? Scientometrics, 19(5–6):481–493, November 1990. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02020709.

Willems:1993:UBE

[WW93] J. Willems and E. Woudstra. The use by biologists and engineers of non-specialist information sources and its relation to their social involvement. Scientometrics, 28(2):205–216, October 1993. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02016900.

Yamazaki:1991:AOF

[Yam91] S. Yamazaki. Academic origin of the first professors in American medical schools before the civil war. Scientometrics, 22(3): 359–368, November 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019768.

Yamazaki:1994:RAL

[Yam94] Shigeaki Yamazaki. Research activities in life sciences in Japan. Scientometrics, 29(2):181–190, February 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017972.

Yamazaki:1995:RSL

[Yam95] S. Yamazaki. Refereeing system of 29 life science journals preferred by Japanese scientists. Scientometrics, 33(1):123– 129, May 1995. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02020778.

Yitzhaki:1991:NRB

[YBT91] M. Yitzhaki and D. Ben-Tamar. Number of references in biochemistry and other fields; a case study of the Journal of Biological Chemistry throughout 1910–1985. Scientometrics, 21(1):3–22, May 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02019179.

Yurtsever:1999:IRP

[YG99] Ersin Yurtsever and Sami Gülgöz. The increase in the rate of publications originating from Turkey. Scientometrics, 46(2): 321–336, October 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02464781.

Yitzhaki:1994:RTL

[Yit94] M. Yitzhaki. Relation of title length of journal articles to number of authors. Scientometrics, 30(1):321–332, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017231.

Yitzhaki:1997:VIT

[Yit97] M. Yitzhaki. Variation in informativity of titles of research papers in selected humanities journals: A comparative study. Scientometrics, 38(2):219–229, February 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457410.

Yitzhaki:1998:LPS

[Yit98] M. Yitzhaki. The ‘language preference’ in sociology: Measures of ‘language self-citation’, ‘relative own-language preference indicator’, and ‘mutual use of languages’. Scientometrics, 41(1–2):243–254, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/BF02457981.

Yuthavong:1993:CST

[YPC^{+}93] Y. Yuthavong, K. Phornsadja, A. Chungcharoen, T. O. Eisemon, and C. H. Davis. Communication strategies in tissue culture and seed research in Thailand. Scientometrics, 28(1): 41–60, September 1993. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02016284.

Zachos:1991:ROE

[Zac91] G. Zachos. Research output evaluation of two University departments in Greece with the use of bibliometric indicators. Scientometrics, 21(2):195–221, June 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02017569.

Zitt:1994:DMD

[ZB94] M. Zitt and E. Bassecoulard. Development of a method for detection and trend analysis of research fronts built by lexical or cocitation analysis. Scientometrics, 30(1):333–351, May 1994. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02017232.

Zitt:1996:RCC

[ZB96] M. Zitt and E. Bassecoulard. Reassessment of co-citation methods for science indicators: Effect of methods improving recall rates. Scientometrics, 37(2):223–244, October 1996. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02093622.

Zitt:1998:ISJ

[ZB98] M. Zitt and E. Bassecoulard. Internationalization of scientific journals: A measurement based on publication and citation scope. Scientometrics, 41(1–2):255–271, January 1998. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02457982.

Zitt:1999:ICV

[ZB99] M. Zitt and E. Bassecoulard. Internationalization of communication a view on the evolution of scientific journals. Scientometrics, 46(3):669–685, November 1999. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:/

/link.springer.com/article/10.1007/BF02459619.

Zitt:1991:SMD

[Zit91] M. Zitt. A simple method for dynamic scientometrics using lexical analysis. Scientometrics, 22(1):229–252, September 1991. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/10.
1007/BF02019282.

Zhu:1991:CDR

[ZMM91] Jialong Zhu, A. J. Meadows, and G. Mason. Citations and departmental research ratings. Scientometrics, 21(2):171–179, June 1991. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02017567.

Zitt:1996:SMI

[ZT96] M. Zitt and N. Teixeira. Science macro-indicators: some aspects of Ost experience. Scientometrics, 35(2):209–222, February 1996. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/BF02018479.

Zumelzu:1997:MEP

[Zum97] E. Zumelzu. Mainstream engineering publishing in Latin America: The Chilean experience. Scientometrics, 40(1):3– 12, September 1997. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/BF02459258.